### v5

# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

Abstract
Let $\tilde{W}$ be an extended affine Weyl group, $\mathbf{H}$ be the corresponding affine Hecke algebra over the ring $\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$, and $J$ be Lusztig’s asymptotic Hecke algebra, viewed as a based ring with basis $\{t_{w}\}$. Viewing $J$ as a subalgebra of the $(\mathbf{q}^{-\frac{1}{2}})$-adic completion of $\mathbf{H}$ via Lusztig’s map $\phi$, we use Harish-Chandra’s Plancherel formula for $p$-adic groups to show that the coefficient of $T_{x}$ in $t_{w}$ is a rational function of $\mathbf{q}$, with denominator depending only on the two-sided cell containing $w$, and dividing a power of the Poincaré polynomial of the finite Weyl group. As an application, we conjecture that these denominators encode more detailed information about the failure of the Kazhdan-Lusztig classification at roots of the Poincaré polynomial than is currently known.
Along the way, we show that upon specializing $\mathbf{q}=q>1$, the map from $J$ to the Harish-Chandra Schwartz algebra is injective. As an application of injectivity, we give a novel criterion for an Iwahori-spherical representation to have fixed vectors under a larger parahoric subgroup in terms of its Kazhdan-Lusztig parameter.
Keywords— Asymptotic Hecke algebra, Iwahori-Hecke algebra, Plancherel formula, parahoric subgroup
1 Introduction
Let $\tilde{W}$ be an affine Weyl group or extended affine Weyl group, and let $\mathbf{H}$ be its associated Hecke algebra over $\mathcal{A}:=\mathbb{C}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]$, where $\mathbf{q}$ is a formal variable. The representation theory of $\mathbf{H}$ is very well understood, behaving well and uniformly when $\mathbf{q}$ is specialized to any $q\in\mathbb{C}^{\times}$ that is not a root of the Poincaré polynomial $P_{W}$ of the finite Weyl group $W\subset\tilde{W}$.
When $\mathbf{q}$ is specialized to a prime power $q$, the category of finite-dimensional modules over the specialized algebra $H$ is equivalent to the category of admissible representations with nonzero Iwahori-fixed vector of some $p$-adic group. A form of local Langlands correspondence, the Deligne-Langlands conjecture, has been established by Kazhdan and Lusztig in [KLDeligneLanglands], where they classified modules over the generic algebra $\mathbf{H}$ using algebraic $K$-theory. A slightly different approach to this classification due to Ginzburg is explained in [CG]. In both treatments, a first step is to fix a central character. In particular, one must chose a complex number $q\in\mathbb{C}^{\times}$ by which $\mathbf{q}$ will act. Decomposing the $K$-theory of certain subvarieties of Springer fibres into irreducible representations of a certain finite group yields the standard modules. It can happen that the standard modules are themselves simple (for example, simple tempered representations, which play an essential role in the present paper, are of this form), but in general simple modules are obtained as a certain unique nonzero quotient of standard modules. This quotient exists when $q\in\mathbb{C}^{\times}$ is not a root of unity, but can be zero otherwise. Lusztig conjectured in [LRepAffHecke] that this classification would in fact hold whenever $q$ was not a root of $P_{W}$, and this result was proven by Xi in [XiJAMS]. Xi also showed that the classification fails in general at roots of the Poincaré polynomial, and presented this failure by giving an example related to a lack of simple $\mathbf{H}|_{\mathbf{q}=q}$-modules attached to the lowest two-sided cell. Our results in this paper explain that the lowest two-sided cell is, in a precise sense, maximally singular with respect to the parameter $\mathbf{q}$.
One way Lusztig expressed the uniformity in $q$ of the representation theory of the various algebras $\mathbf{H}|_{\mathbf{q}=q}$ is via the asymptotic Hecke algebra $J$. This is a $\mathbb{C}$-algebra (in fact, a $\mathbb{Z}$-algebra) $J$ with distinguished basis $\{t_{w}\}_{w\in\tilde{W}}$, equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$. In this way there is a map from $J$-modules to $\mathbf{H}$-modules, and Lusztig has shown in [affineIII] and [affineIV] that when $q$ is not a root of unity (other than $1$), that the specialized map $\phi_{q}$ induces a bijection between simple $\mathbf{H}|_{\mathbf{q}=q}$-modules and simple $J$-modules, these last being defined over $\mathbb{C}$. Moreover, he showed that when $P_{W}(q)\neq 0$, the map $\phi_{q}$ induces an isomorphism
$(\phi_{q})_{*}\colon K_{0}(J-\mathbf{Mod})\to K_{0}(\mathbf{H}|_{\mathbf{q}=q}-\mathbf{Mod})$
of Grothendieck groups. The map $\phi$ becomes a bijection after completing $\mathbf{H}$ and $J\otimes_{\mathbb{C}}\mathcal{A}$ by replacing $\mathcal{A}$ with $\mathbb{C}((\mathbf{q}^{-1/2}))$ and allowing infinite sums convergent in the $(\mathbf{q}^{-1/2})$-adic topology. In this way one can write a basis element $t_{w}$ as an infinite sum
$()^{\dagger}\circ\phi^{-1}(t_{w})=\sum_{x\in\tilde{W}}a_{x,w}T_{x},$ (1)
where each $a_{x,w}$ is a formal Laurent series in $\mathbf{q}^{-1/2}$, and $()^{\dagger}$ is the Goldman involution.
In light of the above, it is natural to ask how $a_{x,w}$ behaves when $\mathbf{q}$ is specialized to a root of unity.
1.1 The asymptotic Hecke algebra and $p$-adic groups
This paper is prompted by the work of Braverman and Kazhdan in [BK], who related the asymptotic Hecke algebra to harmonic analysis on $p$-adic groups. Specifically, in [affineIV], Lusztig relates simple $J$-modules to certain $\mathbf{H}\otimes_{\mathcal{A}}\mathbb{C}(\mathbf{q}^{-1/2})$-modules termed tempered because their definition is made in analogy with Casselman’s criterion for temperedness of $p$-adic groups. In [BK], Braverman and Kazhdan showed essentially that the analytic meaning of the word “tempered” can be substituted into Lusztig’s results from [affineIV].
More precisely, let $\mathbf{G}$ be a connected reductive group defined and split over a non-archimedean local field $F$ whose extended affine Weyl group is $\tilde{W}$. Then in [BK], Braverman and Kazhdan define a map expressing $J$ as sitting between the Iwahori-Hecke algebra of $G=\mathbf{G}(F)$ and the Harish-Chandra Schwartz algebra $\mathcal{C}$, and propose a spectral characterization of $J$ via the operator Payley-Wiener theorem, obtaining a diagram
[…]
where the outer vertical maps are Fourier transform $f\mapsto\pi(f)$ and the rings $\mathcal{E}^{I}$ and $\mathcal{E}^{I}_{t}$ of endomorphisms of forgetful functors to vector spaces are as described by the operator Paley-Wiener theorem, as we recall in Section 2.3.
The map $\eta$ is defined in [BK] and we will recall its definition below. In particular, it induces the map $\tilde{\phi}$, which then associates a Harish-Chandra Schwartz function to every element of $J$ such that $\eta(j)=\pi(\tilde{\phi}(j))$, giving another way of associating to $t_{w}$ an expression similar to (1).
This prompts several questions: whether $\eta$ (equivalently $\tilde{\phi}$) is injective, whether it is surjective, and the nature of the relationship between the Schwartz function $\tilde{\phi}(t_{w})$ and the expression (1).
1.1.1 Denominators in the affine Hecke algebra and injectivity of $\eta$
[…]
In particular, $\phi_{1}$ is injective. Using that the representation theory of $J$ is sufficiently independent of $q$, we then show in Corollary 4 that $\tilde{\phi}$ is injective for any $q>1$. This is obviously equivalent to
Theorem 1 (Corollary 4).
The map $\eta$ is injective for any $q>1$.
A weaker form of the following Theorem, which is the main result of this paper, was conjectured by Kazhdan.
Theorem 2.
Let $\tilde{W}$ be an affine Weyl group, $\mathbf{H}$ its affine Hecke algebra over $\mathcal{A}$, and $J$ its asymptotic Hecke algebra. Let $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ be Lusztig’s map.
  1. 1.
    For all $x,w\in\tilde{W}$, $a_{x,w}$ is a rational function of $\mathbf{q}$. The denominator of $a_{x,w}$ is independent of $x$. As a function of $w$, it is constant on two-sided cells.
  2. 2.
    There exists $N_{\tilde{W}}\in\mathbb{N}$ such that upon writing
    $()^{\dagger}\circ\phi^{-1}(t_{w})=\sum_{x\in\tilde{W}}a_{x,w}T_{x},$
    we have
    $P_{W}(\mathbf{q})^{N_{\tilde{W}}}a_{x,w}\in\mathcal{A}$
    for all $x,w\in\tilde{W}$.
  3. 3.
    If $d$ is a distinguished involution in the lowest two-sided cell, then $a_{1,d}=1/P_{W}(\mathbf{q})$ exactly.
In [D], the author proved Theorem 2 in type $\tilde{A}_{1}$, but with different conventions. To translate to the conventions of this paper, the reader should replace $j$ with the Goldman involution $()^{\dagger}$, and the completion of $\mathbf{H}$ with respect to the $C_{w}$ basis and positive powers of $\mathbf{q}^{1/2}$ with the completion of $\mathbf{H}$ with respect to the basis $\{(-1)^{\ell(w)}C^{\prime}_{w}\}_{w\in\tilde{W}}$ and negative powers of $\mathbf{q}^{1/2}$. Note also that we write $a_{x,w}$ instead of $a_{w,x}$ as in [D]. In [N], Neunhöffer described the coefficients $a_{x,w}$ for finite Weyl groups.
In future, it would be desirable to also treat the case of unequal parameters, where a result like that of [BDD] governing denominators is not yet available. On the other hand, when $G=\mathrm{GL}_{n}$, we are able to be slightly more precise than Theorem 2, while also not appealing to [BDD]. For this reason we treat the case $G=\mathrm{GL}_{n}$ separately as Theorem 13.
Our main tool is Harish-Chandra’s Plancherel formula for the $p$-adic group $G$ associated to $\mathbf{H}$ and the surjection of cocentres induced by Lusztig’s map $\phi$ after inverting $P_{W}(\mathbf{q})$ proved in [BDD] [^1] In fact, as proved by Bezrukavnikov-Braverman-Kazhdan in the appendix of loc. cit., $\phi_{q}$ induces an isomorphism whenever $q$ is not a root of unity, but we will not need this.. We invoke [BDD] only at the very end of our argument, which, absent [BDD], still proves that $a_{x,w}$ are rational functions with denominator depending only the two-sided cell containing $w$. We do so with an eye to future work dealing with Hecke algebras with unequal parameters.
In [D], the author related a conjecture of Kazhdan concerning the positivity of some coefficients related to the coefficients $a_{x,w}$. Historically, proofs of such positivity phenomena have also provided interpretations of the positive quantities in question. While we cannot currently prove the conjecture in [D], our results in Section 4 hint at a possible interpretation of $a_{1,d}$ for certain distinguished involutions $d$.
Remark 1.
It is tempting to conjecture the following more precise version of Theorem 2, based on the factorization $P_{\mathbf{M}_{\mathbf{P}}}(\mathbf{q})P_{\mathbf{G}/\mathbf{P}}(\mathbf{q})=P_{\mathbf{G}/\mathbf{B}}(\mathbf{q})$: for every Levi subgroup $\mathbf{M}_{\mathbf{P}}$ of $\mathbf{G}$ and all $\omega\in\mathcal{E}_{2}(M_{P})$, the formal degree $d(\omega)$ is a rational function of $q$ the denominator of which divides a power of $P_{\mathbf{M}_{\mathbf{P}}}(q)$. The integral over all induced twists $\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$ is a rational function of $q$ with denominator dividing a power of the Poincaré polynomial of the partial flag variety $(\mathbf{G}/\mathbf{P})(\mathbb{C})$. For example, if $G=\mathrm{GL}_{6}(F)$ and $M=\mathrm{GL}_{3}(F)\times\mathrm{GL}_{3}(F)$. In this case the integral itself (omitting the factor $C_{M}$ in the notation of Section 2.5) is
$\frac{1}{2\pi i}\frac{1}{2\pi i}\int_{\mathbb{T}}\int_{\mathbb{T}}\frac{(z_{1}-z_{2})(z_{1}-z_{2})}{(z_{1}-q^{3}z_{2})(z_{1}-q^{-3}z_{2})}\frac{\mathrm{d}z_{1}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}=\frac{(1-q^{3})^{2}}{(1-q^{6})^{2}q^{3}}+1=\frac{(1-q^{3})q^{-3}}{1+q^{3}}+1,$
and by [BottTu] Proposition 23.1, (with $\mathbf{q}=t^{2}$) we have
$P_{\mathbf{G}/\mathbf{P}}(\mathbf{q})=P_{G(3,6)}(\mathbf{q})=(1+\mathbf{q}^{2})(1+\mathbf{q}+\mathbf{q}^{2}+\mathbf{q}^{3}+\mathbf{q}^{4})(1+\mathbf{q}^{3}).$
In examples such as the above, this does indeed happen, but only after cancellation with some terms in the numerator. In general, we will not track numerators precisely enough to show this version of Theorem 2. We shall however see a limited demonstration of this behaviour in Corollary 1.
1.1.2 Denominators in the affine Hecke and the Kazhdan-Lusztig classification at roots of unity
The affine Hecke has a filtration by two-sided ideals
$\mathbf{H}^{\geq i}=\operatorname{span}{\left\{C_{w}\,\middle|\,a(w)\geq i\right\}},$
where $a$ is Lusztig’s $a$-function. As such, for any $q\in\mathbb{C}^{\times}$ and any simple $H=\mathbf{H}|_{\mathbf{q}=q}$-module $M$ there is an integer $a(M)$ such that $H^{\geq i}M\neq 0$ but $H^{\geq i+1}M=0$. Define $a(M)=i$ to be this integer. One can also define $a(E)=a(\mathbf{c}(E))$ where $E$ is a simple $J$-module and $J_{\mathbf{c}(E)}$ is the unique two-sided ideal not annihilating $E$.
The algebra $J$ linearizes the above filtration into an honest direct sum, and implements the almost-independence on $q\in\mathbb{C}^{\times}$ of the representation theory of $H=\mathbf{H}|_{\mathbf{q}=q}$ as follows.
Theorem 3 ([XiJAMS]).
  1. 1.
    Suppose that $q$ is not a root of the Poincaré polynomial of $\mathbf{G}$. Then for each simple $J$-module $E$, the $H$ module ${}^{\phi_{q}}E$ has a unique simple quotient $L$ such that $a(E)=a(L)$. For all other simple subquotients $L^{\prime}$ of $E$, we have $a(L^{\prime})<a(E)$.
    Equivalently, for all admissible triples $(u,s,\rho)$, the representation $K(u,s,\rho,q)$ of $H$ has a unique nonzero simple quotient $L=L(u,s,\rho,q)$ such that $a(L)=a(\mathbf{c}(u))$. That is, the Deligne-Langlands conjecture is true for $\mathbf{H}|_{\mathbf{q}=q}$.
  2. 2.
    If $q$ is a root of the Poincaré polynomial of $\mathbf{G}$, then the Deligne-Langlands conjecture is false for the lowest cell. That is, if $u=\{1\}$, then every simple subquotient $L^{\prime}$ of $K(u,s,\rho,q)$ has $a(L^{\prime})<a(\mathbf{c}_{0})$.
By Theorem 2, the coefficients $a_{1,d}$ have poles at every root of $P_{W}$, for all distinguished involutions $d$ in the lowest two-sided cell $\mathbf{c}_{0}$. On the other hand, as we show in Example 7, there do exist cells $\mathbf{c}\neq\mathbf{c}_{0}$ such that the coefficients $a_{x,w}$ are nonsingular at certain roots of $P_{W}$, for all $w\in\mathbf{c}$ and $x\in\tilde{W}$. We encode the hope that this is no accident as
Conjecture 1.
Let $\tilde{W}$ be an affine Weyl group, and let $q\in\mathbb{C}^{\times}$ be a root of $P_{W}$. Let $\mathbf{c}$ be a two-sided cell such that if $w\in\mathbf{c}$, then $a_{x,w}$ does not have a pole at $\mathbf{q}=q$ for any $x\in\tilde{W}$. Let $u=u(\mathbf{c})$. Then every standard module $K(u,s,\rho,q)$ has a unique simple quotient $L=L(u,s,\rho,q)$ such that $a(L)=a(E)$, where $E$ is the simple $J$ module corresponding to $K(u,s,\rho,q)$, in the notation of [affineIV]. Two such simple modules are isomorphic if and only if their corresponding triples are conjugate.
Note that in type $\tilde{A}_{n}$, the number of two-sided cells grows as $e^{\sqrt{n}}$, whereas the number of subsets of roots of $P_{W}$ is $2^{n(n+1)/2}$. For example in type $\tilde{A}_{1}$, there is only one root of $P_{W}=\mathbf{q}+1$, but there are two two-sided cells (and both are singular at $\mathbf{q}=-1$.) However, already in type $\tilde{A}_{3}$, one can see from Theorem 11 that the two-sided cell corresponding to the partition $4=2+2$ is not singular at two of the roots of $P_{A_{3}}(\mathbf{q})=(1+\mathbf{q})(1+\mathbf{q}+\mathbf{q}^{2})(1+\mathbf{q}+\mathbf{q}^{2}+\mathbf{q}^{3})$; see Example 7.
1.1.3 Application: representations with parahoric-fixed vectors
In Section 4, we use the existence of the action of the asymptotic Hecke algebra on tempered $G$-representations to give a simple criterion for the existence of vectors fixed under a parahoric subgroup of $G$: thm existence of parahoric-fixed vectors
Theorem 14.
Let $\pi=K(u,s,\rho)$ be a simple tempered $I$-spherical representation of $G$. Let $\mathcal{P}$ be a parahoric subgroup of $G$ and let $w_{\mathcal{P}}$ be the longest element in the corresponding subgroup of $\tilde{W}$. Let $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
  1. 1.
    If $\ell(w_{\mathcal{P}})>\dim\mathcal{B}_{u}^{\vee}$, then $\pi^{\mathcal{P}}=\{0\}$.
  2. 2.
    Conversely, let $u_{\mathcal{P}}$ be the unipotent conjugacy class corresponding to the two-sided cell containing $w_{\mathcal{P}}$. Then there exists a semisimple element $s\in Z_{G^{\vee}}(u_{\mathcal{P}})$, a Levi subgroup $M^{\vee}$ of $G^{\vee}$ minimal such that $(u_{\mathcal{P}},s)\in M^{\vee}$, and a discrete series representation $\omega\in\mathcal{E}_{2}(M)$ such that
    $\pi^{\mathcal{P}}=i_{P_{M}}^{G}(\omega\otimes\nu)^{\mathcal{P}}\neq\{0\}$
    for all $\nu$ non-strictly positive and the parameter of $\pi$ is $(u_{\mathcal{P}},s)$.
Thus starting from the regular unipotent class, $\pi(t_{w_{\mathcal{P}}})=0$ until reaching the unipotent attached to $w_{\mathcal{P}}$. At this unipotent, $\mathcal{P}$-fixed vectors are first encountered, and $t_{w_{\mathcal{P}}}$ acts by a nonzero projector with image contained in $\pi^{\mathcal{P}}$. For lower cells, it may still be the case that $\pi^{\mathcal{P}}\neq 0$, but $t_{w_{\mathcal{P}}}$ will act by zero on such representations, too. Therefore the nonzero action of $t_{w_{\mathcal{P}}}$ detects the precisely the “most regular” unipotent attached to $\mathcal{P}$-spherical representations, in the sense that if a representation $\pi$ such that $\pi^{\mathcal{P}}\neq 0$ has the unipotent part of its parameter equal to $u$, then $a(u)\geq a(u_{\mathcal{P}})$. In this way the distinguished involutions $t_{w_{\mathcal{P}}}$ are more exact versions of the corresponding indicator functions $1_{\mathcal{P}}$, at the expense of being more complicated to understand.
Remark 2.
Recall from Section 1.1.2 that for every simple $H$-module $M$ there is a number $a(M)$ such that $H^{\geq i}M=0$ for all $i>a(M)$. However, if $(u,s,\rho)$ is the $KL$-parameter of $M$, then without knowing that $M$ extends to a simple $J$-module, it does not follow that $a(M)=\dim\mathcal{B}_{u}^{\vee}$.
Remark 3.
By [affineIV, Theorem 4.8(d)] and the proof of [affineIV, Lemma 5.5], every two-sided cell contains a distinguished involution contained in a finite parabolic subgroup of $\tilde{W}$, but not every distinguished involution of a finite Coxeter group is the longest word of a parabolic subgroup, i.e. is of the form $w_{\mathcal{P}}$; approximately half of two-sided cells of the finite Weyl group $W\subset\tilde{W}$ do not contain any distinguished involutions contained in proper parabolic subgroups, because of the cell-preserving bijection $w\mapsto w_{0}w$. For example, this happens for the second-lowest cell for $E_{8}$.
The existence of parahoric-fixed vectors is a rigid question, in the sense of the rigid cocentre of Ciubotaru-He [CH]. We investigate this connection further in forthcoming work.
Some time after completing the present paper, we became aware of [GP], which also studies the connection between the asymptotic Hecke algebra and the Plancherel theorem in type $\tilde{G}_{2}$, for unequal parameters. In op. cit. the authors speculate that the “asymptotic Plancherel measure” of op. cit. should be related to the perspective of [BK] on $J$. We defer investigation of this to future work, but note that in light of both the classic work [Morris] of Morris, and recent work [SolleveldProGen] of Solleveld, the unequal parameters case is relevant even to split $p$-adic groups. In particular, establishing results similar to those of the present paper for unequal parameters may provide an effective way to study the algebra $\mathcal{J}$ of Braverman and Kazhdan given in Definition 1.9 of [BK].
1.2 Outline of the argument
This paper is organized according to our strategy for proving Theorems 2, Theorem 13, and Corollary 4.
These results are each simple corollaries of computations with the Plancherel formula and some of Lusztig’s results on $J$. The remainder of this section will introduce $\mathbf{H}$ and $J$ precisely, and recall their basic representation theory. In Section 2, we introduce Harish-Chandra’s Plancherel formula in detail, along with all the numerical constants that appear in it. In Section 2.4, we recall the results of Braverman-Kazhdan from [BK]. There is no original material in the first two sections. In Section 2.8, we prepare to apply the Plancherel formula by proving that, if $f_{w}$ is the Schwartz function associated by Braverman-Kazhdan to $t_{w}$, and $\pi$ is a tempered representation, then $\mathrm{trace}\left({\pi}\,,{f_{w}}\right)$ is sufficiently regular so as not to complicate the denominators of $a_{x,w}$. This section is also mostly a recollection of standard material, the only original result being Lemma 4.
In Section 3, we prove most of our main results. As we are able to be more precise in type $\tilde{A}_{n}$, we perform each step in parallel for type $\tilde{A}_{n}$ and for other types: in Sections 3.1 and 3.2 we prove statements like those of Theorem 2 for the Schwartz functions $f_{w}$. In these sections $\mathbf{q}$ is specialized to a prime power $q$. In Section 3.3 we relate the functions $f_{w}$ to the basis elements $t_{w}$, turning statements that hold for all prime powers $q$ into statements that hold for the formal variable $\mathbf{q}$. We are then able to prove our main results.
In Section 4, we state our application about the existence of parahoric-fixed vectors.
1.3 The affine Hecke algebra
Let $F$ be a non-archimedean local field, $\mathcal{O}$ its ring of integers and $\varpi$ be a uniformizer. Let $q$ be the cardinality of the residue field. Then $q=p^{r}$ is a prime power. We write $|\cdot|_{F}$ for the $p$-adic absolute value on $F$; when necessary, $|\cdot|_{\infty}$ will denote the archimedean absolute value on $\mathbb{C}$.
Let $\mathbf{G}$ be a connected reductive algebraic group defined and split over $F$, $\mathbf{A}$ a maximal torus of $\mathbf{G}$, and $X_{*}=X_{*}(\mathbf{A})$ the cocharacter lattice of $\mathbf{A}$. Let $\mathbf{N}$ be unipotent radical of a chosen Borel subgroup $\mathbf{B}$, so that $\mathbf{B}=\mathbf{A}\mathbf{N}$. Let $W$ be the finite Weyl group of $\mathbf{G}$, and $\tilde{W}=W\ltimes X_{*}(\mathbf{A})$ be the extended affine Weyl group. Write $S$ for the set of simple reflections in $\tilde{W}$. Let $\mathbf{G}^{\vee}$ be the Langlands dual group of $\mathbf{G}$, taken over $\mathbb{C}$. We write $G=\mathbf{G}(F)$, $A=\mathbf{A}(F)$, etc. Where there is no danger of confusion, we also write $G^{\vee}$ for $\mathbf{G}^{\vee}(\mathbb{C})$, $M^{\vee}$ for $\mathbf{M}^{\vee}(\mathbb{C})$, etc. Let $K$ be the maximal compact subgroup $\mathbf{G}(\mathcal{O})$. Also let $I$ be the Iwahori subgroup of $G$ that is the preimage of $\mathbf{B}(\mathbb{F}_{q})$. We sometimes write $P_{\mathbf{G}/\mathbf{B}}$ for $P_{W}$, as this polynomial is also the Poincaré polynomial of the flag variety $(\mathbf{G}/\mathbf{B})(\mathbb{C})$.
We write $\mathbf{H}$ for the affine Hecke algebra of $\tilde{W}$. It is a unital associative algebra over the ring $\mathcal{A}=\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ (in fact, it is defined over $\mathbb{Z}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ but we will work over $\mathbb{C}$ to avoid having to introduce extra notation later), where $\mathbf{q}^{\frac{1}{2}}$ is a formal variable. We will think of $\mathbb{C}^{\times}$ as $\operatorname{Spec}\mathcal{A}$. The algebra $\mathbf{H}$ has the Coxeter presentation with standard basis $\{T_{w}\}_{w\in\tilde{W}}$ with $T_{w}T_{w^{\prime}}=T_{ww^{\prime}}$ if $\ell(ww^{\prime})=\ell(w)+\ell(w^{\prime})$ and quadratic relation $(T_{s}+1)(T_{s}-\mathbf{q})=0$ for $s\in S$. We write $\theta_{\lambda}$ for the generators of the Bernstein subalgebra.
Recall from [KL79] the two Kazhdan-Lusztig bases $\{C_{w}\}_{w\in\tilde{W}}$ and $\{C^{\prime}_{w}\}_{w\in\tilde{W}}$ of $\mathbf{H}$ and write $C_{x}C_{y}=\sum_{z\in\tilde{W}}h_{x,y,z}C_{z}$. The inverse Kazhdan-Lusztig polynomials $Q_{y,x}$ are the unique family of polynomials satisfying
$T_{x}=\sum_{y\leq x}(-1)^{\ell(x)-\ell(y)}\mathbf{q}^{\frac{\ell(y)}{2}}Q_{y,x}(\mathbf{q})C^{\prime}_{y},$
or equivalently,
$\sum_{z\leq y\leq x}(-1)^{\ell(x)-\ell(y)}Q_{y,x}(\mathbf{q})P_{z,y}(\mathbf{q})=\delta_{z,x}$
along with some restrictions on their degrees. For example, we shall use in Section 3.3.2 that $\deg Q_{y,x}\leq\frac{1}{2}(\ell(x)-\ell(y)-1)$. See [BjornerBrenti] for further exposition.
It is well-known that there is an isomorphism of associative $\mathbb{C}$-algebras
$\mathbf{H}|_{\mathbf{q}=q}:=\mathbf{H}\otimes_{\mathcal{A}}\mathbb{C}\to C_{c}^{\infty}(G)^{I\times I}=:H,$
where $\mathbf{q}$ acts on $\mathbb{C}$ by multiplication by $q$.
1.4 The asymptotic Hecke algebra
Definition 1.
Lusztig’s a-function $a\colon\tilde{W}\to\mathbb{Z}_{\geq 0}$ is defined such that $a(w)$ is the minimal value such that $\mathbf{q}^{\frac{a(w)}{2}}h_{x,y,w}\in\mathcal{A}^{+}=\operatorname{Spec}\mathbb{C}[\mathbf{q}^{1/2}]$ for all $x,y\in\tilde{W}$.
The $a$-function is constant of two-sided cells of $\tilde{W}$. Obviously, $a(1)=1$, and the $a$-function obtains its maximum, equal to the number of positive roots, on the two-sided cell containing the longest word $w_{0}\in W$. In general, under the bijection between two-sided cells $\mathbf{c}$ of $\tilde{W}$ and unipotent conjugacy classes $u=u(\mathbf{c})$ in $G^{\vee}$ of [affineIV], we have
$a(\mathbf{c})=\dim_{\mathbb{C}}(\mathcal{B}_{u}^{\vee}),$
where $\mathcal{B}_{u}^{\vee}$ is the Springer fibre of $u$. We have $a(w)\leq\ell(w)$ for all $w\in\tilde{W}$.
In [affineII], Lusztig defined an associative algebra $J$ over $\mathbb{C}$ equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ which becomes an isomorphism after taking a certain completion, to be recalled in Section 3.3.1, of both sides. As an abelian group, $J$ is free with basis $\{t_{w}\}_{w\in\tilde{W}}$. The structure constants of $J$ are obtained from those of $\mathbf{H}$ written in the $\{C_{w}\}_{w\in\tilde{W}}$-basis under the following procedure: first, the integer $\gamma_{x,y,z}$ is defined by the condition
$\mathbf{q}^{\frac{a(z)}{2}}h_{x,y,z^{-1}}-\gamma_{x,y,z}\in\mathbf{q}\mathcal{A}^{+}.$
One then defines
$t_{x}t_{y}=\sum_{z\in\tilde{W}}\gamma_{x,y,z}t_{z^{-1}}.$
Surprisingly, this defines a unital associative algebra with unit
$1_{J}=\sum_{d\in\mathcal{D}}t_{d},$
where $\mathcal{D}$ is the (finite) set of distinguished involutions [affineII]. The elements $t_{d}$ are orthogonal idempotents in $J$, which decomposes as a direct sum indexed by the two-sided cells of $\tilde{W}$ in the sense of [affineI]. Each left cell, again in the sense of op. cit., contains a single distinguished involution which is the unit in the ring $t_{d}Jt_{d}$. If $\mathbf{c}$ is a two-sided cell, then $J_{\mathbf{c}}$ is a is a unital ring with unit
$1_{J_{\mathbf{c}}}=\sum_{d\in\mathcal{D}\cap\mathbf{c}}t_{d}.$
Lusztig further defined a map of algebras
$\phi\colon\mathbf{H}\to J\otimes_{\mathbb{C}}\mathcal{A}$
given by
$\phi(C_{w})=\sum_{z\in W,~{}d\in\mathcal{D},~{}a(z)=a(d)}h_{x,d,z}t_{z}.$
Write $\phi_{q}$ for the specialization of this map when $\mathbf{q}=q$. It is known [affineIII, Proposition 1.7] that $\phi_{q}$ is injective for all $q\in\mathbb{C}^{\times}$.
1.4.1 Deformations of the group ring
Upon setting $\mathbf{q}=1$, $\mathbf{H}|_{\mathbf{q}=1}$ is isomorphic to $\mathbb{C}[\tilde{W}]$, and so $\mathbf{H}$ is a deformation of the group algebra of $\tilde{W}$.
Let, temporarily, $W$ be any finite Coxeter group. Then one can define its Hecke algebra $\mathbf{H}$, an algebra over $\mathbb{Z}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ which deforms the group ring $\mathbb{Z}[W]$. Let $q\in\mathbb{C}^{\times}$. For all but finitely-many values of $q$, all roots of unity, the algebras $\mathbf{H}_{\mathbf{q}=q}$ are trivial deformations of $\mathbb{C}[W]$, and hence are all isomorphic. However, this isomorphism requires choosing a square root of $q$. The affine Hecke algebra provides a canonical isomorphism: away from finitely-many $q$, we have that $\mathbf{H}|_{\mathbf{q}=q}$ is isomorphic to $J$, and $J$ is defined over $\mathbb{Z}$ (although, as stated above, we will view it as a $\mathbb{C}$- algebra to unburden notation), see [LusztigBook, Section 20.1 (e)].
Example 1.
Let $W=\left\langle 1,s\,\middle|\,s^{2}=1\right\rangle$ be the Weyl group of type $A_{1}$. The Kazhdan-Lusztig $C_{w}$-basis elements are $C_{1}=T_{1}$ and $C_{s}=\mathbf{q}^{-\frac{1}{2}}T_{s}-\mathbf{q}^{\frac{1}{2}}T_{1}$. There are two two-sided cells in $W$, and one can easily check that
$\phi\colon C_{1}\mapsto t_{1}+t_{s}$
and
$\phi\colon C_{s}\mapsto-\left(\mathbf{q}^{\frac{1}{2}}+\mathbf{q}^{-\frac{1}{2}}\right)t_{s}.$
Specializing $\mathbf{q}=q$, we see that $\phi$ becomes an isomorphism whenever $\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right)\neq 0$, that is, whenever $q\neq-1$.
1.5 Representation theory of $\mathbf{H}$ and $J$
Recall the classification of finite-dimensional $\mathbf{H}$-modules given in [KLDeligneLanglands]. For an extended exposition with slightly different conventions, we refer the reader to [CG]. The primary difference between the setup we require and that of [CG] is that we must be able to defer specializing $\mathbf{q}$ until the last possible moment, whereas specializing $\mathbf{q}$ is the first step of the construction as given in [CG]. In particular, let $u\in\mathbf{G}^{\vee}(\mathbb{C})$ be a unipotent element, and $s\in\mathbf{G}^{\vee}(\mathbb{C})$ be a semisimple element such that $us=su$. Let $\rho$ be an irreducible representation of the simultaneous centralizer $Z_{G^{\vee}}(s,u)$. The standard $H$-module $K(s,u,\rho)$ is a certain, and in general reducible, $H$-module defined using the geometry of the flag variety of the Langlands dual group. It may be the zero module; we say that $(u,s,\rho)$ is admissible when this does not happen. Having fixed $s$ and $u$, we say that $\rho$ is admissible if $(u,s,\rho)$ is.
We now recall an algebraic version of the Langlands classification. As we do not have access to the notion of absolute value of $\mathbf{q}$, the classical definitions of tempered and discrete-series representations of the corresponding $p$-adic group $G$ are not available to us. However, Kazhdan-Lusztig provide the following algebraic generalization. Let $\kappa=\mathbb{C}(\mathbf{q}^{-\frac{1}{2}})$ and $\bar{\kappa}$ be the algebraic closure of $\kappa$. We write $\mathbf{H}_{\bar{\kappa}}$ for $\mathbf{H}\otimes_{\mathcal{A}}\bar{\kappa}$ and recall another definition of Lusztig’s from [affineIV].
Definition 2.
Let $M$ be a $H_{\bar{\kappa}}$-module finite-dimensional over $\bar{\kappa}$. Say that $m\in M$ is an eigenvector if $\theta_{x}\cdot m=\chi_{m}(x)m$ for all dominant $x$ in $X_{*}$. As $\chi_{m}$ is a character of the coweight lattice, it corresponds to an element $\sigma_{m}\in\mathbf{A}^{\vee}(\bar{\kappa})$ in the sense that, for all cocharacters $x$ of $A$, we have
$\chi_{m}(x)=x(\sigma_{m})$
where $x$ is viewed as a character of $\mathbf{A}^{\vee}$. Then $M$ is of constant type if there is a semisimple element $s^{\prime}\in\mathbf{G}^{\vee}(\mathbb{C})$ and a morphism of algebraic groups
$\phi^{\prime}\colon\mathrm{SL}_{2}(\mathbb{C})\to Z_{G^{\vee}}^{0}(s^{\prime})$
such that for all eigenvectors $m$ of $M$, the element $\sigma_{m}$ is $\mathbf{G}^{\vee}(\bar{\kappa})$-conjugate to
$\phi^{\prime}(\mathrm{diag}(\mathbf{q}^{1/2},\mathbf{q}^{-1/2}))s^{\prime},$
where by abuse of notation we have written $\phi^{\prime}$ again for the base-change to $\bar{\kappa}$.
The idea of the name of the definition is that $s^{\prime}\in\mathbf{G}^{\vee}(\mathbb{C})$ is a “constant element” not depending on $\mathbf{q}$.
Next is a generalization of Casselman’s criterion, which as such, requires a choice of dominant weights. The definitions in both [KLDeligneLanglands] and [affineIV] are both given under the choice of positive roots to be those associated to the opposite Borel (the so-called geometric choice of dominance). On the other hand, for purposes of harmonic analysis, Casselman’s criterion as usually stated uses the standard choice of positive roots.
Following [affineIV], we choose a morphism of groups $V\colon\bar{\kappa}^{\times}\to\mathbb{R}$ such that $V(\mathbf{q}^{-\frac{1}{2}})=1$ and $V(a\mathbf{q}^{-\frac{1}{2}}+b)=0$ for all $a\in\mathbb{C}$, $b\in\mathbb{C}^{\times}$.
Definition 3 ([affineIV], c.f. [KLDeligneLanglands]).
Let $M$ be any finite-dimensional $\mathbf{H}_{\bar{\kappa}}$-module and fix the geometric notion of dominance for $X_{*}(\mathbf{A})$. We say that $M$ is $V$-geometrically anti-tempered if all eigenvalues $\nu$ of $\theta_{\lambda}$ for all geometrically dominant $\lambda\in X_{*}(\mathbf{A})$ satisfy $V(\nu)\geq 0$. We say that $M$ is $V$-geometrically anti-square-integrable if $V(\nu)>0$.
When $\mathbf{q}$ is specialized a prime power $q\in\mathbb{C}^{\times}$ and $V(z)=\log|z|$, we recover that our anti-tempered for the geometric choice of dominance modules are tempered in the traditional analytic sense after twisting by the Goldman involution, and likewise for square-integrability, by the Casselman criterion.
The representation theory of $J$ is very well understood. We shall recall some notation and then state some major classification results of Lusztig, which relate the representation theory of $J$ to certain $\mathbf{H}$-modules defined by Kazhdan-Lusztig.
Definition 4.
Let $E$ be a $J$-module. Then $E\otimes_{\mathbb{C}}\kappa$ is a $J\otimes_{\mathbb{C}}\kappa$-module. Hence $\mathbf{H}_{\kappa}$ acts on $E$ via $\phi$. Denote the resulting $\mathbf{H}_{\kappa}$ module by ${}^{\phi}E$.
1.5.1 Involutions on $\mathbf{H}$
Definition 5.
Let $j\colon\mathbf{H}\to\mathbf{H}$ be the ring (and not $\mathcal{A}$-algebra) involution of $\mathbf{H}$ defined by $j(\sum_{w}a_{w}T_{w})=\sum_{w}\bar{a}_{w}(-1)^{\ell(w)}\mathbf{q}^{-\ell(w)}T_{w}$.
The $j$-involution exchanges the $\{C_{w}\}$ and $\{C_{w}^{\prime}\}$-bases [KL79].
The reason for our choice of conventions, which match those of [affineIII] and [affineIV], is the presence of the Goldman involution and its exchange of temperedness and anti-temperedness in the relationship between $H$-modules and $J$-modules; see Theorem 4 and Lemma 1 below.
Definition 6.
The Goldman involution $h\mapsto h^{\dagger}$ is the $\mathcal{A}$-algebra involution of $\mathbf{H}$ defined by setting
$T_{w}^{\dagger}=(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1}.$
Given an $\mathbf{H}$ (or $\mathbf{H}_{\kappa}$)-module $M$, define ${}^{\dagger}M$ to be the same vector space with the $\mathbf{H}$-action twisted by this involution.
Let $\Delta$ be the usual positive system determined by $\mathbf{B}$ and let $\Delta^{-}$ be the geometric system. Then we have
Lemma 1.
We have
  1. 1.
    $\overline{h^{\dagger}}=(\bar{h})^{\dagger}=j(h)$
    as $\mathcal{A}$-antilinear involutions. In particular, $C_{w}^{\dagger}=j(C_{w})=(-1)^{\ell(w)}C^{\prime}_{w}$.
  2. 2.
    Let $\lambda$ be dominant for $\Delta$ and let $\theta^{\Delta}_{\lambda}$ and $\theta^{\Delta^{-}}_{\lambda}$ denote Bernstein elements with respect to $\Delta$ and $\Delta^{-}$, respectively. Then
    ${\theta^{\Delta}_{\lambda}}^{\dagger}=(-1)^{\ell(\lambda)}\theta^{\Delta^{-}}_{\lambda}.$
Proof.
We compute
$\overline{\sum_{x}b_{x}T_{x}^{\dagger}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\overline{\mathbf{q}^{\ell(x)}\bar{T}_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}=j\left(\sum_{x}b_{x}T_{x}\right)$
whereas
$\left(\overline{\sum_{x}b_{x}T_{x}}\right)^{\dagger}=\left(\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}^{\dagger}\right)^{\dagger}.$
Thus we have $\bar{C_{w}^{\dagger}}=(-1)^{\ell(w)}C^{\prime}_{w}$, whence the first claim.
Next,
${\theta^{\Delta}_{\lambda}}^{\dagger}=(-1)^{\ell(\lambda)}\mathbf{q}^{-\frac{\ell(\lambda)}{2}}\mathbf{q}^{\ell(\lambda)}T_{-\lambda}^{-1}=(-1)^{\ell(\lambda)}\theta^{\Delta^{-}}_{\lambda},$
noting that $-\lambda$ is dominant with respect to $\Delta^{-}$ and so can appear as a label with respect to $\Delta^{-}$. ∎
We now summarize the relationship between representations of $\mathbf{H}$ and of $J$.
Theorem 4 ([affineIV], Prop. 2.11, Thm. 4.2, Prop. 4.4).
Fix the geometric notion of dominance for $X_{*}(\mathbf{A})$. Then there are bijections of sets
[…]
Moreover, for a simple $J$-module $E$,
  1. 1.
    $E$ is finite-dimensional over $\mathbb{C}$;
  2. 2.
    There is a unique two-sided cell $\mathbf{c}=\mathbf{c}(E)$ of $\tilde{W}$ such that $\mathrm{trace}\left({E}\,,{t_{w}}\right)\neq 0$ implies $w\in\mathbf{c}$.
  3. 3.
    $\mathrm{trace}\left({E}\,,{t_{w}}\right)$ is the constant term of the polynomial
    $(-\mathbf{q}^{1/2})^{a(\mathbf{c}(E))}\mathrm{trace}\left({M}\,,{C_{w}}\right)\in\mathbb{C}[\mathbf{q}^{\frac{1}{2}}]$
    where $M\simeq{}^{\phi}E$.
In particular, $\mathrm{trace}\left({E}\,,{t_{w}}\right)$ is independent of $\mathbf{q}$, and upon specializing $\mathbf{q}=q$ a prime power, will be a regular function in the twisting character in the setting of the Payley-Weiner theorem for the Iwahori-Hecke algebra of the $p$-adic group $G$, as we will explain in greater detail below.
That is, in our conventions, simple tempered $\mathbf{H}|_{\mathbf{q}=q}$-modules in the sense of harmonic analysis will extend to simple $J$-modules via ϕ()italic-$\phi\circ()^{\dagger}$.
We will comment even further on the necessity of twisting by the Goldman involution in Section 2.4.
1.6 Acknowledgements
The author thanks Alexander Braverman for introducing him to this material and the problem, for suggesting the use of the Plancherel formula, and for helpful conversations. The author also thanks Roman Bezrukavnikov, Itai Bar-Natan, Gurbir Dhillon, Malors Espinosa-Lara, Julia Gordon, George Lusztig, and Matthew Sunohara for helpful conversations. The author thanks Maarten Solleveld for explaining the generalization in [Solleveld] of the results of [fomFormalDegUnip], suggesting the reference [APM], and for careful reading of an early version of this manuscript.
Some of this work was completed at the Max Planck Institute in Mathematics for Bonn, the directors and staff of which the author thanks for their hospitality. This research was supported by NSERC.
2 Harish-Chandra’s Plancherel formula
We recall the notation and classical results we will need about the Plancherel formula. For Iwahori-biinvariant Schwartz functions, the Plancherel formula is known explicitly for all connected reductive groups, and is due to Opdam in [OpdamSpectral]. In the case of $G=\mathrm{GL}_{n}(F)$, we shall refer instead to [AubPlym] (where in fact the Plancherel formula is computed explicitly in its entirety for $\mathrm{GL}_{n}$). In the case $\mathbf{G}=\mathrm{Sp}_{4}$, we shall refer to the unpublished work [AKSp4] of Aubert and Kim. For $\mathbf{G}=G_{2}$, we will refer to Parkinson [Parkinson].
In this section $q$ is a prime power (or at least a real number of absolute value strictly greater than $1$). The formal variable $\mathbf{q}$ will not appear in this section.
2.1 Tempered and discrete series representations
Definition 7.
A smooth representation $\pi$ of $G$ is tempered if $\pi$ is unitary, of finite length, and all matrix coefficients of $\pi$ belong to $L^{2+\epsilon}(G/Z(G))$ for all $\epsilon>0$. We write $\mathcal{M}_{t}(G)$ for the category of tempered representations of $G$.
In particular, the central character of such a representation takes values in the circle group $\mathbb{T}\subset\mathbb{C}^{\times}$.
Definition 8.
A smooth finite-length representation $\omega$ of $G$ belongs to the discrete series if all matrix coefficients of $\omega$ are square-integrable modulo $Z(G)$. We write $\mathcal{E}_{2}(G)$ for the space of unitary discrete series, and $\mathcal{E}_{2}(G)^{I}$ for the space of unitary discrete series with nontrivial Iwahori-fixed vectors.
The tempered representations are built from the discrete series according to the following theorem of Harish-Chandra, as related in [Waldspurger, Prop. III.4.1].
Theorem 5 (Harish-Chandra).
Let $P$ be a parabolic subgroup of $G$ with Levi subgroup $M$, and let $\omega\in\mathcal{E}_{2}(M)$. Let $\nu$ be a unitary character of $M$. Then $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$ is a tempered representation. Every simple tempered representation is a direct summand of a representation of this form.
By induction we always understand normalized induction.
2.1.1 Formal degrees of discrete series representations
We will soon study the Plancherel decomposition $f=\sum_{M}f_{M}$ of the Schwartz function $f$ determined by an element of $J$ as explained in Section 2.4. As will be explained below, each function $f_{M}$ is given by an integral formula that involves several constants that depend on the Levi subgroup $M$, or are functions on the discrete series of $M$. These constants are rational functions of $q$, the most sensitive of which is the formal degree $d(\omega)$ of $\omega\in\mathcal{E}_{2}(M)^{I}$. Much is known about formal degrees for $I$-spherical $\omega$; the most general current result seems to be
Theorem 6 ([Solleveld], [fomFormalDegUnip], Theorem 5.1 (b), [Gross-Reeder] Proposition 4.1).
Let $\mathbf{G}$ be connected reductive over $F$. Let $\omega$ be any unipotent—in particular, any Iwahori-spherical—discrete series representation of $G=\mathbf{G}(F)$. Then $d(\omega)$ is a rational function of $q$, the numerator and denominator of which are products of factors of the form $q^{m/2}$ with $m\in\mathbb{Z}$ and $(q^{n}-1)$ with $n\in\mathbb{N}$. Moreover, there is a polynomial $\Delta_{G}$ depending only on $G$ and $F$ such that $\Delta_{G}d(\omega)$ is a polynomial in $q$.
This result is proven by first proving that the Hiraga-Ichino-Ikeda conjecture [HII] holds for unipotent discrete series representations. Note that [Solleveld], [fomFormalDegUnip] and [HII] all use the normalization of the Haar measure on $G$ defined in [HII]. This normalization gives in our setting $\mu_{\mathrm{HII}}(K)=q^{\dim\mathbf{G}}\#\mathbf{G}(\mathbb{F}_{q})$. Hence, noting that #$\#\mathbf{G}(\mathbb{F}_{q})=P_{\mathbf{G}/\mathbf{B}}(q)\cdot\#\mathbf{B}(\mathbb{F}_{q})$ and that, as $\mathbb{F}_{q}$ is perfect, #$\#B(\mathbb{F}_{q})$ is a polynomial in $q$, we have
$\mu_{I}=\frac{1}{q^{\dim\mathbf{G}}\#\mathbf{B}(\mathbb{F}_{q})}\mu_{\mathrm{HII}},$
and so this question of normalization cannot affect the denominators of $d(\omega)$, for any Levi subgroup.
In the Iwahori-spherical case, Opdam showed the above result in [OpdamSpectral, Proposition 3.27 (v)], although with less control over the possible factors appearing in the numerator and denominator of $d(\omega)$. We emphasize that op. cit. does not make the splitness assumption we allow ourselves.
Remark 4.
Proposition 4.1 of [Gross-Reeder] studies not the $\gamma$-factor we are interested in, but rather its quotient by the $\gamma$-factor for the Steinberg representation. However, accounting for the use of the Euler-Poincaré measure $\mu_{\mathrm{HII}}$ on $G$, and known formula for the formal degree of the Steinberg representation, one may recover our desired statement about formal degrees from the main theorem and equation (61) of [Gross-Reeder].
2.2 Harish-Chandra’s canonical measure
In this section, we recall the standard coordinates used in [AubPlym], and [Waldspurger]. We follow both references closely. These conventions differ slightly from the original [HCP]. Everything in this section is standard, but we include details because we do require explicit measures with which to compute. We state the below for general $\mathbf{G}$, for application to each standard Levi subgroup of $\mathbf{G}$.
2.2.1 Unramified characters
Let $X(\mathbf{G})=\mathrm{Hom}(\mathbf{G},{\mathbb{G}_{\mathrm{m}}})$ denote the rational characters of $\mathbf{G}$ defined over $F$. Let $\mathfrak{a}_{G}:=(X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{R})^{*}=(X^{*}(\mathbf{G})\otimes_{\mathbb{Z}}\mathbb{R})^{*}$ be the real Lie algebra of the maximal split central torus $\mathbf{A}=\mathbf{A}_{G}$ of $\mathbf{G}$ and let ${\mathfrak{a}_{G}}_{\mathbb{C}}$ be its complexification. We have a map
$X(\mathbf{G})\to\mathrm{Hom}(G/G^{1},\mathbb{C}^{\times})=:\mathcal{X}(G)$
given by $\chi\mapsto|\chi|_{F}$, where $|\chi|_{F}(g)=|\chi(g)|_{F}$ and $G^{1}=\bigcap_{\chi\in X(\mathbf{G})}\ker|\chi|_{F}$. This gives the unramified characters $\mathcal{X}(G)$ a complex manifold structure under which $\mathcal{X}(G)\simeq(\mathbb{C}^{\times})^{\dim_{\mathbb{R}}\mathfrak{a}_{G}}$. For indeed, we have
${{\mathfrak{a}_{G}}_{\mathbb{C}}^{*}}$${\mathcal{X}(G)}$11{1}1
given by
$\chi\otimes s\mapsto(g\mapsto|\chi(g)|_{F}^{s}=q^{-s\operatorname{\mathrm{val}}(\chi(g))}).$
The kernel is spanned by all $\chi\otimes s$ such that $s\operatorname{\mathrm{val}}(\chi(g))\in\frac{2\pi i}{\log q}\mathbb{Z}$ for all $g\in G$. Hence the kernel is $\frac{2\pi i}{\log q}R$, where $R\subset X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{Q}$ is a lattice. In this way the quotient $\mathcal{X}(G)$ is a complex manifold.
2.2.2 Unitary unramified characters
Denote $\mathrm{Im}\,\mathcal{X}(G)$ the set of unitary unramified characters taking values in the unit circle $\mathbb{T}\subset\mathbb{C}^{\times}$. This notation is justified, as if $|\chi|_{F}=|\chi^{\prime}|_{F}$, then $\mathrm{Re}(\chi)=\mathrm{Re}(\chi^{\prime})\in\mathfrak{a}_{G}$. Hence we can define $\mathrm{Im}\,\mathcal{X}(G)$ to be the unramified characters coming from pure imaginary elements of ${\mathfrak{a}_{G}}_{\mathbb{C}}$.
The surjection $\mathrm{Im}\,\mathcal{X}(G)\to\mathrm{Im}\,\mathcal{X}(A)$ has finite kernel, and $\mathrm{Im}\,\mathcal{X}(A)$ is compact. We chose the Haar measure on it with volume one.
2.2.3 Action by twisting and the canonical measure
The group $\mathcal{X}(G)$ acts on admissible representations of $G$ by twisting: $\omega\mapsto\omega\otimes\nu$ for $\nu\in\mathcal{X}(G)$. This restricts to an action the unitary unramified characters on $\mathcal{E}_{2}(G)$.
Pulling this action back to $i\mathfrak{a}_{G}$, and given a representation $\omega$, let $L^{*}$ be its stabilizer in $i\mathfrak{a}_{G}$, so that the orbit $\mathfrak{o}$ of $\omega$ is identified as $i{\mathfrak{a}_{G}}/L^{*}$. This gives $\mathfrak{o}=\mathrm{Im}\,\mathcal{X}(G)\cdot\omega$ the structure of a real submanifold of the larger orbit $\mathfrak{o}_{\mathbb{C}}={\mathfrak{a}_{G}}_{\mathbb{C}}/L^{*}=\mathcal{X}(G)\cdot\omega$.
Definition 9.
Given an orbit $\mathfrak{o}\subset\mathcal{E}_{2}(G)$, the Harish-Chandra canonical measure on $\mathfrak{o}$ is the Euclidean measure on $\mathfrak{o}$ whose pullback to $\mathrm{Im}\,\mathcal{X}(G)$ agrees with the pullback of the Haar measure on $\mathrm{Im}\,\mathcal{X}(A)$.
Hence even though the construction of the canonical measure is slightly involved, in practice it will be easy to recognize as being essentially the Haar measure on the compact torus $\mathrm{Im}\,\mathcal{X}(A)$.
Example 2.
If $G=\mathrm{SL}_{2}(F)$ and $M=A$ is the diagonal torus, we have $\mathfrak{a}_{G}^{*}\simeq\mathbb{R}$ and $R=\mathbb{Z}$ so that a fundamental domain for $\mathfrak{a}_{G}^{*}/\frac{2\pi}{\log q}R$ is $\left[-\frac{\pi}{\log q},\frac{\pi}{\log q}\right)$ and the canonical measure $d\nu=\frac{\log q}{2\pi}\mathrm{d}x$, where $\mathrm{d}x$ is the Lebesgue measure. To obtain quasicharacters of $G$, we associate to $\nu\in{\mathfrak{a}_{G}}_{\mathbb{C}}$ the quasicharacter $\chi_{\nu}(g)=q^{i\left\langle\nu,H_{G}(g)\right\rangle}$.
Therefore to compute the integral of a function $f$ on $\mathcal{E}_{2}(G)$ supported on the unramified unitary characters of $A$, we compute
$\int_{\mathcal{E}_{2}(A)}{f(\omega)}\,\mathrm{d}{\omega}=\int{f(\chi_{\nu})}\,\mathrm{d}{\nu}=\frac{\log q}{2\pi}\int_{\frac{-\pi}{\log q}}^{\frac{\pi}{\log q}}f(e^{it\log q})\,\mathrm{d}t=\frac{1}{2\pi i}\int_{\mathbb{T}}{\frac{f(z)}{z}}\,\mathrm{d}{z}$
if $t\mapsto e^{it\log q}=q^{s}=:z$ (here $s=it$) parameterizes the unit circle $\mathbb{T}$.
In general, $\mathcal{E}_{2}(M_{P})$ is a disjoint union of compact tori, and the Plancherel density descends to the quotients of these compact tori by certain finite groups, namely, the Weyl groups of $(P,A_{P})$. The set $\mathcal{E}_{2}(M_{P})^{I}$ is finite up to twist by unramified characters, by a result of Harish-Chandra [Waldspurger].
2.3 The Harish-Chandra Schwartz algebra
Let $\mathcal{C}=\mathcal{C}(G)$ be the Harish-Chandra Schwartz algebra of $G$; see [HCH], [Waldspurger], or [BK] for the definition and associated notation. In particular, we will record for future comparison with the argument in the proof of Proposition 2 that
$q^{\frac{\ell(w)}{2}}q^{-\#W}\leq\Delta(IwI)\leq q^{\frac{\ell(w)}{2}}.$
Let $\mathcal{C}^{I\times I}$ the subalgebra of Iwahori-biinvariant functions. As explained in [Waldspurger], the Fourier transform $f\mapsto\pi(f)$ defines an endomorphism of $\pi$ for every $f\in\mathcal{C}$ and every tempered representation $\pi$.
The Plancherel theorem is the statement (see [Waldspurger, Thm. VIII.1.1]) that this assignment defines an isomorphism of rings
$\mathcal{C}\to\mathcal{E}_{t}(G),$
where $\mathcal{E}_{t}(G)$ is the subring of the endomorphism ring of the forgetful functor $\mathcal{M}_{t}(G)\to\mathbf{Vect}_{\mathbb{C}}$ defined by the following conditions:
  1. 1.
    For all $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$, the endomorphism $\eta_{\pi}=\eta_{\nu,\omega}$ is a smooth function of the unramified unitary character $\nu$ and $\omega\in\mathcal{E}_{2}(M_{P})$;
  2. 2.
    The endomorphism $\eta_{\pi}$ is biinvariant with respect to some open compact subgroup of $G$.
We have the obvious inclusion $\iota\colon H\mathrel{\hookrightarrow}\mathcal{C}^{I\times I}$.
Define the subring $\mathcal{E}(G)$ of the endomorphism ring of the forgetful functor $\mathcal{M}(G)\to\mathbf{Vect}_{\mathbb{C}}$ from the category of all smooth representations, by replacing, in (1) above, $\mathcal{E}_{2}(M)$ with $\mathcal{M}(M)$, unitary characters with all unramified characters, and “smooth” with ”algebraic.” The matrix Paley-Wiener theorem of Bernstein [Bernstein] says that $f\mapsto\pi(f)$ is an isomorphism from the full Hecke algebra of $G$ onto $\mathcal{E}(G)$. Denote $\mathcal{E}^{I}$ and $\mathcal{E}_{t}$ the subrings of $I$-invariant endomorphisms.
For computational purposes such as ours, we require that these isomorphisms be explicit. In the Iwahori-spherical case, harmonic analysis on $\mathcal{C}^{I\times I}$ can be phrased internally to $H$ and various completions of $H$. In this setting Opdam gave an explicit Plancherel formula in [OpdamSpectral]. In more general settings there are explicit formulas for $\mathrm{GL}_{n}(F)$, $\mathrm{Sp}_{4}(F)$, and $G_{2}(F)$, which we will also make use of.
2.4 The algebra $J$ as a subalgebra of the Schwartz algebra.
In [BK], Braverman and Kazhdan constructed a map of $\mathbb{C}$-algebras $J\to\mathcal{C}^{I\times I}$. We shall review this construction now.
Definition 10 ([BK], Section 1.7).
Let $P=M_{P}N_{P}$. A character $\chi\colon M_{P}\to\mathbb{C}^{\times}$ of $M_{P}$ is non-strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x))|_{\infty}\geq 1$ for $|x|_{F}\geq 1$.
We say a non-strictly positive character $\chi$ is strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x)|_{\infty}>1$ for $|x|_{F}\geq 1$.
Of course, it suffices to test this for $x=\varpi^{-1}$. The condition for $\chi^{-1}$ to be non-strictly positive is then that $|\chi(\alpha^{\vee}(\varpi))|_{\infty}\geq 1$.
Example 3.
For $\mathbf{G}=\mathrm{SL}_{2}$, in the conventions fixed in Example 2, an unramified character $\chi$ of $A$ is non-strictly positive if $\chi^{-1}$ corresponds to $z$ such that $|z|\geq 1$.
If $\mathbf{G}=\mathrm{GL}_{n}$ and $\nu$ corresponds to the vector $(z_{1},\dots,z_{n})\in\mathbb{C}^{n}$, then the condition that $\nu^{-1}$ is non-strictly positive translate to $|z_{n}|\geq|z_{n-1}|\geq\cdots\geq|z_{1}|$. Such conditions divide $\mathbb{C}^{n}$ into chambers, on which the Weyl group $\mathfrak{S}_{n}$ clearly acts simply-transitively. Interior points correspond to $\nu$ such that $\nu^{-1}$ is strictly positive.
Following op. cit., let $\mathcal{E}^{I}_{J}(G)$ denote the subring of $\mathcal{E}_{t}(G)$ defined by the following conditions on the endomorphisms $\eta_{\pi}$:
  1. 1.
    For all $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$, the endomorphism $\eta_{\pi}=\eta_{\nu,\omega}$ is a rational function of $\nu$, regular on the set of $\nu$ such that $\nu^{-1}$ is non-strictly positive.
  2. 2.
    The endomorphism $\eta_{\pi}$ is $I\times I$-biinvariant.
Theorem 7 ([BK], Theorem 1.8).
Let $\mathbf{G}$ be a connected reductive group defined and split over $F$. Then the following statements hold:
  1. 1.
    Let $\pi$ be a tempered representation of $G$. Then the action of $H$ on $\pi^{I}$ extends uniquely to $J$.
  2. 2.
    Let $P=MN$ be a parabolic subgroup of $G$ with Levi subgroup $M$ and let $\omega$ an irreducible tempered representation of $M$. Let $\nu$ be a non-strictly positive character of $M$ and let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu^{-1})$. Then the action of $H(G,I)$ on $\pi^{I}$ extends uniquely to an action of $J$.
  3. 3.
    The map $t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}$ defines a homomorphism
    $\eta\colon J\to\mathcal{E}_{J}^{I}(G).$
Composing this isomorphism with the inverse Fourier transform, Braverman and Kazhdan define an algebra map
$\tilde{\phi}\colon J\to\mathcal{C}^{I\times I}$
sending
$t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}\mapsto f_{w}=\sum_{x\in\tilde{W}}A_{x,w}T_{x}\in\mathcal{C}^{I\times I},$
where $A_{x,w}=f_{w}(IxI)$. By definition, $\eta_{\pi}(w)=\pi(f_{w})$ as endomorphisms of $\pi$. We will show later that $\tilde{\phi}$ is essentially the map $\phi^{-1}$.
Remark 5.
There are gaps in the proofs of injectivity and, as pointed out to us by R. Bezrukavnikov and I. Karpov after an early version of the present paper was completed, of surjectivity of the map $\eta$ in [BK]. In the present paper we prove injectivity by proving injectivity of $\tilde{\phi}$ in Corollary 4. We prove surjectivity in [rigidDet] for all but a small number of cells for exceptional groups; [BKK] proves that $\eta$ is an isomorphism for all two-sided cells.
Implicit in [BK] is
Lemma 2.
We have the commutative diagram (2).
Proof.
Let $\pi$ be a tempered representation of $G$. By [BK], the $H$-action on it extends to a $J$-action, and by Corollary 2.6 of loc. cit., we have $\eta_{\pi}(\phi_{q}(f^{\dagger}))=\pi(f)$ in $\mathcal{E}_{t}^{I}$ for any $f\in H$. Therefore $\tilde{\phi}\circ\phi_{q}(f^{\dagger})=f$ by the Plancherel theorem. ∎
Remark 6.
The presence of the Goldman involution is necessary for tempered representations of $J$ to restrict to tempered representations of $H$. For example, in [BK], Braverman and Kazhdan showed for $G=\mathrm{SL}_{2}(F)$, that $\mathcal{E}^{I}_{J}(G)=\operatorname{End}(\mathrm{St}^{I})\oplus\operatorname{End}_{\Phi,X_{*}}(\mathcal{S}^{I})$, matching the decomposition $J=\mathbb{C}t_{1}\oplus J_{0}$, where $\Phi$ is the Fourier transform on the Schwartz space of the basic affine space $\mathcal{S}$ [BKBasicAffine], and $\mathrm{St}$ is the Steinberg representation. In this case $J$ has three one-dimensional representations, whereas $H$ has one one-dimensional discrete series representation $\mathrm{St}$, two tempered one-dimensional representations not belonging to the discrete series, and one non-tempered one-dimensional representation, the trivial representation. It is easy to compute using the formulas of [D] that $\phi_{q}(T_{s})t_{1}=qt_{1}$, whereas $\phi_{q}(T_{s}^{\dagger})t_{1}=-t_{1}$. Thus for the Steinberg representation to restrict to the Steinberg representation, the twist is required.
2.5 The Plancherel formula for $\mathrm{GL}_{n}$
For $G=\mathrm{GL}_{n}(F)$, we have access to an explicit Plancherel measure and its Bernstein decomposition, thanks to [AubPlym].
Recall that for $G=\mathrm{GL}_{n}(F)$, we have bijections
$\{\text{partitions}~{}\text{of}~{}n\}$$\leftrightarrow\{\text{Standard Levi subgroups}~{}M~{}\text{of}~{}\mathrm{GL}_{n}(F)\}$ (3)
$\leftrightarrow\left\{\text{unipotent conjugacy classes in}~{}\mathrm{GL}_{n}(\mathbb{C})\right\}$ (4)
$\leftrightarrow\mathcal{N}/\mathrm{GL}_{n}(\mathbb{C})$
$\leftrightarrow\{2-\text{sided cells}~{}\mathbf{c}~{}\text{in}~{}\tilde{W}\}$ (5)
$\leftrightarrow\{\text{direct summands}~{}J_{\mathbf{c}}~{}\text{of}~{}J\}$
where (3) $\leftrightarrow$ (4) sends a unipotent conjugacy class $u$ to the standard Levi $M$ such that a member of $u$ is distinguished in $M^{\vee}$, and (4) $\leftrightarrow$ (5) is Lusztig’s bijection from [affineIV].
Definition 11.
Let $u$ be a unipotent element of a semisimple group $S$ over the complex numbers. Then $u$ is distinguished in $S$ if $Z_{S}(u)$ contains no nontrivial torus.
Let $P=M_{P}N_{P}$ be a parabolic subgroup of $G$ and let $\mathcal{O}$ be an orbit in $\mathcal{E}_{2}(M_{P})$ under the action of the unitary unramified characters of $M_{P}$ as explained in Section 2.2. Write $W_{M_{P}}\subset W$ for the finite Weyl group of $(M_{P},A_{P})$. Let $\mathrm{Stab}_{W_{M}}(\mathcal{O})$ be the stabilizer of $\mathcal{O}$. Recall that a parabolic subgroup of $G$ is said to be semistandard if it contains $A$. Then the Plancherel decomposition reads
$f=\sum_{(P=M_{P}N_{P},\mathcal{O})/\text{association}}f_{M_{P},\mathcal{O}}$
where $f\in\mathcal{C}(G)$, the sum is taken over semistandard parabolic subgroups $P$ up to association, and
$f_{M_{P},\mathcal{O}}(g)=c(G/M)^{-2}\gamma(G/M)^{-1}\#\mathrm{Stab}_{W_{M}}(\mathcal{O})^{-1}\int_{\mathcal{O}}{\mu_{G/M}(\omega)d(\omega)\mathrm{trace}\left({\pi}\,,{R_{g}(f)}\right)}\,\mathrm{d}{\omega},$
where $R_{g}f(x)=f(xg)$ is the right translation of $f$ and $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$ is the normalized parabolic induction of the twist of $\omega$ by a unitary unramified character $\nu$ of $M$. In [AubPlym], each term above is explicitly calculated as a rational function of $q$.
Lemma 3.
There is a finite set $S=\left\{(M,\omega)\,\middle|\,\omega\in\mathcal{E}_{2}(M)^{I}\right\}$ such that $\mathrm{trace}\left({\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)}\,,{f}\right)$ is nonzero only for $\omega\in S$, for all $I$-biinvaraiant Schwartz functions $f$.
Proof.
This is entirely standard. As $\mathcal{E}_{2}(M)^{I_{M}}$ is finite for every $M$, and there are finitely-many standard parabolics of $G$, we need only show that $\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)^{I}\neq 0$ only if $\omega^{I_{M}}\neq 0$, where $I_{M}$ is the Iwahori subgroup of the reductive group $M$ relative to the Borel subgroup $M(\mathbb{F}_{q})\cap B(\mathbb{F}_{q})$ of $M(\mathbb{F}_{q})$. Note that $I_{M}$ is naturally a subgroup of $I$. For any representation $\sigma$ of $M$, if $f\in\operatorname{Ind}_{P}^{G}(\sigma)$ is $I$-fixed, then for $i_{M}\in I_{M}$, we have
$f(i_{M})=\sigma(i_{M})\delta^{\frac{1}{2}}_{P}(i_{M})f(1)=f(1),$
and so if $\delta^{\frac{1}{2}}_{P}(i_{M})=1$, then it must be that $f(1)\in\sigma^{I_{M}}$. The Iwahori factorization
$I_{M}=N_{P}^{-}(\pi\mathcal{O})A(\mathcal{O})N_{M}(\mathcal{O})$
and the fact that $\mathrm{Ad}_{\mathfrak{n}_{P}}$ is a representation implies that
$\delta^{\frac{1}{2}}_{P}(i_{M})=|\det\mathrm{Ad}_{\mathfrak{n}_{P}}(N_{P}^{-}(i_{M})A(i_{M})N_{M}(i_{M}))|_{F}=|\det\mathrm{Ad}_{\mathfrak{n}_{P}}(A(i_{M}))|_{F}=1$
by unipotence of the groups $N_{P}^{-}$ and $N_{M}$ and the fact that the image of $i_{M}$ in $M(\mathbb{F}_{q})$ lies in the Borel. ∎
Let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$ be a tempered representation and let $(u,s)$ be the KL-parameter of its discrete support. Then by [KLDeligneLanglands, Theorem 8.3], $\mathbf{M}_{\mathbf{P}}^{\vee}$ is minimal such that $(u,s)\in M_{P}^{\vee}$. By op. cit., this condition is equivalent to $Z_{M}(s)$ being semisimple and $u$ being distinguished in $Z_{M}(s)$.
Proposition 1 (c.f. [OpdamSpectral] Proposition 8.3).
The Plancherel decomposition is compatible with the decomposition $J=\bigoplus_{\mathbf{c}}J_{\mathbf{c}}$ in the sense that if $w\in\mathbf{c}$, $f=f_{w}$ and $u=u(\mathbf{c})$ under Lusztig’s bijection, then $f_{M}\neq 0$ only for those $M$ such that there exists $s\in M^{\vee}(\mathbb{C})$ such that $Z_{M^{\vee}}(s)$ is semisimple and $Z_{M^{\vee}}(s)\cap Z_{M^{\vee}}(u)$ contains no nontrivial torus.
Proof.
Let $\pi:=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$ be a tempered irreducible representation of $G$ induced as usual from a standard parabolic subgroup $P$ with Levi subgroup $M$. Then $\pi^{I}$ is a tempered irreducible $H$-module, and is of the form $K(u,s,\rho,q)$ for $u,s\in\mathbf{G}^{\vee}(\mathbb{C})$ and $\rho$ a representation of $\pi_{0}(Z_{G^{\vee}}(u,s))$ with $s$ compact. By [BK], Corollary 2.6, $K(u,s,\rho,q)$ extends to a $J$-module. By definition of Lusztig’s bijection, $\pi(f_{w})\neq 0$ only if $w$ is in the two-sided cell $\mathbf{c}=\mathbf{c}(u)$ of $\tilde{W}$ corresponding to $u$. On the other hand, $K(u,s,\rho,q)$ is induced from a square-integrable standard module $K_{M}(u,s,\tilde{\rho},q)$ of $H(M,I_{M})$. But now [KLDeligneLanglands, Theorem 8.3] says that $(u,s)$ must be preicsely as in the statement of the proposition. Thus only such summands $(f_{w})_{M}$ in the Plancherel decomposition of $f_{w}$ are nonzero. ∎
Of course, when $\mathbf{G}=\mathrm{GL}_{n}$, the bijections (3), (4), and (5), imply that there is a unique nonzero summand $(f_{w})_{M}$ for each $w$.
2.6 Plancherel measure for $\mathrm{GL}_{n}$
We refer to [AubPlym, Section 5], for a summary of the Bernstein decomposition of the tempered irreducible representations of $\mathrm{GL}_{n}$, in particular we use the description in loc. cit. of the Bernstein component parameterizing the $I$-spherical representations of $G$.
As we shall be applying the Plancherel formula only to Iwahori-biinvariant functions, it suffices to consider only irreducible tempered representations with Iwahori-fixed vectors. For $\mathrm{GL}_{n}$, the only such representations are of the form
$\pi=\operatorname{Ind}_{P}^{G}(\nu_{1}\mathrm{St}_{1}\boxtimes\cdots\boxtimes\nu_{k}\mathrm{St}_{k})$
where $\mathrm{St}_{i}$ is the Steinberg representation of $\mathrm{GL}_{i}$, and $P\supset M=\mathrm{GL}_{l_{1}}\times\cdots\times\mathrm{GL}_{l_{k}}$.
These representations are parameterized as follows. Let $M$ be a Levi subgroup corresponding to the partition $l_{1}+\cdots+l_{k}=n$, and recall that we write $\mathbb{T}$ for the circle group. Define $\gamma\in\mathfrak{S}_{n}$ by $\gamma=(1\dots l_{1})(1\dots l_{2})\cdots(1\dots l_{k})$, so that the fixed-point set $(\mathbb{T}^{n})^{\gamma}=\{(z_{1},\dots,z_{1},\dots,z_{k},\dots,z_{k})\}\simeq\mathbb{T}^{k}$. Then the irreducible tempered representations with Iwahori-fixed vectors induced from $M$ are parameterized by the compact orbifold $(\mathbb{T}^{n})^{\gamma}/Z_{\mathfrak{S}_{n}}(\gamma)$.
Theorem 8 ([AubPlym], Remark 5.6).
Let $G=\mathrm{GL}_{n}$ and $M=\mathrm{GL}_{l_{1}}\times\cdots\times\mathrm{GL}_{l_{k}}$ be a Levi subgroup. Then the Plancherel measure of $H$ on $(\mathbb{T}^{n})^{\gamma}/Z_{\mathfrak{S}_{n}}(\gamma)$ is
$\mathrm{d}\nu_{H}(\omega)=\prod_{i=1}^{k}\frac{q^{l_{i}^{2}-l_{i}}(q-1)^{l_{i}}}{l_{i}(q^{l_{i}}-1)}\cdot q^{\frac{n-n^{2}}{2}}\cdot\prod_{(i,j,g)}\frac{q^{2g+1}(z_{i}-z_{j}q^{g})(z_{i}-q^{-g}z_{j})}{(z_{i}-z_{j}q^{-g-1})(z_{i}-q^{g+1}z_{j})},$
where the tuples $(i,j,g)\in\mathbb{Z}\times\mathbb{Z}\times\frac{1}{2}\mathbb{Z}$ are tuples such that $1\leq i<j\leq k$ and $|g_{i}-g_{j}|\leq g\leq g_{i}+g_{j}$, where $g_{i}=\frac{l_{i}-1}{2}$.
This is the measure that we will integrate against, by successively applying the residue theorem. When carrying out explicit calculations, we will usually elide the constant
$\prod_{i=1}^{k}\frac{q^{l_{i}^{2}-l_{i}}(q-1)^{l_{i}}}{l_{i}(q^{l_{i}}-1)}\cdot q^{\frac{n-n^{2}}{2}}$
as it depends only on $M$. We shall abbreviate
$\Gamma_{i,j,g}:=q^{-2g-1}\left|\Gamma_{F}\left(q^{-g}\frac{z_{i}}{z_{j}}\right)\right|^{2}=\frac{(z_{i}-q^{g}z_{j})(z_{i}-q^{-g}z_{j})}{(z_{i}-z_{j}q^{-g-1})(z_{i}-q^{g+1}z_{j})},$
and recall that, as noted in the proof of Theorem 5.1 in [AubPlym], the function $(z_{1},\dots,z_{k})\mapsto\prod_{(i,j,g)}\Gamma_{i,j,g}$ is $Z_{\mathfrak{S}_{n}}(\gamma)$-invariant. Hence for the purposes of integration, we may allow ourselves to integrate simply over $\mathbb{T}^{k}$. Moreover, there are many cancellations between the $\Gamma_{i,j,g}$ for a fixed pair $i<j$ as $g$ varies. Indeed, putting $q_{ij}=q^{|g_{i}-g_{j}|}$, $q^{ij}=q^{g_{i}+g_{j}+1}$, and
$\Gamma^{ij}:=\frac{(z_{i}-q_{ij}z_{j})(z_{i}-(q_{ij})^{-1}z_{j})}{(z_{i}-q^{ij}z_{j})(z_{i}-(q^{ij})^{-1}z_{j})}$
we have
$\prod\Gamma_{i,j,g}=\Gamma^{ij},$
where the product is taken over all integers $g$ appearing in triples $(i,j,g)$ for $i<j$ fixed. We set
$c_{M}:=\prod_{(i,j,g)}q^{2g+1}\cdot\prod_{i=1}^{k}\frac{q^{l_{i}^{2}-l_{i}}(q-1)^{l_{i}}}{l_{i}(q^{l_{i}}-1)}\cdot q^{\frac{n-n^{2}}{2}},$
where the first product is taken over $(i,j,g)$ such that $1\leq i<j\leq k$ and $|g_{i}-g_{j}|\leq g\leq g_{i}+g_{j}$.
2.7 Beyond type $A$: the Plancherel formula following Opdam
Beyond type $A$, we still have available Opdam’s explicit Plancherel formula for the Iwahori-Hecke algebra [OpdamSpectral]. Let $\mathbf{G}$ be a connected reductive algebraic group defined and split over $F$, of Dynkin type other than type $A$ (to avoid redundancy). Let $f$ be an Iwahori-biinvariant Schwartz function on $G$ and let $M$ be a Levi subgroup of $G$. Given a parabolic subgroup $\mathbf{P}$ of $\mathbf{G}$, let $R_{1,+}$ and $R_{P,1,+}$ be defined as in [OpdamSpectral], Section 2.3. Recall that the group of unramified characters of a Levi subgroup $M$ has the structure of a complex torus, and is in fact a maximal torus of $\mathbf{M}^{\vee}(\mathbb{C})$. In particular, if $P$ is a parabolic subgroup and $\alpha$ is a root of $(M_{P},A_{P})$, then it makes sense to write $\alpha(\nu)$ for any unramified character $\nu$ of $M$. Then, altering Opdam’s notation to match our own from Section 2.5, the Plancherel formula reads
Theorem 9 ([OpdamSpectral] Thm. 4.43).
$f_{M,\mathcal{O}}(1)=\frac{q^{-\ell(w^{P})}}{\#\mathrm{Stab}_{W_{M}}(\mathcal{O})}\int_{\mathcal{O}}d(\omega)\prod_{\alpha^{\vee}\in R_{1,+}\setminus R_{P,1,+}}\frac{|1-\alpha^{\vee}(\nu)|^{2}}{|1+q^{\frac{1}{2}}_{\alpha}\alpha^{\vee}(\nu)^{1/2}|^{2}|1-q^{\frac{1}{2}}_{\alpha}q_{2\alpha}\alpha^{\vee}(\nu)^{1/2}|^{2}}\mathrm{trace}\left({\pi}\,,{f}\right)\,\mathrm{d}\omega,$ (6)
where $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$, $P\supset M$, and where $q_{\alpha}$ and $q_{2\alpha}$ are powers of $q$, and $w^{P}$ is the longest element in the complement $W^{P}$ to the parabolic subgroup $W_{P}$ of $W$.
Note that whenever $q_{2\alpha}=1$, which holds whenever $\alpha^{\vee}\not\in 2X_{*}$, the factor for $\alpha$ reduces to
$\frac{|1-\alpha^{\vee}(\nu)|^{2}}{|1-q_{\alpha}\alpha^{\vee}(\nu)|^{2}}.$ (7)
In types $A$ (as we have used above) and $D$, this simplification always occurs. In types $B$ and $C$, it happens for all roots except $\alpha=2\varepsilon_{i}\in R_{1,+}(B_{n})$ and $\alpha=4\varepsilon_{i}\in R_{1,+}(C_{n})$, where $\varepsilon_{i}$ is the character $\mathrm{diag}(a,\dots,a_{n})\mapsto a_{i}$.
For explicit evaluation we rewrite (6) in coordinates as follows. Recalling the setup of Section 2.2.1, a chose basis $\{\beta_{i}^{\vee}\}$ of the coweight lattice of $\mathbf{G}$. Then we obtain coordinates $z_{i}$, such that if $\alpha^{\vee}=\sum_{i}e_{i}\beta_{i}^{\vee}$, then the factor in (6) labelled by $\alpha^{\vee}$ is
$\frac{|1-z_{1}^{e_{1}}\cdots z_{n}^{e_{n}}|^{2}}{|1+q^{\frac{1}{2}}_{\alpha}(z_{1}^{e_{1}}\cdots z_{n}^{e_{n}})^{1/2}|^{2}|1-q^{\frac{1}{2}}_{\alpha}q_{2\alpha}(z_{1}^{e_{1}}\cdots z_{n}^{e_{n}})^{1/2}|^{2}}.$ (8)
When integrating, the coordinates $z_{i}$ are restricted to the residual coset corresponding to $\mathcal{O}$, in the sense of [OpdamSpectral]. For $\mathbf{G}=\mathrm{GL}_{n}$, we used the basis afforded by the characters $\varepsilon_{i}$.
2.8 Regularity of the trace
In order to extract information about the expansion of the elements $t_{w}$ in terms of the basis $T_{x}$ via the Plancherel formula, we must establish a regularity property of $\mathrm{trace}\left({\pi}\,,{f_{w}}\right)$ where $\pi$ is an irreducible tempered representation of $G$. The needed property follows trivially from Theorem 7.
2.8.1 Intertwining operators
The goal is to use the property that elements of $\mathcal{E}^{I}_{J}(G)$ commute with all intertwining operators in $\mathcal{M}_{t}(G)$, and regularity of the trace for unitary and non-strictly positive characters of $M$ to deduce regularity of the trace at all characters of $M$.
Let $\omega$ be a discrete series representation of a Levi subgroup $M$ of $G$, and let $\nu$ be any unramified character of $M$, not necessarily unitary. Then we may form the representation $\pi=\operatorname{Ind}_{P}^{B}(\nu\otimes\omega)$ of $G$, where $P$ is a parabolic with Levi factor $M$. We will now recall some well-known facts about the action of the Weyl group of $M$ on such representations $\pi$. Let $\theta$ and $\theta^{\prime}$ be two subsets of $\Delta$ corresponding to Levi subgroups $M$ and $M^{\prime}$. Let $w\in W$ be such that $w\theta=\theta^{\prime}$. Then there is an intertwining operator
$J_{P|P^{\prime}}(\omega,\nu)\colon\operatorname{Ind}_{P}^{B}(\nu\otimes\omega)\to\operatorname{Ind}_{P^{\prime}}^{B}(\nu\otimes\omega)$
for each $w\in W(\theta,\theta^{\prime})=\left\{w\in W\,\middle|\,w(\theta)=\theta^{\prime}\right\}$. For $\mathbf{G}=\mathrm{GL}_{n}$, this set is nonempty only if
$M_{\theta}=\mathrm{GL}_{l_{1}}\times\cdots\times\mathrm{GL}_{l_{N}}~{}~{}M_{\theta^{\prime}}=\mathrm{GL}_{l^{\prime}_{1}}\times\cdots\times\mathrm{GL}_{l^{\prime}_{N}}$
and $\{l_{1},\dots,l_{N}\}=\{l_{1}^{\prime}\,\dots,l^{\prime}_{N}\}$ are equal multisets. In this case , $W(\theta,\theta^{\prime})\simeq\mathfrak{S}_{N}$ can be viewed as acting by permuting the blocks of $M$. It is well-known that $J_{P|P^{\prime}}(\omega,\nu)$ is a meromorphic function of $\nu$ with simple poles. The poles of these operators have been studied by Shahidi in [Shahidi], and in the language of modules over the full Hecke algebra by Arthur in [Art]. The results will be stated for certain renormalizations $A(\nu,\omega,w)$ of the operators $J_{P|P^{\prime}}(\omega,\nu)$ as explained in equation (2.2.1) of [Shahidi].
2.8.2 Conventions on parabolic subgroups
We now recall the notation of Shahidi [Shahidi]. Given a subset $\theta$, we set $\Sigma_{\theta}=\operatorname{span}_{\mathbb{R}}\theta$, and $\Sigma_{\theta}^{+}=\Psi^{+}\cap\Sigma_{\theta}$ and likewise define $\Sigma_{\theta}^{-}$. Let $\Sigma(\theta)$ be the roots of $(P,A_{P})$. Define the positive roots $\Sigma^{+}(\theta)$ to be the roots obtained by restriction of an element of $\Psi^{+}\setminus\Sigma_{\theta}^{-}$.
Given two subsets $\theta,\theta^{\prime}\subset\Delta$, following [Shahidi] we set
$W(\theta,\theta^{\prime})=\left\{w\in W\,\middle|\,w(\theta)=\theta^{\prime}\right\},$
and then for $w\in W(\theta,\theta^{\prime})$, we define
$\Sigma(\theta,\theta^{\prime},w)=\left\{[\beta]\in\Sigma^{+}(\theta)\,\middle|\,\beta\in\Psi^{+}-\Sigma_{\theta}^{+}~{}\text{and}~{}w(\beta)\in\Psi^{-}\right\},$
and then
$\Sigma^{\circ}(\theta,\theta^{\prime},w):=\left\{[\beta]\in\Sigma(\theta,\theta^{\prime},w)\,\middle|\,w_{[\beta]}\in W(A_{P})\right\}.$
Remark 7.
In [Shahidi] additional care about the relative case is taken in the notation. In our simple case this is of course unnecessary, and we omit it.
In the case $\mathbf{G}=\mathrm{GL}_{n}$, this specializes as follows. Let $\alpha_{ij}\colon\mathrm{diag}(t_{i})\mapsto t_{i}t_{j}^{-1}$ be characters of $T$. The $\alpha_{ij}$ for $j>i$ are the positive roots of $(\mathbf{B},\mathbf{A})$ and $\beta_{i}:=\alpha_{ii+1}$ are our chosen simple roots. If $P$ corresponds to the partition $n_{1}+\cdots+n_{p}$ of $n$ and subset $\theta\subset\Delta$, then
$\Sigma(\theta)=\operatorname{span}{\{\beta_{n_{1}},\beta_{n_{1}+n_{2}},\dots\}},$
where we view $\beta_{i}$ as restricted to $\mathfrak{a}_{P}\mathrel{\hookrightarrow}\mathfrak{a}_{G}$. Note that all the positive roots $\Sigma^{+}(\theta)$ of $(P_{\theta},N_{\theta})$ are in $N_{P}$. Denoting by $[\alpha]$ the coset representing a root $\alpha$ of $G$ restricted to $P$, the positive roots in $N_{P}$ are the $\alpha_{ij}$ such that $[\alpha_{ij}]=[\beta_{n_{1}+\cdots+n_{k}}]$.
Example 4.
If $\mathbf{G}=\mathrm{GL}_{6}$ and $P$ be the parabolic subgroup of block upper-triangular matrices corresponding to the partition $6=2+2+1+1$ and $\theta=\{\alpha_{12},\alpha_{34}\}$, then we have
$A_{P}=\left\{\mathrm{diag}(t_{1},t_{1},t_{2},t_{2},t_{3},t_{4})\right\}.$
The positive simple roots are $\Sigma^{+}(\theta)=\{[\beta_{2}],[\beta_{4}],[\beta_{5}]\}$. The Weyl group $W(A_{P})\simeq\mathfrak{S}_{2}\times\mathfrak{S}_{2}$ acts by permuting the blocks. Note that the simple reflection sending $\beta_{4}\mapsto-\beta_{4}$ does not arise by permuting the blocks (i.e. $\begin{pmatrix}4&5\end{pmatrix}$ does not send blocks to blocks) , hence $w_{[\beta_{4}]}\not\in W(A_{P})$. Hence for any $w\in W(\theta,\theta^{\prime})$ we have $\Sigma^{\circ}(\theta,\theta^{\prime},w)\subseteq\left\{[\beta_{2}],[\beta_{5}]\right\}$.
2.8.3 Regularity of the trace
We have the following information about the poles of intertwining operators, due, according to [Shahidi], to Harish-Chandra:
Theorem 10 ([Shahidi], Theorem 2.2.1).
Let $\omega$ be an irreducible unitary representation of $M$. Say that $\omega$ is a subrepresentation of $\operatorname{Ind}_{P_{*}}^{M}(\omega_{*})$ for a parabolic subgroup $P_{*}=M_{*}N_{*}$ and $\omega_{*}$ is an irreducible supercuspidal representation of $M_{*}$. Let $\theta_{*}\subset\Delta$ be such that $P_{*}=P_{\theta_{*}}$ as parabolic subgroups of $M_{*}$.
Then the operator
$\prod_{\alpha\in\Sigma^{0}_{r}(\theta_{*},w\theta_{*},w)}(1-\chi_{\omega,\nu}^{2}(h_{\alpha}))A(\nu,\omega,w)$
is holomorphic on $\mathfrak{a}_{\theta\mathbb{C}^{*}}$. Here $\chi_{\omega,\nu}$ is the central character of the twisted representation $\omega\otimes q^{\left\langle\nu,H_{\theta}(-)\right\rangle}$.
In particular for the purposes of the Plancherel formula, the only relevant $\omega$ are unitary, hence have unitary central characters. Therefore $A(\nu,\omega,w)$ is holomorphic at $\nu$ if $|q^{\left\langle\nu,H_{\theta}(-)\right\rangle}|\neq 1$, or equivalently if
$\Re(\left\langle\nu,H_{\theta}(-)\right\rangle)\neq 0.$
In particular, there is a finite union of hyperplanes away from which each operator $A(\nu,\pi,w)$ is holomorphic, for any $w\in W$.
Lemma 4.
Let $M$ be a Levi subgroup of $G=\mathbf{G}(F)$ and let $\omega$ be a discrete series representation of $M$. Let $P=M_{P}A_{P}N_{P}$ be the standard parabolic subgroup containing $M=M_{P}$ and let $k=\operatorname{rk}A_{P}$. Let $z_{1},\dots,z_{k}\in(\mathbb{C}^{\times})^{k}=X^{*}(A_{P})\otimes_{\mathbb{Z}}\mathbb{C}$ define an unramified quasicharacter $\nu=\nu(z_{1},\dots,z_{k})$ of $A_{P}$ as in Section 2.2. Let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$. Let $f\in J$. Then
$\mathrm{trace}\left({\pi}\,,{f}\right)\in\mathbb{C}[z_{1},\dots,z_{k},z_{1}^{-1},\dots,z_{k}^{-1}].$
That is, the trace is a regular function on (×$(\mathbb{C}^{\times})^{k}$.
Proof.
We know a priori that
$\mathrm{trace}\left({\pi}\,,{f}\right)\in\mathbb{C}(z_{1},\dots,z_{k})$
is a rational function of $\nu$, as the operator $\pi(f)$ itself depends rationally on the variables $z_{i}$ by Theorem 7. Therefore
$\mathrm{trace}\left({\pi}\,,{f}\right)=\frac{p(\nu)}{h(\nu)}\in\mathbb{C}(z_{1},\dots,z_{k}).$
By Theorem 10 and the discussion following it, there is an open subset $U$ of the unitary characters, such that for all $\nu\in U$, and all $w\in W$ we have
$\frac{p(\nu)}{h(\nu)}=\frac{p(w(\nu))}{h(w(\nu))}.$ (9)
Holding $z_{2},\dots,z_{n}$ constant and in $U$, (9) becomes an equality of meromorphic functions of $z_{1}$ that holds on a set with an accumulation point, and hence (9) holds for all $z_{1}\in\mathbb{C}^{\times}$. Now holding $z_{1}\in\mathbb{C}^{\times}$ constant and arbitrary, and $z_{3},\dots,z_{k}$ constant and in $U$, we see that (9) holds also for all $z_{2}\in\mathbb{C}^{\times}$. Therefore (9) actually holds for all $\nu$, i.e. $\mathrm{trace}\left({\pi}\,,{f}\right)$ is a $W$-invariant rational function of $\nu$.
When $\nu$ is non-strictly positive with respect to $M$, by Theorem 7, $\mathrm{trace}\left({\pi}\,,{f}\right)$ has poles only of the form $z_{i}^{n_{i}}=0$. The claim now follows from the $W$-invariance, and thus $\mathrm{trace}\left({\pi}\,,{f}\right)$ is a regular function on (×$(\mathbb{C}^{\times})^{k}$. ∎
We will therefore allow ourselves to write $\mathrm{trace}\left({\pi}\,,{f}\right)$ for functions $f\in J$ even for $\nu$ such that the operator $\pi(f)$ itself is not defined.
Lemma 5.
Let $d\in\tilde{W}$ be a distinguished involution and $\pi$ be a tempered representation induced from one of the Levi subgroups attached to the two-sided cell containing $d$ in the sense of Proposition 1. Then $\mathrm{trace}\left({\pi}\,,{f_{d}}\right)$ is constant and a natural number. In fact, the same is true for any idempotent in $J$.
Proof.
Let $j$ be an idempotent in $J$. We have $\mathrm{trace}\left({\pi}\,,{f_{j}}\right)=\operatorname{rank}(\pi(f_{j}))$. The trace is continuous in $\nu$ by Lemma 4, and the present lemma follows as $\mathbb{T}$ is connected. ∎
3 Proof of Theorem 13 for $\mathrm{GL}_{n}$, and of Theorem 2 for general $\mathbf{G}$
3.1 The functions $f_{w}$ for $\mathrm{GL}_{n}$
In this section, $q>1$.
To compute with the Plancherel formula, we will need to apply the residue theorem successively in each variables $z_{i}$, and in doing so will we will need to sum over a certain tree that will track, for each variable, at which residues we evaluated. Upon integrating with respect to each variable $z_{i}$, we will have poles of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$. For example, if we have $4$ variables $z_{1},z_{2},z_{3},z_{4}$ corresponding to a Levi subgroup $\mathrm{GL}_{l_{1}}\times\mathrm{GL}_{l_{2}}\times\mathrm{GL}_{l_{3}}\times\mathrm{GL}_{l_{4}}$, then some of the summands obtained by successively applying the residue theorem are labelled by paths on the tree
{forest}
[$z_{1}=0$ [$z_{2}=0$ [$z_{3}=0$] [$z_{3}=(q^{34})^{-1}z_{4}$] ] [$z_{2}=(q^{23})^{-1}z_{3}$ [$z_{3}=0$] [$z_{3}=(q^{34})^{-1}z_{4}$] ] [$z_{2}=(q^{24})^{-1}z_{4}$ [$z_{4}=0$] [$z_{4}=(q^{43})^{-1}z_{3}$] ] ]
Of course, to evaluate the entire integral for $M$, we would also need to consider trees whose roots are decorated with $z_{1}=(q^{12})^{-1}z_{2}$, and so on, for a total of four trees.
Definition 12.
Given a Levi subgroup $M$ with $N+1$ blocks, a bookkeeping tree $T$ for $M$ is a rooted tree with $N$ levels such that the vertices on the $i$-th level below the root each have $N-i$ child vertices, and each vertex is decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$, where the index $j$ does not appear along the path from the vertex to the root, and the parent root is decorated with an equation $z_{k}=(q^{ki})^{-1}z_{i}$ for some $k$. Moreover, we require that the root be decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$ for $i$ minimal. A branch of $T$ is a simple path in $T$ from the root to one of the leaves.
Definition 13.
Given a branch $B$ of a bookkeeping tree, a clump in $B$ is an ordered subset of indices $i$ appearing in the decorations of successive parent-child vertices, such that no decoration of the form $z_{i}=0$ occurs along the path from the closest index to the root to the farthest index from the root. We write $C\prec B$ if $C$ is a clump of $B$.
Example 5.
The sets of indices $\{3,4\}$, $\{2,3,4\}$, $\{2,3\}$, $\{2,4\}$, and $\{2,4,3\}$ (note the ordering) are all the clumps of the above tree. The sets $\{1,2,3\}$ and $\{1,2,4\}$ are not clumps.
Theorem 11.
Let $G=\mathrm{GL}_{n}(F)$ and let $d$ be a distinguished involution such that the two-sided cell containing $d$ corresponds to the Levi subgroup $M$. Let $N+1$ be the number of blocks in $M$ such that the $i$-th block has size $l_{i}$. Let $m_{j}$ be the number $l_{i}$ that are equal to $j$. Define for $r\leq k$
$Q_{rk}=q^{i_{k}i_{k+1}}q^{i_{k-1}i_{k}}\cdots q^{i_{r}i_{r+1}}$
in the notation of 2.6. Then if $f_{d}$ is as in Sections 2.4,
$f_{d}(1)=\frac{\operatorname{rank}(\pi(f_{d}))}{m_{1}!\cdots m_{n}!}c_{M}\sum_{\mathrm{trees}~{}T}\sum_{\mathrm{branches}~{}B~{}\mathrm{of}~{}T}\prod_{\begin{subarray}{c}C\prec B\\ C=\{i_{0},\dots,i_{t}\}\end{subarray}}\frac{(1-q^{l_{i_{0}}})(1-q^{l_{i_{1}}})}{1-q^{l_{i_{0}}+l_{i_{1}}}}\\ \cdot\prod_{k=1}^{t-1}\frac{(1-q^{l_{i_{k+1}}})}{(1-q^{l_{i_{k}}+l_{i_{k+1}}})}\prod_{r=0}^{k-1}\frac{R_{rk}}{1-Q_{rk}q^{i_{r}i_{k+1}}},$ (10)
where
$R_{rk}=\begin{cases}1-Q_{rk}q^{g_{i_{r}}-g_{i_{k}}}&\text{if}~{}k<t-1\\ (1-Q_{r,t-1}q^{g_{i_{r}}-g_{i_{t}-1}})(1-Q_{r,t-1}q^{g_{i_{k}}-g_{i_{r}}})&\text{if}~{}k=t-1\end{cases}.$
Corollary 1.
The denominator of $f_{d}(1)$ divides a power of the Poincaré polynomial $P_{\mathbf{G}/\mathbf{B}}(q)=P_{\mathfrak{S}_{n}}(q)$ of $\mathbf{G}$. Moreover, when $d$ is in lowest two-sided cell, corresponding to $M=T$, $f_{d}(1)=\operatorname{rank}(\pi(f_{d}))/P_{\mathbf{G}/\mathbf{B}}(q)$ exactly.
Proof of Corollary 1.
We will show that each of the three forms of denominator that appear in the conclusion of Theorem 11 divide $P_{\mathbf{G}/\mathbf{B}}(q)$, and thus that their product divides a power of $P_{\mathbf{G}/\mathbf{B}}(q)$. The denominators of $c_{M}$ are all of the form $1+q+\cdots+q^{l_{i}-1}$, and so divide $P_{\mathbf{G}/\mathbf{B}}(q)$ as $l_{i}\leq n$ for all $i$. Note that as $l_{i_{i}}+l_{i_{k+1}}\leq n$, the leftmost denominators in (10) satisfy the conclusion of the corollary also. Finally, $Q_{rk}q^{i_{r}i_{k+1}}=q^{l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}}$, and $R_{rk}$ is likewise always a polynomial in $q$ (as opposed to $q^{-1}$) divisible by $1-q$. Again using that $l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}\leq n$, we are done with the first statement.
We now take up the second statement, the proof [^2] We thank A. Braverman for explaining to us the observation whose obvious generalization we present below. of which does not require any computations at all and which holds for any $w\in\mathbf{c}_{0}$. Recalling that the only $K$-spherical tempered representations of $G$ are principal series representations [MacdonaldZonal], let $\mu_{K}$ be the Haar measure on $G$ such that $\mu_{K}(K)=1$, $\mu_{I}$ be the Haar measure such that $\mu_{I}(I)=1$, and denote temporarily $\mathrm{d}\pi_{\mu}$ the Plancherel measure normalized according to a chosen Haar measure $\mu$ on $G$. Likewise write $\pi_{K}(f)$ for the Fourier transform with respect to $\mu_{K}$ and $\pi_{I}(f)$ for the Fourier transform with respect to $\mu_{I}$.
As we have
$\mu_{I}=P_{\mathbf{G}/\mathbf{B}}(q)\mu_{K},$
and formal degrees scale inversely to the Haar measure, we have
$d\pi_{\mu_{I}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}d\pi_{\mu_{K}}$ (11)
for tempered principal series. Now let $w\in\mathbf{c}_{0}$. By Lemma 4 and the Satake isomorphism, there is a function $h_{w}$ in the spherical Hecke algebra such that for all principal series $\pi$,
$\mathrm{trace}\left(\pi_{I}(f_{w})\right)=\mathrm{trace}\left(\pi_{K}(h_{w})\right)$
as regular Weyl-invariant functions of the Satake parameter. By the Plancherel formula and (11),
$f_{w}(1)=\int{\mathrm{trace}\left(\pi_{I}(f_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{K}}}=\frac{h_{w}(1)}{P_{\mathbf{G}/\mathbf{B}}(q)}.$ (12)
In particular, if $w=d$ is a distinguished involution in the lowest two-sided cell, then
$\mathrm{trace}\left(\pi_{I}(f_{d})\right)=\operatorname{rank}\left(\pi_{I}(f_{d})\right)=\operatorname{rank}\left(\pi_{I}(f_{d})\right)\cdot\mathrm{trace}\left(\pi_{K}(1_{K})\right),$
and (12) becomes
$f_{d}(1)=\frac{\operatorname{rank}(\pi(f_{d}))}{P_{\mathbf{G}/\mathbf{B}}(q)}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}.$ (13)
Here the rank is given by [XiI], Proposition 5.5. (In op. cit. there is the assumption of simple-connectedness, but it is easy to see that the distinguished involutions for the extended affine Weyl group $\tilde{W}(\tilde{\mathbf{G}}(F))$ of the universal cover $\tilde{\mathbf{G}}$ are distinguished involutions for $\tilde{W}$ using the definition in [affineII] and uniqueness of the $\{C_{w}\}$-basis, and that the lowest cell is just the lowest cell of $\tilde{W}(\tilde{\mathbf{G}}(F))$ intersected with $\tilde{W}$.) ∎
Once we have established injectivity of $\tilde{\phi}$, we will show by a counting argument that $\operatorname{rank}\left(\pi(t_{d})\right)=1$ for any distinguished involution $d$, in the case $G=\mathrm{GL}_{n}$, see Theorem 13. Note that (13) is an example of the behaviour conjectured in Remark 1.
Remark 8.
It would be interesting to find $I$-biinvariant Schwartz functions $h$ playing the role of $h_{w}$ for the other two-sided cells, namely such that $\mathrm{trace}\left({\pi}\,,{h}\right)$ was regular, nonzero only for a single pair $(M,\omega)$, and the value $h(1)$ was known as a function of $q^{1/2}$.
Remark 9.
If $\mathcal{P}$ is a maximal parahoric subgroup of $G$, then the longest word $w_{\mathcal{P}}$ is a distinguished involution in the lowest two-sided cell. As above, and as we will explain again in Section 4, $\pi(f_{w_{\mathcal{P}}})$ is nonzero only for the principal series representations, and its image is contained in $\pi^{\mathcal{P}}$. Thus after we will have shown injectivity, [KeysClassI] gives another proof that $\operatorname{rank}(\pi(f_{w_{\mathcal{P}}}))=1$, and more generally that $\pi(t_{d})$ has rank 1 on the principal series for any distinguished involution $d\in\mathbf{c}_{0}$.
Lemma 6.
Let $e_{0},\dots,e_{n}\in\mathbb{Z}$. Then
$\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}z_{0}^{e_{0}}\cdots z_{N}^{e_{N}}\prod_{i<j}\Gamma^{ij}\mathrm{d}z_{0}\cdots\mathrm{d}z_{N}=0$ (14)
unless $e_{0}+\cdots+e_{N}=-N$.
Although we do not use the Lemma in the sequel, we include it because it illustrates an efficient way to compute the functions $f_{w}$ in practice, as we explain after its proof.
Corollary 2.
Let $w\in\tilde{W}$. Then $f_{w}(1)$ is a rational function of $q$ with denominator dividing a power of the Poincaré polynomial of $\mathbf{G}$. The numerator is a Laurent polynomial $p_{1}(q^{1/2})+p_{2}(q^{-1/2})$ in $q^{1/2}$, where the degree of $p_{1}$ is bounded uniformly in terms of $\tilde{W}$. The denominator of $f_{w}(1)$ depends only on the two-sided cell containing $w$.
Remark 10.
In light of Lemma 4, Lemma 6 and Corollary 2 have the following interpretation. Let $\Gamma$ be a left cell in a two-sided cell $\mathbf{c}$. Then in [XiAMemoir], Xi shows that all the rings $J_{\Gamma\cap\Gamma^{-1}}$ are isomorphic to the representation ring of the associated Levi subgroup $M_{\mathbf{c}}$, and $J_{\mathbf{c}}$ is a matrix algebra over $J_{\Gamma\cap\Gamma^{-1}}$. Therefore $w\in\Gamma\cap\Gamma^{-1}$ are labelled by dominant weights of $M_{\mathbf{c}}$, and if $t_{\lambda}$ is such an element, Xi’s results show that if $\mathbf{q}=q$ and $\pi=\operatorname{Ind}_{B}^{G}(\nu)$ is an irreducible representation of $J_{\mathbf{c}}$, then
$\mathrm{trace}\left({\pi}\,,{t_{\lambda}}\right)=\mathrm{trace}\left({V(\lambda)}\,,{\nu}\right),$
where we view $\nu$ as a semisimple conjugacy class in $M_{\mathbf{c}}$, and $V(\lambda)$ is the irreducible representation of $M_{\mathbf{c}}$ of highest weight $\lambda$. Then we have that $f_{\lambda}(1)\neq 0$ only if $\lambda$ is of height 00 with respect to the basis $\varepsilon_{i}\colon\mathrm{diag}(a_{1},\dots,a_{n})\mapsto a_{i}$.
The proofs of Lemma 6 and Corollary 2 will use the notation of the proof of Theorem 11, and we defer them until after the proof of the theorem.
3.1.1 Example computations and a less singular cell
In this section we provide two example computations to elucidate the coming proof of Theorem 11.
Example 6.
Let $G=\mathrm{GL}_{2}$ and $M=A$. Then $l_{1}=l_{2}=1$ and $g_{1}=g_{2}=0$. Let $d=s_{0}$ or $s_{1}$. Then
$f_{d}(1)$ $=\frac{1}{2}\frac{1}{2\pi i}\frac{1}{2\pi i}q^{-1}\int_{\mathbb{T}}\int_{\mathbb{T}}q\frac{(z_{1}-z_{2})(z_{1}-z_{2})}{(z_{1}-q^{-1}z_{2})(z_{1}-qz_{2})}\frac{\mathrm{d}z_{1}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}$
$=\frac{1}{2}\frac{1}{2\pi i}\int_{\mathbb{T}}\mathrm{Res}_{z_{1}=q^{-1}z_{2}}\frac{(z_{1}-z_{2})(z_{1}-z_{2})}{(z_{1}-q^{-1}z_{2})(z_{1}-qz_{2})}\frac{1}{z_{1}}+\mathrm{Res}_{z_{1}=0}\frac{(z_{1}-z_{2})(z_{1}-z_{2})}{(z_{1}-q^{-1}z_{2})(z_{1}-qz_{2})}\frac{1}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}$
$=\frac{1}{2}\frac{1}{2\pi i}\int_{\mathbb{T}}\frac{(q^{-1}z_{2}-z_{2})(q^{-1}z_{2}-z_{2})}{(q^{-1}z_{2}-qz_{2})q^{-1}z_{2}}+\frac{z_{2}^{2}}{z_{2}^{2}}\frac{\mathrm{d}z_{2}}{z_{2}}$
$=\frac{1}{2}\frac{1}{2\pi i}\int_{\mathbb{T}}\frac{(q^{-1}-1)^{2}}{q^{-2}-1}+1\frac{\mathrm{d}z_{2}}{z_{2}}$
$=\frac{1}{q+1}.$
The factor $\frac{1}{2}$ reflects the fact that we integrate with the respect to the pushforward of the above $\mathfrak{S}_{2}$-invariant measure to the quotient $\mathbb{T}\times\mathbb{T}/\mathfrak{S}_{2}$.
This agrees with the theorem, which instructs us to calculate $f_{d}(1)$ as follows: There are two trees, each of which has one vertex and no edges. The trees are $z_{1}=0$ and $z_{1}=q^{-1}z_{2}$. Each has one branch. The first has no clumps, so the entire is product is empty. The second tree has one clump $C=\{1,2\}$ for which $t=1$, and we obtain
$f_{d}(1)=\frac{1}{2}\left(1+\frac{1-q}{1+q}\right)=\frac{1}{2}\frac{2}{1+q}=\frac{1}{1+q}.$
Now we give an example of a less singular cell.
Example 7.
Let $G=\mathrm{GL}_{4}$, and let $\mathbf{c}$ be the two-sided cell corresponding to the partition $4=2+2$. Then $P_{W}(\mathbf{q})=(1+\mathbf{q})(1+\mathbf{q}+\mathbf{q}^{2})(1+\mathbf{q}+\mathbf{q}^{2}+\mathbf{q}^{3})$ has five distinct roots $\mathbf{q}=-1,\pm i,\zeta_{1},\zeta_{2}$, where $\zeta_{i}$ are primitive third roots of unity. We compute $P_{\mathbf{c}}$ using the Plancherel formula. We have
We have $l_{1}=l_{2}=2$, $q_{12}=1$, $q^{12}=q^{2}$,
$c_{M}=\frac{q^{4}(q-1)^{2}}{2(q+1)^{2}},$
and
$\left(\frac{1}{2\pi i}\right)^{2}\iint_{\mathbb{T}^{2}}\Gamma^{12}\frac{\mathrm{d}z_{1}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}$ $=\left(\frac{1}{2\pi i}\right)^{2}\iint_{\mathbb{T}^{2}}\frac{(z_{1}-z_{2})(z_{1}-z_{2})}{(z_{1}-q^{2}z_{2})(z_{1}-q^{-2}z_{2})}\frac{\mathrm{d}z_{1}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}$
$=\frac{1}{2\pi i}\int_{\mathbb{T}}\mathrm{Res}_{z_{1}=0}\frac{\Gamma^{12}}{z_{1}}+\mathrm{Res}_{z_{1}=q^{-2}z_{2}}\frac{\Gamma^{12}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}$
$=1+\frac{1}{2\pi i}\int_{\mathbb{T}}\frac{(q^{-2}-1)^{2}q^{2}}{(q^{-2}-q^{2})}\frac{\mathrm{d}z_{2}}{z_{2}}$
$=1+\frac{(1+q)(1-q^{2})}{1+q+q^{2}+q^{3}}$
$=\frac{2}{1+q^{2}}.$
Accounting for $c_{M}$, we see that $J_{\mathbf{c}}$ is regular at $\mathbf{q}=\zeta_{1},\zeta_{2}$.
3.1.2 Proofs of Theorem 11, Lemma 6, and Corollary 2
Proof of Theorem 11.
Let $n=l_{0}+\cdots+l_{N}$. We may assume that $l_{0}\leq l_{2}\leq\cdots\leq l_{N}$. It suffices to evaluate the integral
$\left(\frac{1}{2\pi i}\right)^{N+1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod\Gamma_{i,j,g}\frac{\mathrm{d}z_{0}}{z_{0}}\cdots\frac{\mathrm{d}z_{N}}{z_{N}}=\left(\frac{1}{2\pi i}\right)^{N+1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{i<j}\Gamma^{ij}\frac{\mathrm{d}z_{0}}{z_{0}}\cdots\frac{\mathrm{d}z_{N}}{z_{N}}.$
We claim that the value of this integral is
$c_{M}\sum_{\mathrm{trees}~{}T}\sum_{\mathrm{branches}~{}B\mathrm{of}~{}T}\prod_{\begin{subarray}{c}C\prec B\\ C=\{i_{0},\dots,i_{t}\}\end{subarray}}\prod_{k=0}^{t-1}\frac{(1-q^{i_{k}i_{k+1}}q_{i_{k}i_{k+1}})(1-q^{i_{k}i_{k+1}}(q_{i_{k}i_{k+1}})^{-1})}{1-(q^{i_{k}i_{k+1}})^{2}}\\ \cdot\prod_{r=0}^{k-1}\frac{(1-Q_{rk}q_{i_{r}i_{k+1}})(1-Q_{rk}(q_{i_{r}i_{k+1}})^{-1})}{(1-Q_{rk}q^{i_{r}i_{k+1}})(1-Q_{rk}(q^{i_{r}i_{k+1}})^{-1})},$ (15)
where the sum over trees is taken over all bookkeeping trees for the integral. When $k=0$, we interpret the product over $r$ as being empty.
First we explain how (15) simplifies to (10). All cancellations will take place within the same clump $C$ of some branch $B$, which we now fix. We have
$1-Q_{rk}(q^{i_{r}i_{k+1}})^{-1}=1-q^{g_{i_{k}}+g_{i_{k+1}}+1}q^{i_{k}i_{k+1}}\cdots q^{i_{r}i_{r+1}}q^{-g_{i_{r}}-g_{i_{k+1}}-1}=1-Q_{r,k-1}q^{g_{i_{k}}-g_{i_{r}}},$
which is one of the factors in the product $(1-Q_{r,k-1}q_{i_{r}i_{k}})(Q_{r,k-1}(q_{i_{r}i_{k}})^{-1})$. The surviving factor in the numerator at index $(k-1,r)$ is then equal to $(1-Q_{r,k-1}q^{g_{i_{r}}-g_{i_{k}}})$. In short, the above factors in the denominator cancel with a numerator occurring with the same $r$-index but $k$-index one lower. Such a factor occurs whenever $r<k-1$ (note that this inequality does not hold when $k=1$ and $r=0$). When $r=k-1$, we have
$1-Q_{k-1,k}(q^{i_{k}i_{i+1}})^{-1}=1-q^{i_{k}i_{k+1}}q^{i_{k-1}i_{k}}(q^{i_{k-1}i_{k+1}})^{-1}=q^{g_{i_{k}}+g_{i_{k+1}}+1}q^{g_{i_{k-1}}-g_{i_{k+1}}}=1-q^{2g_{i_{k}}+1},$
which is one of the factors in $(1-q^{i_{k+1}i_{k}}q_{i_{k}i_{k+1}})(1-q^{i_{k}i_{k+1}}(q_{i_{k}}i_{k+1})^{-1})$. The cancellation leaves behind the factor $1-q^{l_{i_{k+1}}}$ in the numerator, except for $k=0$; this term keeps both its denominators. At this point we have shown that the factor corresponding to $C$ in (15) simplifies to
$\frac{(1-q^{l_{i_{0}}})(1-q^{l_{i_{1}}})}{1-q^{l_{i_{0}}+l_{i_{1}}}}\prod_{k=1}^{t-1}\frac{(1-q^{l_{i_{k+1}}})}{(1-q^{l_{i_{k}}+l_{i_{k+1}}})}\prod_{r=0}^{k-1}\frac{R_{rk}}{1-Q_{rk}q^{i_{r}i_{k+1}}},$
where
$R_{rk}=\begin{cases}1-Q_{rk}q^{g_{i_{r}}-g_{i_{k}}}&\text{if}~{}k<t-1\\ (1-Q_{r,t-1}q^{g_{i_{r}}-g_{i_{t}-1}})(1-Q_{r,t-1}q^{g_{i_{k}}-g_{i_{r}}})&\text{if}~{}k=t-1\end{cases}.$
This means that (15) simplifies to (10).
To prove (15), we will use the residue theorem for each variable consecutively, keeping track of the constant expressions in $q$ that we extract after integrating with respect to each variable $z_{i}$. More precisely, we will track what happens in a single summand corresponding to some set of successive choices of poles to take residues at. Note that all the rational functions that will appear, namely the $\Gamma^{ij}$ or the rational functions that result from substitutions into the $\Gamma^{ij}$, become equal to $1$ once $z_{i}$ or $z_{j}$ is set to zero. Therefore poles at $z_{i}=0$ serve simply to remove all factors involving $z_{i}$ from inside the integrand (we will see what these rational functions are below). It follows that the summand whose branch we are computing is a product over clumps in the corresponding branch, so it suffices to compute the value of a given clump for some ordered subsets $\{i_{0},i_{1},\dots,i_{l}\}$ of the indices $\{0,\dots,N\}$. As we are inside a clump, we will consider only poles occurring at nonzero complex numbers. Thus we are left only to determine what happens within a single clump.
We first integrate with respect to the variable $z_{i_{0}}$. The residue theorem gives
$\left(\frac{1}{2\pi i}\right)^{l+1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{i<j}\Gamma^{ij}\frac{\mathrm{d}z_{i_{0}}}{z_{i_{0}}}\cdots\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}$
$=\left(\frac{1}{2\pi i}\right)^{l}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\sum_{l\neq i_{0}}\mathrm{Res}_{z_{i_{0}}=(q^{i_{0}l})^{-1}z_{l}}\frac{1}{z_{i_{0}}}\prod_{i<j}\Gamma^{ij}\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\cdots\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}+\left(\frac{1}{2\pi i}\right)^{l}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{\begin{subarray}{c}i<j\\ i,j\neq i_{0}\end{subarray}}\Gamma^{ij}\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\cdots\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}.$
As noted above, the second integral belongs to a different branch (in fact, in the case of $i_{0}$, to a different tree); our procedure will deal with it separately, and we will now consider what happens with the first integral.
For the first integral, consider one of the summands corresponding to $z_{i_{0}}=(q^{i_{0}i_{1}})^{-1}z_{i_{1}}$ for some $i_{1}$. We have
$\left(\frac{1}{2\pi i}\right)^{l}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\mathrm{Res}_{z_{i_{0}}=(q^{i_{0}i_{1}})^{-1}z_{i_{1}}}\frac{1}{z_{i_{0}}}\prod_{i<j}\Gamma^{ij}\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\cdots\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}$
$=\frac{(1-q^{i_{0}i_{1}}q_{i_{0}i_{1}})(1-q^{i_{0}i_{1}}(q_{i_{0}i_{1}})^{-1})}{1-(q^{i_{0}i_{1}})^{2}}\left(\frac{1}{2\pi i}\right)^{l}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{j\neq i_{1},i_{0}}\frac{(z_{i_{1}}-q^{i_{0}i_{1}}q_{i_{0}j}z_{j})(z_{i_{1}}-q^{i_{0}i_{1}}(q_{i_{0}j})^{-1}z_{j})}{(z_{i_{1}}-q^{i_{0}i_{1}}q^{i_{0}j}z_{j})(z_{i_{1}}-q^{i_{0}i_{1}}(q^{i_{0}j})^{-1}z_{j})}$
$\prod_{\begin{subarray}{c}i<j\\ i,j\neq i_{0}\end{subarray}}\Gamma^{ij}\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\frac{\mathrm{d}z_{i_{2}}}{z_{i_{2}}}\cdots\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}.$
Recall that, if $i_{1}>2$, even though formally we have defined the symbols $q_{ij}$ and $q^{ij}$ only for $i<j$, the symmetry of the factors allows us to write $q_{ij}$ even if $i>j$, thanks to the factor with $(q_{ij})^{-1}$ also present in the numerator.
Now we integrate with respect to $z_{i_{1}}$. Observe that the leftmost product over $j\neq i_{1},i_{0}$ does not contribute poles. Indeed, the first factor each denominator does not have its zero contained in $\mathbb{T}$, and the second factor in each denominator has its zero at $z_{i_{1}}=q^{i_{0}i_{1}}(q^{i_{0}i_{2}})^{-1}z_{i_{2}}$. The power of $q$ appearing is
$g_{i_{0}}+g_{i_{1}}+1-(g_{i_{0}}+g_{i_{2}}+1)=g_{i_{1}}-g_{i_{2}},$
and so $q^{i_{0}i_{1}}(q^{i_{0}i_{2}})^{-1}$ is equal to $q_{i_{1}i_{2}}$ or $(q_{i_{2}i_{1}})^{-1}$, whichever is defined. Thus this zero cancels with a zero in the numerator of $\Gamma^{i_{1}i_{2}}$ or $\Gamma^{i_{2}i_{1}}$, whichever is defined. Therefore we need only consider the simple poles at $z_{i_{1}}=(q^{i_{1}i_{2}})^{-1}z_{i_{2}}$ for $i_{1}<i_{2}$ and $z_{i_{1}}=(q^{i_{2}i_{1}})^{-1}z_{i_{2}}$ for $i_{2}<i_{1}$. Observe that the residues will be the same for either inequality. In the case $i_{2}<i_{1}$, for example, $\Gamma^{i_{2}i_{1}}$ needs to be rewritten so that its simple pole inside $\mathbb{T}$ is in the correct format to calculate the residue by substitution:
$\mathrm{Res}_{z_{i_{1}}=(q^{i_{2}i_{1}})^{-1}z_{i_{2}}}\frac{(z_{i_{2}}-q_{i_{2}i_{1}}z_{i_{1}})(z_{i_{2}}-(q_{i_{2}i_{1}})^{-1}z_{i_{1}})}{(z_{i_{2}}-q^{i_{2}i_{1}}z_{i_{1}})(z_{i_{2}}-(q^{i_{2}i_{1}})^{-1}z_{i_{1}})}\frac{1}{z_{i_{1}}}$
$=\mathrm{Res}_{z_{i_{1}}=(q^{i_{2}i_{1}})^{-1}z_{i_{2}}}\frac{-(q^{i_{2}i_{1}})^{-1}(z_{i_{2}}-q_{i_{2}i_{1}}z_{i_{1}})(z_{i_{2}}-(q_{i_{2}i_{1}})^{-1}z_{i_{1}})}{(z_{i_{2}}-q^{i_{2}i_{1}}z_{i_{1}})(z_{i_{2}}-(q^{i_{2}i_{1}})^{-1}z_{i_{1}})}\frac{1}{z_{i_{1}}}$
$=\frac{(1-q^{i_{2}i_{1}}q_{i_{2}i_{1}})(1-q^{i_{2}i_{1}}(q_{i_{2}i_{1}})^{-1})}{1-(q^{i_{2}i_{1}})^{2}}.$
Therefore after integrating within the clump at hand with respect to $z_{i_{0}}$ and then $z_{i_{1}}$, we have the expression
$\frac{(1-q^{i_{0}i_{1}}q_{i_{0}i_{1}})(1-q^{i_{0}i_{1}}(q_{i_{0}i_{1}})^{-1})}{1-(q^{i_{0}i_{1}})^{2}}\frac{(1-q^{i_{2}i_{1}}q^{i_{0}i_{1}}q_{i_{0}i_{2}})(1-q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q_{i_{0}i_{2}})^{-1})}{(1-q^{i_{2}i_{1}}q^{i_{0}i_{1}}q^{i_{0}i_{2}})(1-q^{i_{2}i_{1}}q^{i_{0}i_{1}}(q^{i_{0}i_{2}})^{-1})}$ (16)
$\cdot$ $\frac{(1-q^{i_{2}i_{1}}q_{i_{2}i_{1}})(1-q^{i_{2}i_{1}}(q_{i_{2}i_{1}})^{-1})}{(1-(q^{i_{2}i_{1}})^{2})}$
$\cdot$ $\left(\frac{1}{2\pi i}\right)^{l-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{j\neq i_{0},i_{1},i_{2}}\frac{(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}q_{i_{0}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q_{i_{0}j})^{-1}z_{j})}{(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}q^{i_{0}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q^{i_{0}j})^{-1}z_{j})}$
$\cdot$ $\prod_{j\neq i_{0},i_{1},i_{2}}\frac{(z_{i_{2}}-q^{i_{1}i_{2}}q_{i_{1}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}(q_{i_{1}j})^{-1}z_{j})}{(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{1}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}(q^{i_{1}j})^{-1}z_{j})}\prod_{\begin{subarray}{c}i<j\\ i,j\neq i_{0},i_{1}\end{subarray}}\Gamma^{ij}\frac{\mathrm{d}z_{i_{2}}}{z_{i_{2}}}\cdots\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}.$
Now we integrate with respect to $z_{i_{2}}$. Again, only poles from the product of $\Gamma^{ij}$’s occur with nonzero residues: in total we have simple poles contained in $\mathbb{T}$ possibly at $z_{i_{2}}=q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q^{i_{0}j})^{-1}z_{j}$, at $z_{i_{2}}=q^{i_{1}i_{2}}(q^{i_{1}j})^{-1}z_{j}$ and at $z_{i_{2}}=(q^{i_{2},j})^{-1}z_{j}$ for $j\neq i_{0},i_{1},i_{2}$. It may happen that these poles are not all distinct, but all zeros in the denominator of the former two types are in fact cancelled by zeros of denominator anyway. The necessary factors occur in the product immediately adjacent on the right. Indeed, we have
$g_{i_{1}}+g_{i_{2}}+1+g_{i_{0}}+g_{i_{1}}+1-g_{i_{0}}-g_{j}-1=g_{i_{1}}+g_{i_{2}}+1+g_{i_{1}}-g_{j}$
and so either $q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q_{i_{0}j})^{-1}=q^{i_{1}i_{2}}q_{i_{1}j}$ or $q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q_{i_{0}j})^{-1}=q^{i_{1}i_{2}}(q_{i_{1}j})^{-1}$ (or $q_{ji_{1}}$ or $(q_{ji_{1}})^{-1}$).
Likewise we have
$g_{i_{1}}+g_{i_{2}}+1-g_{i_{1}}-g_{j}-1=g_{i_{2}}-g_{j}$
as happened when we integrated with respect to $z_{i_{0}}$.
Now we see the following pattern: At each stage, we integrate with respect to the variable we took a residue at in the previous step. When integrating with respect to $z_{i_{r}}$, there will be $r+1$ products of rational functions, and each rational function in the leftmost $r$ products will contribute a pole, in addition to the pole contributed by $\Gamma^{i_{r}i_{r+1}}$ or $\Gamma^{i_{r+1}i_{r}}$, whichever is defined. However, each is nullified by having zero residue thanks to a denominator in the product immediately to the right. Integrating with respect to $z_{i_{r}}$ will result in extracting $r+1$ new rational factors, each of the form claimed in the theorem. When it comes to integrating with respect to $z_{i_{l}}$, all variables $z_{i_{l}}$ will cancel from the remains of the Gamma functions by homogeneity, and the factor $\frac{1}{z_{i_{l}}}$ will result in the final integral contributing just the remaining $l+1$ rational factors. ∎
Lemma 6 is a porism of the preceding proof.
Proof of Lemma 6.
We will evaluate the integral (14) by applying the residue theorem successively for each variable as in the proof of Theorem 11. Note that as functions of any variable $z_{k}$, the functions $\prod_{i<j}\Gamma^{ij}$ and all the other products, for example those appearing in (16), have numerator and denominator with equal degrees. Thus the overall sum of powers of all $z_{i}$ in the integrand of (14) is $e_{0}+\cdots+e_{N}$, and in general, the sum of powers of all $z_{i}$ in the integrand of an expression like (16) is the sum of the degrees of the monomial $z_{i_{j}}$ terms.
When evaluating (14) along a single branch, we find again that the only poles that appear are of the form $z_{i}=(q^{ij})^{-1}z_{j}$, or $z_{i}^{e_{i}}=0$ for $e_{i}<0$. We will track the effect that evaluating each successive residue has on the total degree of the integrand, and then conclude using the fact that $\int_{\mathbb{T}}z^{r}\,\mathrm{d}z=2\pi i\delta_{r,-1}$. First, observe that evaluating a residue of the form $z_{i}=(q^{ij})^{-1}z_{j}$ increases the sum of all powers by $1$; a factor $z_{j}$ is contributed to the resulting integrand. The sum of all powers is likewise increased by $1$ when evaluating the residue at a simple pole of $z_{i}^{-1}$ at 00. To compute the residue at a pole of $z_{i}^{-e_{i}}$ at 00 for $e_{i}>1$, consider the Taylor expansion at 00 of $\Gamma^{ij}$. We have
$\frac{1}{z_{i}-q^{ij}z_{j}}=-\frac{1}{q^{ij}z_{j}}-\frac{z_{i}}{(q^{ij}z_{j})^{2}}-\frac{z_{i}^{2}}{(q^{ij}z_{j})^{3}}-\cdots$
and
$\frac{1}{z_{i}-(q^{ij})^{-1}z_{j}}=-\frac{q^{ij}}{z_{j}}-\left(\frac{q^{ij}}{z_{j}}\right)^{2}z_{i}-\left(\frac{q^{ij}}{z_{j}}\right)^{3}z_{i}^{2}-\cdots.$
Multiplying these series and further multiplying by the denominator $z_{i}^{2}-(q_{ij}+q_{ij}^{-1})z_{i}z_{j}+z_{j}$, it follows that the Taylor expansion of $\Gamma^{ij}$ is
$1+\frac{q^{ij}+(q^{ij})^{-1}-q_{ij}-q_{ij}^{-1}}{z_{j}}z_{i}+\frac{(q^{ij})^{2}+2+(q^{ij})^{-2}-(q_{ij}+q_{ij}^{-1})(q^{ij}+q^{-ij})}{z_{j}}z_{i}^{2}+\cdots\\ +\frac{-(q_{ij}+q_{ij}^{-1})((q^{ij})^{n-1}+\cdots+(q^{ij})^{-n+1})+(q^{ij})^{n-2}+\cdots+(q^{ij})^{-n+2}+(q^{ij})^{n}+\cdots+(q^{ij})^{-n}}{z_{j}^{n}}z_{i}^{n}\\ +\cdots.$ (17)
It is clear that the salient point of (17), that $z_{i}^{n}$ appears with a coefficient proportional to $z_{j}^{-n}$, holds also for the all the products like those in (16). Thus computing a residue at the pole of $z_{i}^{-e_{i}}$ at 00 will result in a new integrand, the total degree of which has increased by $e_{i}-e_{i}+1=1$. It now follows that after integrating with respect to $N-1$ variables, the final integral to be computed will be a constant times $(2\pi i)^{-1}\int_{\mathbb{T}}z_{N}^{e_{0}+\cdots+e_{N}+N-1}\mathrm{d}z_{N}$. This is nonzero if and only if $e_{0}+\cdots+e_{N}=-N$. Summing over all branches of a bookkeeping tree, we see that (14) is nonzero only if $e_{0}+\cdots+e_{N}=-N$. ∎
The lemma is helpful for computing explicit examples, as it points out that one can always avoid dealing with higher-order poles. Indeed, given an integral of the form (14), we may assume by Lemma 6 that $e_{0}+\cdots+e_{N}=-N$. If $e_{i}=-1$ for all $i$ then the integral (14) is just the integral from Theorem 11. Otherwise there is some $e_{i_{0}}\geq 0$. We may assume that $i_{0}=0$. Then the only poles in $z_{0}$ are of the form $z_{0}=(q^{0i_{1}})^{-1}z_{i_{1}}$ for indices $i_{1}>0$. Therefore we compute that (14) is equal to a sum of terms of the form
$\frac{(1-q^{0i_{1}}q_{0i_{1}})(1-q^{0i_{1}}(q_{0i_{1}})^{-1})}{1-(q^{0i_{1}})^{2}}(q^{0i_{1}})^{-e_{0}-1}\left(\frac{1}{2\pi i}\right)^{N-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}z_{i_{1}}^{e_{i_{1}}+1+e_{0}}z_{i_{2}}^{e_{i_{2}}}\cdots z_{i_{N}}^{e_{i_{N}}}\\ \cdot\prod_{j\neq i_{1},0}\frac{(z_{i_{1}}-q^{0i_{1}}q_{0j}z_{j})(z_{i_{1}}-q^{0i_{1}}(q_{0j})^{-1}z_{j})}{(z_{i_{1}}-q^{0i_{1}}q^{0j}z_{j})(z_{i_{1}}-q^{0i_{1}}(q^{0j})^{-1}z_{j})}\prod_{\begin{subarray}{c}i<j\\ i,j\neq 0\end{subarray}}\Gamma^{ij}\mathrm{d}z_{i_{1}}\cdots\mathrm{d}z_{i_{N}}.$ (18)
The total degree of the integrand is now $e_{0}+\cdots+e_{N}+1=-(N-1)$. Therefore either $e_{i_{0}}+e_{i_{1}}+1=e_{i_{2}}=\cdots=e_{-N}=-1$, or we may again assume without loss of generality that the exponent of some $z_{i_{j}}$ is nonnegative, and proceed with evaluating (18) by integrating with respect to $z_{i_{j}}$. We may continue in this way, never having to deal with more than a simple pole at 00. (Of course, the resulting order of integration need not be the same as the order used in the proofs of Theorem 11 and Corollary 2.)
We can now prove Corollary 2.
Proof of Corollary 2.
In light of Lemma 4, it is enough to prove that the conclusions of the present corollary hold for integrals of the form (14). We again follow the algorithm from the proof of Theorem 11, so it is sufficient to consider a single branch, and for this it suffices to observe that, given, for example, a variant
$\left(\frac{1}{2\pi i}\right)^{l-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{j\neq i_{0},i_{1},i_{2}}\frac{(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}q_{i_{0}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q_{i_{0}j})^{-1}z_{j})}{(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}q^{i_{0}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{0}i_{1}}(q^{i_{0}j})^{-1}z_{j})}$
$\cdot\prod_{j\neq i_{0},i_{1},i_{2}}\frac{(z_{i_{2}}-q^{i_{1}i_{2}}q_{i_{1}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}(q_{i_{1}j})^{-1}z_{j})}{(z_{i_{2}}-q^{i_{1}i_{2}}q^{i_{1}j}z_{j})(z_{i_{2}}-q^{i_{1}i_{2}}(q^{i_{1}j})^{-1}z_{j})}\prod_{\begin{subarray}{c}i<j\\ i,j\neq i_{0},i_{1}\end{subarray}}\Gamma^{ij}z_{i_{2}}^{e_{i_{2}}}\cdot\cdots\cdot z_{i_{l}}^{e_{i_{l}}}\mathrm{d}z_{i_{2}}\cdots\mathrm{d}z_{i_{l}}$
of (16) in which $e_{i_{2}}<-1$, the quotient rule gives that the residue at $z_{i_{2}}=0$ is a equal to a linear combination over $\mathbb{Q}[q^{\frac{1}{2}},q^{-\frac{1}{2}}]$ of integrals of the form
$\left(\frac{1}{2\pi i}\right)^{l-2}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{\begin{subarray}{c}i<j\\ i,j\neq i_{0},i_{1},i_{2}\end{subarray}}\Gamma^{ij}z_{i_{3}}^{e_{i_{3}}^{\prime}}\cdot\cdots\cdot z_{i_{l}}^{e_{i_{l}}^{\prime}}\mathrm{d}z_{i_{3}}\cdots\mathrm{d}z_{i_{l}}$
for new exponents $e^{\prime}_{i_{k}}$. This and the calculations in the proof of Theorem 11 make clear that the only positive powers of $q^{1/2}$ that appear in the any denominator are those that appeared as in Theorem 11. These are controlled by the possible block sizes of $M$, and hence are bounded in terms of $\tilde{W}$. Throughout this procedure, the denominator has been contributed to only by the Plancherel density itself, which depends only on $M$. The last claim of the corollary now follows from Proposition 1. ∎
3.2 The functions $f_{w}$ for general $\mathbf{G}$
For general $\mathbf{G}$, we will follow the same plan as for $\mathbf{G}=\mathrm{GL}_{n}$. The only difference is that we have less control over which denominators can appear. Indeed, this is true even for formal degrees, but complications are also introduced by residual coset we integrate over, or equivalently, by lack of explicit control of the Satake parameter of $\pi=i_{P}^{G}(\omega\otimes\nu)$ for arbitrary discrete series representations $\omega\in\mathcal{E}_{2}^{I}(M)$.
Theorem 12.
Let $w\in\tilde{W}$. Then $f_{w}(1)$ is a rational function of $q$ with poles drawn from a finite set of roots of unity depending only on $\tilde{W}$. The numerator is a Laurent polynomial in $q^{1/2}$. The denominator depends only on the two-sided cell containing $w$. If $w$ is in the lowest two-sided cell, then the denominator divides the Poincaré polynomial of $\mathbf{G}$.
In the proof, we do not attempt to record in any information about degrees of the numerators. It will therefore be necessary to control the possible numerators of $f_{w}(1)$ in a different manner than for $\mathbf{G}=\mathrm{GL}_{n}$ in order to prove Proposition 2 below.
Proof of Theorem 12.
First, we note that the reasoning for the lowest cell used in the proof of Corollary 1 holds for arbitrary $G$. Therefore (12) proves the last claim just as for $\mathrm{GL}_{n}$.
More generally, let $w\in\tilde{W}$ and $M_{P}$ be a Levi subgroup corresponding to the two-sided cell containing $w$. Let $N=\operatorname{rk}A_{P}$. Given a coroot $\alpha^{\vee}$ and a basis of the cocharacter lattice as explained in Section 2.7, we write $\alpha^{\vee}=z_{1}^{e_{1}}\cdots z_{n}^{e_{n}}$ for integers $e_{i}=e_{i}(\alpha)$.
By Lemma 4 and Theorem 6, it suffices to show the conclusions of the theorem hold for integrals of the form
$\left(\frac{1}{2\pi i}\right)^{n}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\prod_{\alpha\in R_{1,+}\setminus R_{P,1,+}}\frac{q_{\alpha}(z_{1}^{e_{1}}z_{2}^{e_{2}}\cdots z_{n}^{e_{n}}-q_{\omega})(z_{1}^{e_{1}}z_{2}^{e_{2}}\cdots z_{n}^{e_{n}}-q_{\omega}^{-1})}{\left(z_{1}^{\frac{e_{1}}{2}}z_{2}^{\frac{e_{2}}{2}}\cdots z_{n}^{\frac{e_{n}}{2}}+q_{\omega}^{1/2}q_{\alpha}^{1/2}\right)\left(z_{1}^{\frac{e_{1}}{2}}z_{2}^{\frac{e_{2}}{2}}\cdots z_{n}^{\frac{e_{n}}{2}}+q_{\omega}^{-1/2}q_{\alpha}^{-1/2}\right)}\\ \cdot\frac{1}{\left(z_{1}^{\frac{e_{1}}{2}}z_{2}^{\frac{e_{2}}{2}}\cdots z_{n}^{\frac{e_{n}}{2}}-q_{\omega}^{1/2}q_{\alpha}^{1/2}q_{2\alpha}\right)\left(z_{1}^{\frac{e_{1}}{2}}z_{2}^{\frac{e_{2}}{2}}\cdots z_{n}^{\frac{e_{n}}{2}}-q_{\omega}^{1/2}q_{\alpha}^{1/2}q_{2\alpha}\right)}z_{1}^{f_{1}}\cdots z_{n}^{f_{n}}\,\frac{\mathrm{d}z_{1}}{z_{1}}\cdots\frac{\mathrm{d}z_{n}}{z_{n}},$ (19)
where $f_{i}\in\mathbb{Z}$ and $q_{\omega}=q_{\omega,\alpha}$ is the value of $\alpha^{\vee}$ on the Satake parameter of $\omega$—this is a positive power of $q^{1/2}$ (of $q$ if $\alpha^{\vee}/2$ is a coroot) by Sections 7 and 8 of [OpdamSpectral]. We have $q_{2\alpha}=1$ and the resulting simplification (7) whenever $\alpha^{\vee}/2$ is not a coroot.
With notation fixed, the theorem is essentially an observation. Indeed, suppose that we integrate with respect to $z_{1}$, and wish to compute the contribution to the residue at a pole
$z_{1}=\xi q_{\omega}^{-\frac{1}{e_{1}}}q_{\alpha}^{-\frac{1}{e_{1}}}z_{2}^{\frac{e_{2}}{e_{1}}}\cdots z_{n}^{\frac{e_{n}}{e_{1}}}$
arising from the factor
$\frac{q_{\alpha}(z_{1}^{e_{1}}z_{2}^{e_{2}}\cdots z_{n}^{e_{n}}-q_{\omega})(z_{1}^{e_{1}}z_{2}^{e_{2}}\cdots z_{n}^{e_{n}}-q_{\omega}^{-1})}{(z_{1}^{e_{1}}z_{2}^{e_{2}}\cdots z_{n}^{e_{n}}-q_{\omega}q_{\alpha})(z_{1}^{e_{1}}z_{2}^{e_{2}}\cdots z_{n}^{e_{n}}-q_{\omega}^{-1}q_{\alpha}^{-1})}\frac{1}{z_{1}}=\\ \frac{q_{\alpha}(z_{1}^{e_{1}}-q_{\omega}z_{2}^{-e_{2}}\cdots,z_{n}^{-e_{n}})(z_{1}^{e_{1}}-q_{\omega}^{-1}z_{2}^{-e_{2}}\cdots,z_{n}^{-e_{n}})}{(z_{1}^{e_{1}}-q_{\omega}q_{\alpha}z_{2}^{-e_{2}}\cdots,z_{n}^{-e_{n}})\prod_{\zeta}(z_{1}-\zeta q_{\omega}^{-\frac{1}{e_{1}}}q_{\alpha}^{-\frac{1}{e_{1}}}z_{2}^{\frac{e_{2}}{e_{1}}}\cdots z_{n}^{\frac{e_{n}}{e_{1}}})}\frac{1}{z_{1}},$
where $\xi$ and the $\zeta$ are primitive $e_{1}$-st roots of unity. The contribution is then
$\frac{q_{\alpha}(q_{\omega}^{-1}q_{\alpha}^{-1}z_{2}^{-e_{2}}\cdots z_{n}^{-e_{n}}-q_{\omega}z_{2}^{-e_{2}}\cdots z_{n}^{-e_{n}})(q_{\omega}^{-1}q_{\alpha}^{-1}z_{2}^{-e_{2}}\cdots z_{n}^{-e_{n}}-q_{\omega}^{-1}z_{2}^{-e_{2}}\cdots z_{n}^{-e_{n}})}{(q_{\omega}^{-1}q_{\alpha}^{-1}z_{2}^{-e_{2}}\cdots z_{n}^{-e_{n}}-q_{\omega}q_{\alpha}z_{2}^{-e_{2}}\cdots z_{n}^{-e_{n}})\prod_{\zeta\neq\xi}(\xi q_{\omega}^{\frac{-1}{e_{1}}}q_{\alpha}^{\frac{-1}{e_{1}}}z_{2}^{\frac{-e_{2}}{e_{1}}}\cdots z_{n}^{\frac{-e_{n}}{e_{1}}}-\zeta q_{\omega}^{\frac{-1}{e_{1}}}z_{2}^{\frac{-e_{2}}{e_{1}}}\cdots z_{n}^{\frac{-e_{n}}{e_{1}}})}\\ \cdot\frac{1}{\xi q_{\omega}^{-\frac{1}{e_{1}}}q_{\alpha}^{-\frac{1}{e_{1}}}z_{2}^{\frac{e_{2}}{e_{1}}}\cdots z_{n}^{\frac{e_{n}}{e_{1}}}},$ (20)
which simplifies to
$\frac{(1-q_{\alpha}q_{\omega}^{2})(1-q_{\alpha})}{(1-q_{\alpha}^{2}q_{\omega}^{2})\prod_{\zeta\neq\xi}(1-\xi^{-1}\zeta)}.$ (21)
We have used no special properties of the integers $e_{i}$ or $f_{j}$, and thus after integrating with respect to $z_{1}$, (19) is equal to a sum of integrals with respect to $z_{2},\dots,z_{n}$ again of the form (19), except the factors in the denominator will now involve powers $q_{\omega}(\alpha)q_{\omega}(\alpha^{\prime})^{1/2}$ and $q_{\omega}(\alpha)q_{\omega}(\alpha^{\prime})^{-1/2}$, as for $\mathbf{G}=\mathrm{GL}_{n}$. The coefficients of the this sum are the form
$\frac{Q}{(1\pm q^{r_{1}})^{c_{1}}\cdots(1\pm q^{r_{k}})^{c_{k}}},$
where $Q$ is a Laurent polynomial in $q^{\frac{1}{2}}$, $r_{i}\in\frac{1}{2}\mathbb{N}$ with, a priori, complex coefficients, and $c_{i}\in\mathbb{N}$. Indeed, as we never required any cancellations with the numerator to extract rational functions of $q$ of the required form, it is clear that the simplification of (20) to (21) works essentially the same way for any higher order poles that appear—again thanks to the quotient rule—and that when integrating with respect to subsequent variables, the additional rational functions of $q$ appearing have the same shape as in the proof of Theorem 11. It is also clear that the factors corresponding to non-reduced roots behave similarly.
Therefore (19) has poles in $q$ only at a finite number of roots of unity. Again by the quotient rule, exponents $r_{i}$ and $c_{i}$ depend only on $M$. Clearly there are only finitely-many exponents $r_{i}$ that appear for any $M$. The theorem now follows. ∎
3.3 Relating $t_{w}$ and $f_{w}$
[…]
3.3.1 Completions of $\mathbf{H}$ and $J\otimes_{\mathbb{C}}\mathcal{A}$
[…]
where we say that $b_{x}\to 0$ as $\ell(x)\to\infty$ if for all $N>0$, $b_{x}\in(\mathbf{q}^{-1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $x$ sufficiently long.
Consider also the completions
[…]
and
[…]
of $\mathbf{H}$ (note the difference between $C^{\prime}_{x}$ and $C_{x}$), as well as the completion
[…]
of $J\otimes_{\mathbb{C}}\mathcal{A}$.
[…]
Lemma 7.
[…]
Proof.
Given an infinite sum $\sum_{x}b_{x}C^{\prime}_{x}$, upon rewriting this sum in the standard basis, the coefficient of some $T_{y}$ is
$a_{y}:=\sum_{x\geq y}b_{x}\mathbf{q}^{-\frac{\ell(x)}{2}}P_{y,x}(\mathbf{q}).$
As $\deg P_{y,x}\leq\frac{1}{2}(\ell(x)-\ell(y)-1)$, we have that $\mathbf{q}^{-\frac{\ell(x)}{2}}P_{y,x}(\mathbf{q})$ is a polynomial in $\mathbf{q}^{-1/2}$. Therefore the above sum defines a formal Laurent series. Moreover, as $\ell(y)\to\infty$, it is clear that $a_{y}\to 0$. Continuity of the inclusion is clear from the formula for $a_{y}$. ∎
3.3.2 The functions $f_{w}$ and the basis elements $t_{w}$
[…]
Proposition 2.
[…]
The most difficult part of the proof of the proposition is showing that $a_{x,w}\to 0$ as $\ell(x)\to\infty$. To this end we have
Lemma 8.
Let $w\in\tilde{W}$. Then the degree in $q$ of the numerator of $f_{w}(1)$ is bounded uniformly in $w$ by some $N$ depending only on $\tilde{W}$, and hence in the notation of Proposition 2 above, we have $a_{1,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $w\in\tilde{W}$.
Remark 11.
In type $A$, the Lemma follows immediately from the order of integration given after the proof of Lemma 6.
Proof of Lemma 8.
Let $\pi_{\omega}=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$ be a tempered $I$-spherical representation of $G$ arising by induction from a parabolic $P$ such that $\operatorname{rank}A_{P}=k$. Let $\mathbb{T}_{\omega}^{k}$ be the compact torus parametrizing twists of $\omega$, and let $t_{w}$ be given. By Lemma 4, $\mathrm{trace}\left({\pi}\,,{f_{w}}\right)$ is a regular function on $\mathbb{T}^{k}$, i.e. a Laurent polynomial in the coordinates $z_{1},\dots,z_{k}$ on $\mathbb{T}^{k}$. The coefficients of this Laurent polynomial are independent of $q$, as $J$ and its representation theory are independent of $q$. As we have
$|1-q^{r}|\leq|z_{i}-q^{r}z_{1}^{e_{1}}\cdots z_{k}^{e_{k}}|\leq|1+q^{r}|$
for all $z_{i}\in\mathbb{T}$ and $r$, we may bound, for large $q$, the absolute value $|\mu_{M_{P}}(z)|$ of the Plancherel density for $M_{P}$ by a rational function $U_{M_{P},\omega}(q)$. By Theorem 6, we may do the same for formal degrees $d(\omega)$, bounding $|d(\omega)|$ by a rational function $D_{\omega}(q)$. Thus we define a rational function of $q$ by
$\frac{h_{1}(q)+h_{2}(q^{-1})}{k_{1}(q)+k_{2}(q^{-1})}=\sum_{\omega}U_{M_{P},\omega}(q)D_{\omega}(q)\max_{z\in\mathbb{T}_{\omega}^{k}}|\mathrm{trace}\left({\pi_{\omega}}\,,{f_{w}}\right)|,$
where $h_{1},k_{1}\in\mathbb{C}[q]$ and $h_{2},k_{2}\in q^{-1}\mathbb{C}[q^{-1}]$, and the sum is over the finitely many, up to unitary twist, pairs $(M,\omega)$, for $\omega\in\mathcal{E}_{2}(M)$, such that $\mathrm{trace}\left({\pi_{\omega}}\,,{f_{w}}\right)\neq 0$.
Fix $w\in\tilde{W}$. Then
$|f_{w}(1)|=\left|\sum_{(P,\omega)}\int_{\mathcal{O}(\omega)}\mathrm{trace}\left({\pi_{\omega}}\,,{f_{w}}\right)d(\omega)\,\mathrm{d}\mu_{M_{P}}(z)\right|\leq\frac{h_{1}(q)+h_{2}(q^{-1})}{k_{1}(q)+k_{2}(q^{-1})}.$
Crucially, this expression holds for all $q\gg 1$.
On the other hand, by Corollary 2 and Theorem 12, $f_{w}(1)$ is a rational function of $q$, and for $q$ sufficiently large, we may write
$|f_{w}(1)|=\frac{f_{1}(q)+f_{2}(q^{-1})}{d_{1}(q)+d_{2}(q^{-1})}$
where $f_{1},d_{1}\in\mathbb{C}[q]$ and $f_{2},d_{2}\in q^{-1}\mathbb{C}[q^{-1}]$. Therefore for all $q\gg 1$ we have
$|(f_{1}(q)+f_{2}(q^{-1}))||(k_{1}(q)+k_{2}(q^{-1}))|\leq|(d_{1}(q)+d_{2}(q^{-1})||(h_{1}(q)+h_{1}(q^{-1}))|.$ (22)
We claim that this implies
$|f_{1}(q)k_{1}(q)|\leq|d_{1}(q)h_{1}(q)|$ (23)
for $q$ sufficiently large. Indeed, let $\epsilon>0$ be given and choose $q\gg 1$ such that
$\left||(f_{1}(q)+f_{2}(q^{-1}))||(k_{1}(q)+k_{2}(q^{-1}))|-|f_{1}(q)k_{1}(q)|\right|\leq\epsilon$
and
$\left||(d_{1}(q)+d_{2}(q^{-1}))||(h_{1}(q)+h_{2}(q^{-1}))|-|d_{1}(q)h_{1}(q)|\right|\leq\epsilon,$
and (22) holds. Then we have
$|f_{1}(q)||k_{1}(q)|$ $\leq|(f_{1}(q)+f_{2}(q^{-1}))||(k_{1}(q)+k_{2}(q^{-1}))|+\epsilon$
$\leq|(d_{1}(q)+d_{2}(q^{-1}))||(h_{1}(q)+h_{2}(q^{-1}))|+\epsilon$
$\leq|d_{1}(q)||h_{1}(q)|+2\epsilon,$
which proves (23).
Therefore
$\deg f_{1}\leq\deg f_{1}+\deg k_{1}\leq\deg d_{1}+\deg h_{1}.$
Now, the denominator of $f_{w}(1)$, and hence $\deg d_{1}$, depends only on the two-sided cell containing $w$, again by Corollary 2 and Theorem 12. We can also bound $\deg h_{1}$ uniformly in terms of $\tilde{W}$, as it depends only on the Plancherel measure and the finitely many possible formal degrees appearing in the parametrization of the $I$-spherical part of the tempered dual of $G$. This proves the lemma. ∎
If $t_{w}\in t_{d}Jt_{d^{\prime}}$ for $d\neq d^{\prime}$, then $t_{w}$ is a commutator. We record this observation as
Lemma 9.
We have that $f_{w}(1)=0$ unless $f_{d}\star f_{w}\star f_{d}\neq 0$ for some distinguished involution $d$.
Now we can prove the proposition.
Proof of Proposition 2.
[…]
By Corollary 2 and Theorem 12 , there is a formal power series in $\hat{\mathcal{A}}^{-}$ with constant term equal to $1$ that specializes to the denominator of $A_{1,w}$ when $\mathbf{q}=q$. Moreover, there is a unique formal Laurent series $a_{1,w}\in\hat{\mathcal{A}}$ such that $a_{1,w}(q)=A_{1,w}$ for all prime powers. Indeed, $a_{1,w}$ is convergent for $\mathbf{q}=q$, and the difference of any two such series defines a meromorphic function of $\mathbf{q}^{-1/2}$ outside the unit disk with zeros at $q=p^{r}$ for every $r\in\mathbb{N}$. As these prime powers accumulate at $\infty$, such a meromorphic function must be identically zero.
If $f\in\mathcal{C}^{I\times I}$ is a Harish-Chandra Schwartz function, then
$q^{-\ell(x)}(f\star T_{x^{-1}})(1)=q^{-\ell(x)}\int_{G}{f(g)T_{x^{-1}}(g^{-1})}\,\mathrm{d}{\mu_{I}(g)}\\ =q^{-\ell(x)}\int_{IxI}{f(g)}\,\mathrm{d}{\mu_{I}(g)}=q^{-\ell(x)}\mu_{I}(IxI)f(x)=f(x).$
By definition, $f_{w}(x)=A_{x,w}$. On the other hand, according to Lemma 2, we have
$q^{-\ell(x)}(f_{w}\star T_{x^{-1}})(1)$ $=q^{-\ell(x)}\left(\tilde{\phi}(t_{w})\star\tilde{\phi}\left(\phi_{q}(T_{x^{-1}}^{\dagger})\right)\right)(1)$
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}(T_{x^{-1}}^{\dagger})\right)(1)$
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})-\ell(y)}Q_{y,x^{-1}}(q)(C^{\prime}_{y})^{\dagger}\right)\right)(1)$ (24)
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)(C_{y})\right)\right)(1)$ (25)
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\\ d\in\mathcal{D}\\ a(d)=a(r)\end{subarray}}h_{y,d,r}t_{r}\right)(1)$ (26)
$=q^{-\ell(x)}\tilde{\phi}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ d\in\mathcal{D}\\ a(d)=a(r)=a(w)\end{subarray}}h_{y,d,r}t_{w}t_{r}\right)(1)$
$=q^{-\ell(x)}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ a(r)=a(w)\end{subarray}}h_{y,d_{w},r}(f_{w}\star f_{r})(1),$ (27)
where $d_{w}$ is the unique distinguished involution in the right cell containing $w$.
In line (24), we rewrote $T_{x^{-1}}$ in terms of the $C^{\prime}$-basis of $H$, using the inverse Kazhdan-Lusztig polynomials $Q_{y,x^{-1}}$. In line (25), we applied the Goldman involution (see Lemma 1). In line (26) we applied Lusztig’s map $\phi_{q}$, and then in line (27), we applied the map $\tilde{\phi}$. Also in line (27), we used that left (respectively right) cells give left (respectively right) ideals of $J$, and so $t_{w}t_{r}$ is an integral linear combination of $t_{z^{-1}}$, with $d\sim_{L}z^{-1}\sim_{R}d_{w}$. By lemma 9, $f_{z^{-1}}(1)\neq 0$ only if $d=d_{w}$, that is, $z^{-1}\sim_{L}d_{w}$.
We use (27) to define $a_{x,w}\in\hat{\mathcal{A}}$. By the same arguments as above, $a_{x,w}$ is unique and defines a meromorphic function of $\mathbf{q}^{-1/2}$. It remains to show that as $\ell(x)\to\infty$, $a_{x,w}\to 0$ in the $(\mathbf{q}^{-1/2})$-adic topology. This follows in fact from (27). Indeed, the product $f_{w}\star f_{r}$ is an $\mathbb{N}$-linear combination of functions $f_{z}$, and the values $f_{z}(1)$ are rational functions of $q$, the numerators of which have uniformly bounded degree in $q$ by Lemma 8. The polynomials $h_{y,d_{w},r}$ have bounded degree in $q$ (for example in terms of the $a$-function). Finally, the degree in $q$ of
$q^{-\ell(x)}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)$
is at most
$q^{-\ell(x)}q^{\frac{\ell(y)}{2}}q^{\frac{\ell(x^{-1})-\ell(y)-1}{2}}=q^{-\ell(x)}q^{\frac{\ell(y)}{2}}q^{\frac{\ell(x)-\ell(y)-1}{2}}=q^{\frac{-\ell(x)-1}{2}}\to 0$ (28)
as $\ell(x)\to\infty$. This completes the definition of $\phi_{1}$ as a map of $\mathcal{A}$-modules.
It is easy to see that $\phi_{1}$ is a morphism of rings, essentially because $\tilde{\phi}$ is. Indeed, we have
$\phi_{1}(t_{w}t_{w^{\prime}})=\sum_{z}\gamma_{w,w^{\prime},z^{-1}}\sum_{x}a_{x,z}T_{x}$ (29)
while on the other hand
$\phi_{1}(t_{w})\phi_{1}(t_{w^{\prime}})=\sum_{x}a_{x,w}T_{x}\cdot\sum_{y}a_{y,w^{\prime}}T_{y}$ (30)
and when $\mathbf{q}=q$, we have that (30) becomes by definition
$\tilde{\phi}(t_{w})\star\tilde{\phi}(t_{w^{\prime}})=\tilde{\phi}\left(\sum_{z}\gamma_{w,w^{\prime},z^{-1}}t_{z}\right)=\sum_{z}\sum_{x}\gamma_{w,w^{\prime},z^{-1}}A_{x,z}T_{x}.$
[…]
Remark 12.
[…]
Proposition 3.
There is a commutative diagram
[…]
Proof.
[…]
Note that as $\tilde{\phi}\circ\phi_{q}=()^{\dagger}$ on $H$ for all $q$, we have that $\phi_{1}=()^{\dagger}\circ\phi^{-1}$ on $\phi(\mathbf{H})$. This says that the diagram commutes.
[…]
$\phi_{1}\left(\sum_{w}b_{w}t_{w}\right)=\sum_{y}b^{\prime}_{y}T_{y},$
where
$b_{y}^{\prime}=\sum_{w}b_{w}a_{y,w}.$
We must first show that this infinite sum of elements of $\hat{\mathcal{A}}$ is well-defined. By Lemma 8 and (27), we have that there is $M\in\mathbb{N}$ such that $a_{y,w}\in(\mathbf{q}^{1/2})^{M}\hat{\mathcal{A}}^{-}$ for all $y,w$. Therefore $b^{\prime}_{y}$ is well-defined, and as $a_{y,w}\to 0$ as $\ell(y)\to\infty$, we have $b^{\prime}_{y}\to 0$ as $\ell(y)\to\infty$. Therefore $\phi_{1}$ extends to $\mathbfcal{J}$.
To show continuity, it suffices to show that if $\{\sum_{w}b_{w,n}t_{w}\}_{n}$ is a sequence of elements of $\mathbfcal{J}$ tending to 00 as $n\to\infty$, then
$\sum_{w}b_{w,n}\phi_{1}(t_{w})=\sum_{y}b^{\prime}_{y,n}T_{y}\to 0$
[…]
Corollary 3.
We have
$(\phi\circ()^{\dagger})^{-1}(t_{1})=f_{1}(1)\sum_{w\in\tilde{W}}(-1)^{\ell(w)}q^{-\ell(w)}T_{w}.$
Proof.
If $w=1$, everything in (27) reduces to $r=d_{w}=y=1$, and we need only recall that $Q_{1,x}(\mathbf{q})=1$ for any $x\in\tilde{W}$. This follows from unicity of the inverse Kazhdan-Lusztig polynomials and the identity $\sum_{x\leq w}(-1)^{\ell(x)}P_{x,w}(\mathbf{q})=0$ for $w\neq 1$ [BjornerBrenti, Exercise 5.17]. ∎
Corollary 4.
The map $\tilde{\phi}$ defined in [BK] and recalled in diagram (2) is injective.
Proof.
Let $q>1$ and let $j\neq 0$ be an element of $J$. We must show that $\tilde{\phi}(j)\neq 0$. By injectivity of $\phi_{1}$, we have $\phi_{1}(j)\neq 0$. By definition of the map $\phi_{1}$, this means that there exists $q_{0}>1$, $u,s\in G^{\vee}$ such that $us=su$ with $s$ compact, and a representation $\rho$ of $\pi_{0}(Z_{G^{\vee}}(u,s))$ such that $jK(u,s,\rho,q_{0})\neq 0$. But then $jK(u,s,\rho,q)\neq 0$, $K(u,s,\rho,q)$ being a different specialization of the restriction of the same $J$-module $E(u,s,\rho)$ as for $K(u,s,\rho,q_{0})$, and $K(u,s,\rho,q)$ is also tempered. It follows that $\tilde{\phi}(j)\neq 0$. ∎
3.4 Proof of Theorem 13 for $\mathrm{GL}_{n}$
Theorem 13.
Let $\tilde{W}$ be of type $\tilde{A}_{n}$. Then statements 1 and 2 in the statement of Theorem 2 are true, together with a stronger version of statement 3: Let $u=(r_{1},\dots,r_{k})$ be a unipotent conjugacy class in $\mathrm{GL}_{n}(\mathbb{C})$. Let $d$ be a distinguished involution in the two-sided cell $\mathbf{c}(u)\subset\tilde{W}(\mathrm{GL}_{n})$ corresponding to $u$. Let $\pi$ be the unique family of parabolic inductions that $t_{d}$ does not annihilate. Then $\operatorname{rank}\pi(t_{d})=1$.
Proof.
By Corollary 2 and Propositions 2 and 3, $a_{1,x}$ is a rational function of $\mathbf{q}$ for all $w$. Then equation (27) implies that $a_{x,w}$, being a sum of rational functions with Laurent polynomial coefficients, is a rational function of $\mathbf{q}$ for all $x$. The same equation, together with the fact that $J_{\mathbf{c}}$ is a two-sided ideal for each cell $\mathbf{c}$ shows that the denominator of $a_{x,w}$ depends only on the two-sided cell containing $w$. This proves the first claim. The second claim now follows from the first claim and the first statement of Corollary 2 and the fact that $\tilde{W}$ has finitely-many two-sided cells.
Finally, let $u=(r_{1},\dots,r_{k})$. We have $\pi=\operatorname{Ind}_{P}^{G}(\mathrm{St}_{M_{P}}\otimes\nu)$, for the unique Levi $M_{P}$ such that $u$ is distinguished in $M_{P}^{\vee}$, and as $\dim\mathrm{St}_{M_{P}}^{I}=1$, clearly we have
$\dim\pi^{I}=\frac{n!}{r_{1}!\cdots r_{k}!}.$
This is also the number of distinguished involutions in $\mathbf{c}(u)$, by [ShiLNM]. The claim follows from Corollary 4. ∎
3.5 Proof of Theorem 2
Proof of Theorem 2.
Theorem 12 together with Propositions 2 and 3 show that $a_{1,w}$ is a rational function of $\mathbf{q}$ with denominator depending only on the two-sided cell containing $w$. Equation (27) again shows that $a_{x,w}$ is a rational function of $\mathbf{q}$ with denominator depending only on the two-sided cell containing $w$; up to twists the set $\mathcal{E}_{2}(M)^{I}$ is finite for every Levi subgroup $M$, so we may multiply through to include the denominators of all required formal degrees, which are in fact rational of the correct form by Theorem 6. Therefore there is a polynomial $P_{\mathbf{G}}(\mathbf{q})$ that clears denominators of all $a_{x,w}$. This proves the first statement of the Theorem.
Now, by Proposition 9 and Remark 2 of [BDD], $\phi$ induces an surjection
$\bar{\phi}\colon\mathbf{H}/[\mathbf{H},\mathbf{H}]\left[\frac{1}{P_{W}(\mathbf{q})}\right]\mathrel{\twoheadrightarrow}J/[J,J]\otimes\mathcal{A}.$
Therefore for every $j\in J$, there is $N=N(j)\in\mathbb{N}$ and $h\in\mathbf{H}$ such that
$j\equiv h\frac{1}{P_{W}(\mathbf{q})^{N}}$
in $J/[J,J]\otimes\mathcal{A}$. Considering traces and invoking Proposition 3, we see that the denominator of every $a_{x,w}$ divides a power of $P_{W}(\mathbf{q})$. Therefore there is $N=N_{\tilde{W}}$ depending only $\tilde{W}$ such that we can take $P^{1}_{\mathbf{G}}(\mathbf{q})=P_{W}(\mathbf{q})^{N}$. This proves the second statement of the Theorem. The third statement was proven in Corollary 1. ∎
4 Representations with fixed vectors under parahoric subgroups
In this section we will give an application of the $J$-action on the tempered $H$-modules. The first statement is an immediate corollary of the existence of this action, but the second statement relies on Corollary 4.
Theorem 14.
Let $\mathcal{P}$ be a parahoric subgroup of $G$. Let $\pi$ be a simple tempered representation of $G$ with $I$-fixed vectors with Kazhdan-Lusztig parameter $(u,s,\rho)$. Let $w_{\mathcal{P}}$ be the longest element in the parabolic subgroup of $\tilde{W}$ defined by $\mathcal{P}$ and $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
  1. 1.
    If
    $\ell(w_{\mathcal{P}})>a(u)=\dim_{\mathbb{C}}\mathcal{B}_{u}^{\vee},$
    then
    $\pi^{\mathcal{P}}=\{0\}.$
  2. 2.
    Conversely, let $u_{\mathcal{P}}$ be the unipotent conjugacy class corresponding to the two-sided cell containing $w_{\mathcal{P}}$. Then there exists $s\in Z_{G^{\vee}}(u_{\mathcal{P}})$, a Levi subgroup $M^{\vee}$ of $G^{\vee}$ minimal such that $(u_{\mathcal{P}},s)\in M^{\vee}$, and a discrete series representation $\omega\in\mathcal{E}_{2}(M)$ such that
    $\pi^{\mathcal{P}}=i_{P_{M}}^{G}(\omega\otimes\nu)^{\mathcal{P}}\neq\{0\}$
    for all $\nu$ non-strictly positive and the parameter of $\pi$ is $(u_{\mathcal{P}},s)$.
Proof.
Let $\mathcal{P}$ and $w_{\mathcal{P}}$ be as in the statement. Then $C^{\prime}_{w_{\mathcal{P}}}$ is proportional by a power of $q$ to the indicator function $1_{\mathcal{P}}$ in $H$. Moreover, $w_{\mathcal{P}}$ is a distinguished involution, with $a(w_{\mathcal{P}})=\ell(w_{\mathcal{P}})$. By Proposition 3 and the fact that $\phi$ is “upper-triangular” with respect to the $a$-function, we have $(\phi\circ()^{\dagger})(C^{\prime}_{w_{\mathcal{P}}})J_{\mathbf{c}}=0$ for $\mathbf{c}$ corresponding to $u$ if $a(w_{\mathcal{P}})>a(u)$.
For the second statement, by Corollary 4, there is a tempered representation $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$ with $\pi(t_{w_{\mathcal{P}}})\neq 0$; the unipotent part of its parameter is $u_{\mathcal{P}}$. In particular, there is a vector $v\in\pi$ such that $\pi(t_{w_{\mathcal{P}}})v=v$. We have
$\left(\phi\circ()^{\dagger}\right)(C^{\prime}_{w_{\mathcal{P}}})t_{w_{P}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{\overset{d\in\mathcal{D}}{z\sim_{L}d}}h_{w_{\mathcal{P}},d,z}t_{z}t_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{z}h_{w_{\mathcal{P}},w_{\mathcal{P}},z}t_{z}=\operatorname{vol}(\mathcal{P})t_{w_{P}},$
as $t_{z}t_{w_{\mathcal{P}}}\neq 0$ only if $z\sim_{L}w_{\mathcal{P}}$, and $C_{w_{\mathcal{P}}}C_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\operatorname{vol}(\mathcal{P})C_{w_{\mathcal{P}}}$. Thus $v\in\pi^{\mathcal{P}}$. As $\mathrm{trace}\left({\pi}\,,{t_{d}}\right)$ is constant with respect to $\nu$, the last part of the claim follows. ∎
A version of this statement for enhanced parameters in the case $\mathcal{P}=G(\mathcal{O})$ appears in [APM, Prop. 10.1].
Example 8.
The principal series representations have $u=\{1\}$, maximal $a$-value, and fixed vectors under every maximal compact subgroup of $G$. On the other extreme, the Steinberg representation has $u$ regular, and does not have fixed vectors under any proper parahoric.
In addition to the interpretation given in Section 1.1.3, in the examples below, the $J$-action also detects which direct summands of a reducible $\mathcal{P}$-spherical tempered representation are themselves $\mathcal{P}$-spherical.
Example 9.
Let $G=\mathrm{SL}_{2}(F)$. As remarked in [D], the distinguished involutions in the lowest two-sided cell are the simple reflections $s_{0}$ and $s_{1}$, and each $t_{s_{i}}$ is invariant under one of the two conjugacy classes of maximal parahoric subgroup of $G$. For unitary principal series representations $\pi$, one has $\mathrm{trace}\left({\pi}\,,{t_{s_{0}}}\right)=\mathrm{trace}\left({\pi}\,,{t_{s_{1}}}\right)=1$. At the quadratic character, the corresponding principal series representation is reducible, and each summand contains fixed vectors under precisely one of the maximal parahorics. Indeed, in [D] this computation is carried out at the level of the Schwartz space of the basic affine space.
Example 10.
Let $G=\mathrm{SO}_{5}(F)=\mathrm{PGSp}_{4}(F)$, with affine Dynkin diagram labelled as
[…]
[labels=0,1,2] B[1]2.
There are five conjugacy classes of parahoric subgroup, each obtained by projection from $\mathrm{GSp}_{4}(F)$: the maximal parahoric $\mathrm{PGSp}_{4}(\mathcal{O})$, the image of the paramodular group $\mathsf{K}$ corresponding to $\{0,2\}$, the image of the Siegel parahoric subgroup $\mathsf{Si}$ corresponding to $\{1\}$, the image of the Klingen parahoric $\mathsf{Kl}$ corresponding to $\{0\}$, and the Iwahori subgroup.
The columns of the below table give all the $I$-spherical tempered representations of $G$ by denoting the representation $\operatorname{Ind}_{P}^{G}(\omega)$ by $\omega$. We recall the few cases of reducibility of these inductions immediately below. The rows list unipotent conjugacy classes in $G^{\vee}=\mathrm{Sp}_{4}(\mathbb{C})$ such that all tempered standard modules $K(u,s,\rho)$ are in row $u$. That is, the rows record which summand of $J$ acts on each representation.
[…]
$(1,\dots,1)$ $\nu$
$(2,1,1)$ $\nu\otimes\pm\mathrm{St}_{\mathrm{SO}_{3}}$
$(2,2)$ $\xi\mathrm{St}_{\mathrm{GL}_{2}}$$\pm\tau_{2}$
$(4)$$\pm\mathrm{St}$
The discrete series are as in [ReederFormalDeg]. The only reducibility, by [Matic, Prop. 3.3], is the reducibility
$\operatorname{Ind}_{P}^{G}(\xi\mathrm{St}_{\mathrm{GL}_{2}})=\tau_{\mathrm{triv}}\oplus\tau_{\mathrm{sgn}}$
for $\xi^{2}=1$.
We can now compute the traces of some elements $t_{d}$ using the description of the simple $J$-modules given in [XiBook]. For the cells $(1,\dots,1)$, $(2,1,1)$, and $(4)$, we have $\mathrm{trace}\left({\pi}\,,{t_{d}}\right)=1$ for all $\pi$. We have
$\mathrm{trace}\left({\tau_{\mathrm{triv}}}\,,{t_{s_{0}}}\right)=\mathrm{trace}\left({\tau_{2}}\,,{t_{s_{0}}}\right)=1=\mathrm{trace}\left({\tau_{\mathrm{sgn}}}\,,{t_{s_{1}},}\right)=\mathrm{trace}\left({\tau_{\mathrm{triv}}}\,,{t_{s_{1}}}\right).$
In the below table, $\mathrm{trace}\left({\pi}\,,{t_{d}}\right)$ is recorded in bold face, whereas the dimension of $\pi^{\mathcal{P}}$, taken from [BrooksSchmidt, Table A.15], is recorded in normal face. Representations attached to the same cell but belonging to different packets are separated with a dotted line.
$t_{w_{\mathcal{P}}}$\  $\pi=\operatorname{Ind}_{P}^{G}(-)$ $a$ $\nu$ $\nu\otimes\pm\mathrm{St}_{\mathrm{SO}_{3}}$ $\tau_{\mathrm{triv}}$ $\tau_{\mathrm{sgn}}$$\pm\tau_{2}$$\pm\mathrm{St}$ $\mathcal{P}$
$t_{w_{0}}$ 4 1 $\mathrm{PGSp}_{4}(\mathcal{O})$
$t_{s_{0}s_{2}}$ 2 2 1 $\mathsf{K}$
$t_{s_{0}}$ 1 4 2 1 1 $\mathsf{Kl}$
$t_{s_{1}}$ 1 4 1 1 1 $\mathsf{Si}$
$t_{1}$ 0 8 4 3 1 4 1 $I$

### v6

@@
 Abstract
-Let $\tilde{W}$ be an extended affine Weyl group, $\mathbf{H}$ be the corresponding affine Hecke algebra over the ring $\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$, and $J$ be Lusztig’s asymptotic Hecke algebra, viewed as a based ring with basis $\{t_{w}\}$.
+Let $W_{\mathrm{aff}}$ be an extended affine Weyl group, $\mathbf{H}$ be the corresponding affine Hecke algebra over the ring $\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$, and $J$ be Lusztig’s asymptotic Hecke algebra, viewed as a based ring with basis $\{t_{w}\}$.
 Viewing $J$ as a subalgebra of the $(\mathbf{q}^{-\frac{1}{2}})$-adic completion of $\mathbf{H}$ via Lusztig’s map $\phi$, we use Harish-Chandra’s Plancherel formula for $p$-adic groups to show that the coefficient of $T_{x}$ in $t_{w}$ is a rational function of $\mathbf{q}$, with denominator depending only on the two-sided cell containing $w$, and dividing a power of the Poincaré polynomial of the finite Weyl group.
@@
 1 Introduction
-Let $\tilde{W}$ be an affine Weyl group or extended affine Weyl group, and let $\mathbf{H}$ be its associated Hecke algebra over $\mathcal{A}:=\mathbb{C}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]$, where $\mathbf{q}$ is a formal variable.
-The representation theory of $\mathbf{H}$ is very well understood, behaving well and uniformly when $\mathbf{q}$ is specialized to any $q\in\mathbb{C}^{\times}$ that is not a root of the Poincaré polynomial $P_{W}$ of the finite Weyl group $W\subset\tilde{W}$.
+Let $W_{\mathrm{aff}}$ be an affine Weyl group or extended affine Weyl group, and let $\mathbf{H}$ be its associated Hecke algebra over $\mathcal{A}:=\mathbb{C}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]$, where $\mathbf{q}$ is a formal variable.
+The representation theory of $\mathbf{H}$ is very well understood, behaving well and uniformly when $\mathbf{q}$ is specialized to any $q\in\mathbb{C}^{\times}$ that is not a root of the Poincaré polynomial $P_{W}$ of the finite Weyl group $W\subset W_{\mathrm{aff}}$.
 When $\mathbf{q}$ is specialized to a prime power $q$, the category of finite-dimensional modules over the specialized algebra $H$ is equivalent to the category of admissible representations with nonzero Iwahori-fixed vector of some $p$-adic group.
@@
 One way Lusztig expressed the uniformity in $q$ of the representation theory of the various algebras $\mathbf{H}|_{\mathbf{q}=q}$ is via the asymptotic Hecke algebra $J$.
-This is a $\mathbb{C}$-algebra (in fact, a $\mathbb{Z}$-algebra) $J$ with distinguished basis $\{t_{w}\}_{w\in\tilde{W}}$, equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$.
+This is a $\mathbb{C}$-algebra (in fact, a $\mathbb{Z}$-algebra) $J$ with distinguished basis $\{t_{w}\}_{w\in W_{\mathrm{aff}}}$, equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$.
 In this way there is a map from $J$-modules to $\mathbf{H}$-modules, and Lusztig has shown in [affineIII] and [affineIV] that when $q$ is not a root of unity (other than $1$), that the specialized map $\phi_{q}$ induces a bijection between simple $\mathbf{H}|_{\mathbf{q}=q}$-modules and simple $J$-modules, these last being defined over $\mathbb{C}$.
@@
 In this way one can write a basis element $t_{w}$ as an infinite sum
-$()^{\dagger}\circ\phi^{-1}(t_{w})=\sum_{x\in\tilde{W}}a_{x,w}T_{x},$ (1)
-where each $a_{x,w}$ is a formal Laurent series in $\mathbf{q}^{-1/2}$, and $()^{\dagger}$ is the Goldman involution.
+$^{\dagger}(-)\circ\phi^{-1}(t_{w})=\sum_{x\in W_{\mathrm{aff}}}a_{x,w}T_{x},$ (1)
+where each $a_{x,w}$ is a formal Laurent series in $\mathbf{q}^{-1/2}$, and $^{\dagger}(-)$ is the involution of $\mathbf{H}$ defined in Definition 6.
+It agrees with the Goldman involution of $\mathbf{H}$ when $G$ is simply-connected.
+In this paper we will almost exclusively work with $\phi\circ{}^{\dagger}(-)$, for reasons explained in Section 1.1.
 In light of the above, it is natural to ask how $a_{x,w}$ behaves when $\mathbf{q}$ is specialized to a root of unity.
@@
 In [BK], Braverman and Kazhdan showed essentially that the analytic meaning of the word “tempered” can be substituted into Lusztig’s results from [affineIV].
-More precisely, let $\mathbf{G}$ be a connected reductive group defined and split over a non-archimedean local field $F$ whose extended affine Weyl group is $\tilde{W}$.
+More precisely, let $\mathbf{G}$ be a connected reductive group defined and split over a non-archimedean local field $F$ whose extended affine Weyl group is $W_{\mathrm{aff}}$.
 Then in [BK], Braverman and Kazhdan define a map expressing $J$ as sitting between the Iwahori-Hecke algebra of $G=\mathbf{G}(F)$ and the Harish-Chandra Schwartz algebra $\mathcal{C}$, and propose a spectral characterization of $J$ via the operator Payley-Wiener theorem, obtaining a diagram
@@
 1.1.1 Denominators in the affine Hecke algebra and injectivity of $\eta$
-[…]
-[…]
-[…]
+In the first part of this paper we prove that $\eta$ is injective.
+Along the way, we prove in Proposition 3 that $\tilde{\phi}$ is essentially the map $\phi^{-1}$.
+Our strategy is to determine that the Schwartz functions $f_{w}$ on the $p$-adic group $G$ satisfy the statements of Theorem 2 below, and are in addition sufficiently well-behaved so as to lift to elements of a certain completion $\mathbfcal{H}^{-}$ of $\mathbf{H}$, thus defining a map
+$\phi_{1}\colon J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}.$
+We therefore obtain two maps $J\subset J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$: the inverse $^{\dagger}(-)\circ\phi^{-1}$ of Lusztig’s map, and our map $\phi_{1}$ induced by the construction in [BK].
+We prove in Proposition 3 that these maps agree, at which point Theorem 13 and the first statement of Theorem 2 follow.
 In particular, $\phi_{1}$ is injective.
@@
 Theorem 2.
-Let $\tilde{W}$ be an affine Weyl group, $\mathbf{H}$ its affine Hecke algebra over $\mathcal{A}$, and $J$ its asymptotic Hecke algebra.
+Let $W_{\mathrm{aff}}$ be an affine Weyl group, $\mathbf{H}$ its affine Hecke algebra over $\mathcal{A}$, and $J$ its asymptotic Hecke algebra.
 Let $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ be Lusztig’s map.
   1. 1.
-    For all $x,w\in\tilde{W}$, $a_{x,w}$ is a rational function of $\mathbf{q}$. The denominator of $a_{x,w}$ is independent of $x$. As a function of $w$, it is constant on two-sided cells.
+    For all $x,w\in W_{\mathrm{aff}}$, $a_{x,w}$ is a rational function of $\mathbf{q}$. The denominator of $a_{x,w}$ is independent of $x$. As a function of $w$, it is constant on two-sided cells.
   2. 2.
-    There exists $N_{\tilde{W}}\in\mathbb{N}$ such that upon writing
-    $()^{\dagger}\circ\phi^{-1}(t_{w})=\sum_{x\in\tilde{W}}a_{x,w}T_{x},$
+    There exists $N_{W_{\mathrm{aff}}}\in\mathbb{N}$ such that upon writing
+    $^{\dagger}(-)\circ\phi^{-1}(t_{w})=\sum_{x\in W_{\mathrm{aff}}}a_{x,w}T_{x},$
     we have
-    $P_{W}(\mathbf{q})^{N_{\tilde{W}}}a_{x,w}\in\mathcal{A}$
-    for all $x,w\in\tilde{W}$.
+    $P_{W}(\mathbf{q})^{N_{W_{\mathrm{aff}}}}a_{x,w}\in\mathcal{A}$
+    for all $x,w\in W_{\mathrm{aff}}$.
   3. 3.
@@
 In [D], the author proved Theorem 2 in type $\tilde{A}_{1}$, but with different conventions.
-To translate to the conventions of this paper, the reader should replace $j$ with the Goldman involution $()^{\dagger}$, and the completion of $\mathbf{H}$ with respect to the $C_{w}$ basis and positive powers of $\mathbf{q}^{1/2}$ with the completion of $\mathbf{H}$ with respect to the basis $\{(-1)^{\ell(w)}C^{\prime}_{w}\}_{w\in\tilde{W}}$ and negative powers of $\mathbf{q}^{1/2}$.
+To translate to the conventions of this paper, the reader should replace $j$ with the involution $^{\dagger}(-)$, and the completion of $\mathbf{H}$ with respect to the $C_{w}$ basis and positive powers of $\mathbf{q}^{1/2}$ with the completion of $\mathbf{H}$ with respect to the basis $\{(-1)^{\ell(w)}C^{\prime}_{w}\}_{w\in W_{\mathrm{aff}}}$ and negative powers of $\mathbf{q}^{1/2}$.
 Note also that we write $a_{x,w}$ instead of $a_{w,x}$ as in [D].
@@
 On the other hand, when $G=\mathrm{GL}_{n}$, we are able to be slightly more precise than Theorem 2, while also not appealing to [BDD].
-For this reason we treat the case $G=\mathrm{GL}_{n}$ separately as Theorem 13.
+For this reason we treat the case $G=\mathrm{GL}_{n}$ separately as Theorem 13 in Section 3.4.
 Our main tool is Harish-Chandra’s Plancherel formula for the $p$-adic group $G$ associated to $\mathbf{H}$ and the surjection of cocentres induced by Lusztig’s map $\phi$ after inverting $P_{W}(\mathbf{q})$ proved in [BDD] [^1] In fact, as proved by Bezrukavnikov-Braverman-Kazhdan in the appendix of loc.
@@
 By Theorem 2, the coefficients $a_{1,d}$ have poles at every root of $P_{W}$, for all distinguished involutions $d$ in the lowest two-sided cell $\mathbf{c}_{0}$.
-On the other hand, as we show in Example 7, there do exist cells $\mathbf{c}\neq\mathbf{c}_{0}$ such that the coefficients $a_{x,w}$ are nonsingular at certain roots of $P_{W}$, for all $w\in\mathbf{c}$ and $x\in\tilde{W}$.
+On the other hand, as we show in Example 7, there do exist cells $\mathbf{c}\neq\mathbf{c}_{0}$ such that the coefficients $a_{x,w}$ are nonsingular at certain roots of $P_{W}$, for all $w\in\mathbf{c}$ and $x\in W_{\mathrm{aff}}$.
 We encode the hope that this is no accident as
@@
 Let $\tilde{W}$ be an affine Weyl group, and let $q\in\mathbb{C}^{\times}$ be a root of $P_{W}$.
-Let $\mathbf{c}$ be a two-sided cell such that if $w\in\mathbf{c}$, then $a_{x,w}$ does not have a pole at $\mathbf{q}=q$ for any $x\in\tilde{W}$.
+Let $\mathbf{c}$ be a two-sided cell such that if $w\in\mathbf{c}$, then $a_{x,w}$ does not have a pole at $\mathbf{q}=q$ for any $x\in W_{\mathrm{aff}}$.
 Let $u=u(\mathbf{c})$.
-Then every standard module $K(u,s,\rho,q)$ has a unique simple quotient $L=L(u,s,\rho,q)$ such that $a(L)=a(E)$, where $E$ is the simple $J$ module corresponding to $K(u,s,\rho,q)$, in the notation of [affineIV].
+Let $K(u,s,\rho)$ be a standard module in the notation of [KLDeligneLanglands].
+Then the module $^{\dagger}{}^{*}K(u,s,\rho,q)$ (see Definition 6 (a), (b) and the discussion following Theorem 7) has a unique simple quotient $L=L(u,s,\rho,q)$ such that $a(L)=a(E)$, where $E$ is the simple $J$ module corresponding via $\phi$ to $(u,s,\rho)$ under [affineIV, Thm.
+4.2].
 Two such simple modules are isomorphic if and only if their corresponding triples are conjugate.
@@
 1.1.3 Application: representations with parahoric-fixed vectors
-In Section 4, we use the existence of the action of the asymptotic Hecke algebra on tempered $G$-representations to give a simple criterion for the existence of vectors fixed under a parahoric subgroup of $G$: thm existence of parahoric-fixed vectors
+In Section 4, we use the existence of the action of the asymptotic Hecke algebra on tempered $G$-representations to give a simple criterion for the existence of vectors fixed under a parahoric subgroup of $G$:
 Theorem 14.
 Let $\pi=K(u,s,\rho)$ be a simple tempered $I$-spherical representation of $G$.
-Let $\mathcal{P}$ be a parahoric subgroup of $G$ and let $w_{\mathcal{P}}$ be the longest element in the corresponding subgroup of $\tilde{W}$.
+Let $\mathcal{P}$ be a parahoric subgroup of $G$ and let $w_{\mathcal{P}}$ be the longest element in the corresponding subgroup of $W_{\mathrm{aff}}$.
 Let $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
@@
 Remark 3.
-By [affineIV, Theorem 4.8(d)] and the proof of [affineIV, Lemma 5.5], every two-sided cell contains a distinguished involution contained in a finite parabolic subgroup of $\tilde{W}$, but not every distinguished involution of a finite Coxeter group is the longest word of a parabolic subgroup, i.e.
-is of the form $w_{\mathcal{P}}$; approximately half of two-sided cells of the finite Weyl group $W\subset\tilde{W}$ do not contain any distinguished involutions contained in proper parabolic subgroups, because of the cell-preserving bijection $w\mapsto w_{0}w$.
+By [affineIV, Theorem 4.8(d)] and the proof of [affineIV, Lemma 5.5], every two-sided cell contains a distinguished involution contained in a finite parabolic subgroup of $W_{\mathrm{aff}}$, but not every distinguished involution of a finite Coxeter group is the longest word of a parabolic subgroup, i.e.
+is of the form $w_{\mathcal{P}}$; approximately half of two-sided cells of the finite Weyl group $W\subset W_{\mathrm{aff}}$ do not contain any distinguished involutions contained in proper parabolic subgroups, because of the cell-preserving bijection $w\mapsto w_{0}w$.
 For example, this happens for the second-lowest cell for $E_{8}$.
@@
 We write $|\cdot|_{F}$ for the $p$-adic absolute value on $F$; when necessary, $|\cdot|_{\infty}$ will denote the archimedean absolute value on $\mathbb{C}$.
-Let $\mathbf{G}$ be a connected reductive algebraic group defined and split over $F$, $\mathbf{A}$ a maximal torus of $\mathbf{G}$, and $X_{*}=X_{*}(\mathbf{A})$ the cocharacter lattice of $\mathbf{A}$.
+Let $\mathbf{G}$ be a connected reductive algebraic group defined and split over $F$, $\mathbf{A}$ a maximal $F$-split torus of $\mathbf{G}$, and $X_{*}=X_{*}(\mathbf{A})$ the cocharacter lattice of $\mathbf{A}$.
+Let $\pi_{1}(G)=X_{*}/\mathbb{Z}\Phi^{\vee}$ be the fundamental group, the quotient of the cocharacter lattice by the coroot lattice.
 Let $\mathbf{N}$ be unipotent radical of a chosen Borel subgroup $\mathbf{B}$, so that $\mathbf{B}=\mathbf{A}\mathbf{N}$.
-Let $W$ be the finite Weyl group of $\mathbf{G}$, and $\tilde{W}=W\ltimes X_{*}(\mathbf{A})$ be the extended affine Weyl group.
-Write $S$ for the set of simple reflections in $\tilde{W}$.
+Let $W$ be the finite Weyl group of $\mathbf{G}$, and $W_{\mathrm{aff}}=W\ltimes X_{*}(\mathbf{A})$ be the extended affine Weyl group.
+Write $S$ for the set of simple reflections in $W_{\mathrm{aff}}$.
 Let $\mathbf{G}^{\vee}$ be the Langlands dual group of $\mathbf{G}$, taken over $\mathbb{C}$.
@@
 Let $K$ be the maximal compact subgroup $\mathbf{G}(\mathcal{O})$.
-Also let $I$ be the Iwahori subgroup of $G$ that is the preimage of $\mathbf{B}(\mathbb{F}_{q})$.
+Also let $I$ be the Iwahori subgroup of $G$ that is the preimage of $\mathbf{B}(\mathbb{F}_{q})$ in $K$.
 We sometimes write $P_{\mathbf{G}/\mathbf{B}}$ for $P_{W}$, as this polynomial is also the Poincaré polynomial of the flag variety $(\mathbf{G}/\mathbf{B})(\mathbb{C})$.
-We write $\mathbf{H}$ for the affine Hecke algebra of $\tilde{W}$.
+We write $\mathbf{H}$ for the affine Hecke algebra of $W_{\mathrm{aff}}$.
 It is a unital associative algebra over the ring $\mathcal{A}=\mathbb{C}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ (in fact, it is defined over $\mathbb{Z}[\mathbf{q}^{\frac{1}{2}},\mathbf{q}^{-\frac{1}{2}}]$ but we will work over $\mathbb{C}$ to avoid having to introduce extra notation later), where $\mathbf{q}^{\frac{1}{2}}$ is a formal variable.
@@
 We write $\theta_{\lambda}$ for the generators of the Bernstein subalgebra.
-Recall from [KL79] the two Kazhdan-Lusztig bases $\{C_{w}\}_{w\in\tilde{W}}$ and $\{C^{\prime}_{w}\}_{w\in\tilde{W}}$ of $\mathbf{H}$ and write $C_{x}C_{y}=\sum_{z\in\tilde{W}}h_{x,y,z}C_{z}$.
+Recall from [KL79] the two Kazhdan-Lusztig bases $\{C_{w}\}_{w\in W_{\mathrm{aff}}}$ and $\{C^{\prime}_{w}\}_{w\in W_{\mathrm{aff}}}$ of $\mathbf{H}$, where
+$C^{\prime}_{w}=\mathbf{q}^{-\frac{\ell(w)}{2}}\sum_{x\leq w}P_{x,w}(\mathbf{q})T_{x}$
+for the Kazhdan-Lusztig polynomials $P_{x,w}$.
+Write $C_{x}C_{y}=\sum_{z\in W_{\mathrm{aff}}}h_{x,y,z}C_{z}$.
 The inverse Kazhdan-Lusztig polynomials $Q_{y,x}$ are the unique family of polynomials satisfying
 $T_{x}=\sum_{y\leq x}(-1)^{\ell(x)-\ell(y)}\mathbf{q}^{\frac{\ell(y)}{2}}Q_{y,x}(\mathbf{q})C^{\prime}_{y},$
-or equivalently,
+or equivalently, satisfying
 $\sum_{z\leq y\leq x}(-1)^{\ell(x)-\ell(y)}Q_{y,x}(\mathbf{q})P_{z,y}(\mathbf{q})=\delta_{z,x}$
 along with some restrictions on their degrees. For example, we shall use in Section 3.3.2 that $\deg Q_{y,x}\leq\frac{1}{2}(\ell(x)-\ell(y)-1)$. See [BjornerBrenti] for further exposition.
+If
+$\varphi\colon(W_{\mathrm{aff}},S)\to(W_{\mathrm{aff}},S)$
+is a Coxeter group automorphism of $W_{\mathrm{aff}}$, then
+$T_{w}\mapsto T_{\varphi(w)}$
+is an algebra automorphism of $\mathbf{H}$ commuting with the bar involution, and therefore given equivalently by
+$C^{\prime}_{w}\mapsto C^{\prime}_{\varphi(w)}$
+and
+$C_{w}\mapsto C_{\varphi(w)}.$ (3)
 It is well-known that there is an isomorphism of associative $\mathbb{C}$-algebras
@@
 Definition 1.
-Lusztig’s a-function $a\colon\tilde{W}\to\mathbb{Z}_{\geq 0}$ is defined such that $a(w)$ is the minimal value such that $\mathbf{q}^{\frac{a(w)}{2}}h_{x,y,w}\in\mathcal{A}^{+}=\operatorname{Spec}\mathbb{C}[\mathbf{q}^{1/2}]$ for all $x,y\in\tilde{W}$.
-The $a$-function is constant of two-sided cells of $\tilde{W}$.
-Obviously, $a(1)=1$, and the $a$-function obtains its maximum, equal to the number of positive roots, on the two-sided cell containing the longest word $w_{0}\in W$.
-In general, under the bijection between two-sided cells $\mathbf{c}$ of $\tilde{W}$ and unipotent conjugacy classes $u=u(\mathbf{c})$ in $G^{\vee}$ of [affineIV], we have
+Lusztig’s a-function $a\colon\tilde{W}\to\mathbb{Z}_{\geq 0}$ is defined such that $a(w)$ is the minimal value such that $\mathbf{q}^{\frac{a(w)}{2}}h_{x,y,w}\in\mathcal{A}^{+}=\mathbb{C}[\mathbf{q}^{1/2}]$ for all $x,y\in\tilde{W}$.
+The $a$-function is constant of two-sided cells of $W_{\mathrm{aff}}$.
+Obviously, $a(1)=0$, and the $a$-function obtains its maximum, equal to the number of positive roots, on the two-sided cell containing the longest word $w_{0}\in W$.
+In general, under the bijection between two-sided cells $\mathbf{c}$ of $W_{\mathrm{aff}}$ and unipotent conjugacy classes $u=u(\mathbf{c})$ in $G^{\vee}$ of [affineIV], we have
 $a(\mathbf{c})=\dim_{\mathbb{C}}(\mathcal{B}_{u}^{\vee}),$
-where $\mathcal{B}_{u}^{\vee}$ is the Springer fibre of $u$. We have $a(w)\leq\ell(w)$ for all $w\in\tilde{W}$.
+where $\mathcal{B}_{u}^{\vee}$ is the Springer fibre of $u$. We have $a(w)\leq\ell(w)$ for all $w\in W_{\mathrm{aff}}$.
 In [affineII], Lusztig defined an associative algebra $J$ over $\mathbb{C}$ equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ which becomes an isomorphism after taking a certain completion, to be recalled in Section 3.3.1, of both sides.
-As an abelian group, $J$ is free with basis $\{t_{w}\}_{w\in\tilde{W}}$.
-The structure constants of $J$ are obtained from those of $\mathbf{H}$ written in the $\{C_{w}\}_{w\in\tilde{W}}$-basis under the following procedure: first, the integer $\gamma_{x,y,z}$ is defined by the condition
+As a $\mathbb{C}$-vector space, $J$ has basis $\{t_{w}\}_{w\in W_{\mathrm{aff}}}$.
+The structure constants of $J$ are obtained from those of $\mathbf{H}$ written in the $\{C_{w}\}_{w\in W_{\mathrm{aff}}}$-basis under the following procedure: first, the integer $\gamma_{x,y,z}$ is defined by the condition
 $\mathbf{q}^{\frac{a(z)}{2}}h_{x,y,z^{-1}}-\gamma_{x,y,z}\in\mathbf{q}\mathcal{A}^{+}.$
 One then defines
-$t_{x}t_{y}=\sum_{z\in\tilde{W}}\gamma_{x,y,z}t_{z^{-1}}.$
+$t_{x}t_{y}=\sum_{z\in W_{\mathrm{aff}}}\gamma_{x,y,z}t_{z^{-1}}.$
 Surprisingly, this defines a unital associative algebra with unit
@@
 where $\mathcal{D}$ is the (finite) set of distinguished involutions [affineII].
-The elements $t_{d}$ are orthogonal idempotents in $J$, which decomposes as a direct sum indexed by the two-sided cells of $\tilde{W}$ in the sense of [affineI].
+The elements $t_{d}$ are orthogonal idempotents in $J$, which decomposes as a direct sum indexed by the two-sided cells of $W_{\mathrm{aff}}$ in the sense of [affineI].
 Each left cell, again in the sense of op.
@@
 given by
-$\phi(C_{w})=\sum_{z\in W,~{}d\in\mathcal{D},~{}a(z)=a(d)}h_{x,d,z}t_{z}.$
+$\phi(C_{w})=\sum_{z\in W,~d\in\mathcal{D},~a(z)=a(d)}h_{w,d,z}t_{z}.$
 Write $\phi_{q}$ for the specialization of this map when $\mathbf{q}=q$. It is known [affineIII, Proposition 1.7] that $\phi_{q}$ is injective for all $q\in\mathbb{C}^{\times}$.
+Lemma 1.
+Let
+$\varphi\colon(W_{\mathrm{aff}},S)\to(W_{\mathrm{aff}},S)$
+be a Coxeter group automorphism. Then
+  1. 1.
+    The map $t_{w}\mapsto t_{\varphi(w)}$ defines a based ring automorphism of $J$, which we also denote $\varphi$.
+  2. 2.
+    The map $\phi$ is $\varphi$-linear, in the sense that it commutes with the automorphism from the first statement and the automorphism (3).
+Proof.
+The first statement is [LX, 2.2(g)].
+For the second statement, note that $\varphi$ acts on $\mathcal{D}$ and that
+$\phi(C_{\varphi(w)})=\sum_{d_{1}\sim_{L}z_{1}}h_{\varphi(w),d_{1},z_{1}}t_{z_{1}}$
+and
+$\varphi\left(\phi(C_{w})\right)=\sum_{d_{2}\sim_{L}z_{2}}h_{w,d_{2},z_{2}}t_{\varphi(z_{2})}.$
+Now if $z_{1}=\varphi(z_{2})$, then $d_{1}=\varphi(d_{2})$, and
+$h_{\varphi(w),d_{1},z_{1}}=h_{\varphi(w),\varphi(d_{2}),\varphi(d_{2})}=h_{w,d_{2},z_{2}}$
+by (3). ∎
 1.4.1 Deformations of the group ring
-Upon setting $\mathbf{q}=1$, $\mathbf{H}|_{\mathbf{q}=1}$ is isomorphic to $\mathbb{C}[\tilde{W}]$, and so $\mathbf{H}$ is a deformation of the group algebra of $\tilde{W}$.
+Upon setting $\mathbf{q}=1$, $\mathbf{H}|_{\mathbf{q}=1}$ is isomorphic to $\mathbb{C}[\tilde{W}$], and so $\mathbf{H}$ is a deformation of the group algebra of $W_{\mathrm{aff}}$.
 Let, temporarily, $W$ be any finite Coxeter group.
@@
 Let $W=\left\langle 1,s\,\middle|\,s^{2}=1\right\rangle$ be the Weyl group of type $A_{1}$.
-The Kazhdan-Lusztig $C_{w}$-basis elements are $C_{1}=T_{1}$ and $C_{s}=\mathbf{q}^{-\frac{1}{2}}T_{s}-\mathbf{q}^{\frac{1}{2}}T_{1}$.
+The Kazhdan-Lusztig $C_{w}$-basis elements are $C_{1}=T_{1}$ and $C_{s}=\mathbf{q}^{-\frac{1}{2}}T_{s}-\mathbf{q}^{\frac{1}{2}}T_{1}$, and $\mathcal{D}=W$ in this case.
 There are two two-sided cells in $W$, and one can easily check that
@@
 However, Kazhdan-Lusztig provide the following algebraic generalization.
-Let $\kappa=\mathbb{C}(\mathbf{q}^{-\frac{1}{2}})$ and $\bar{\kappa}$ be the algebraic closure of $\kappa$.
-We write $\mathbf{H}_{\bar{\kappa}}$ for $\mathbf{H}\otimes_{\mathcal{A}}\bar{\kappa}$ and recall another definition of Lusztig’s from [affineIV].
+Let $\mathcal{K}=\mathbb{C}(\mathbf{q}^{-\frac{1}{2}})$ and $\bar{\mathcal{K}}$ be the algebraic closure of $\mathcal{K}$.
+We write $\mathbf{H}_{\bar{\mathcal{K}}}$ for $\mathbf{H}\otimes_{\mathcal{A}}\bar{\mathcal{K}}$ and recall another definition of Lusztig’s from [affineIV].
 Definition 2.
-Let $M$ be a $H_{\bar{\kappa}}$-module finite-dimensional over $\bar{\kappa}$.
+Let $M$ be a $H_{\bar{\mathcal{K}}}$-module finite-dimensional over $\bar{\mathcal{K}}$.
 Say that $m\in M$ is an eigenvector if $\theta_{x}\cdot m=\chi_{m}(x)m$ for all dominant $x$ in $X_{*}$.
-As $\chi_{m}$ is a character of the coweight lattice, it corresponds to an element $\sigma_{m}\in\mathbf{A}^{\vee}(\bar{\kappa})$ in the sense that, for all cocharacters $x$ of $A$, we have
+As $\chi_{m}$ is a character of the coweight lattice, it corresponds to an element $\sigma_{m}\in\mathbf{A}^{\vee}(\bar{\mathcal{K}})$ in the sense that, for all cocharacters $x$ of $A$, we have
 $\chi_{m}(x)=x(\sigma_{m})$
@@
 $\phi^{\prime}\colon\mathrm{SL}_{2}(\mathbb{C})\to Z_{G^{\vee}}^{0}(s^{\prime})$
-such that for all eigenvectors $m$ of $M$, the element $\sigma_{m}$ is $\mathbf{G}^{\vee}(\bar{\kappa})$-conjugate to
+such that for all eigenvectors $m$ of $M$, the element $\sigma_{m}$ is $\mathbf{G}^{\vee}(\bar{\mathcal{K}})$-conjugate to
 $\phi^{\prime}(\mathrm{diag}(\mathbf{q}^{1/2},\mathbf{q}^{-1/2}))s^{\prime},$
-where by abuse of notation we have written $\phi^{\prime}$ again for the base-change to $\bar{\kappa}$.
+where by abuse of notation we have written $\phi^{\prime}$ again for the base-change to $\bar{\mathcal{K}}$.
 The idea of the name of the definition is that $s^{\prime}\in\mathbf{G}^{\vee}(\mathbb{C})$ is a “constant element” not depending on $\mathbf{q}$.
 Next is a generalization of Casselman’s criterion, which as such, requires a choice of dominant weights.
-The definitions in both [KLDeligneLanglands] and [affineIV] are both given under the choice of positive roots to be those associated to the opposite Borel (the so-called geometric choice of dominance).
-On the other hand, for purposes of harmonic analysis, Casselman’s criterion as usually stated uses the standard choice of positive roots.
-Following [affineIV], we choose a morphism of groups $V\colon\bar{\kappa}^{\times}\to\mathbb{R}$ such that $V(\mathbf{q}^{-\frac{1}{2}})=1$ and $V(a\mathbf{q}^{-\frac{1}{2}}+b)=0$ for all $a\in\mathbb{C}$, $b\in\mathbb{C}^{\times}$.
+Following [KLDeligneLanglands] (see the proof of Prop.
+1.6 of loc.
+cit.) and [affineIV, Section 1.6], we choose the positive roots to be those occurring in $\mathfrak{g}/\mathfrak{b}$.
+Following [affineIV], we choose a morphism of groups $V\colon\bar{\mathcal{K}}^{\times}\to\mathbb{R}$ such that $V(\mathbf{q}^{\frac{1}{2}})=1$ and $V(a\mathbf{q}^{\frac{1}{2}}+b)=0$ for all $a\in\mathbb{C}$, $b\in\mathbb{C}^{\times}$.
 Definition 3 ([affineIV], c.f. [KLDeligneLanglands]).
-Let $M$ be any finite-dimensional $\mathbf{H}_{\bar{\kappa}}$-module and fix the geometric notion of dominance for $X_{*}(\mathbf{A})$.
-We say that $M$ is $V$-geometrically anti-tempered if all eigenvalues $\nu$ of $\theta_{\lambda}$ for all geometrically dominant $\lambda\in X_{*}(\mathbf{A})$ satisfy $V(\nu)\geq 0$.
-We say that $M$ is $V$-geometrically anti-square-integrable if $V(\nu)>0$.
-When $\mathbf{q}$ is specialized a prime power $q\in\mathbb{C}^{\times}$ and $V(z)=\log|z|$, we recover that our anti-tempered for the geometric choice of dominance modules are tempered in the traditional analytic sense after twisting by the Goldman involution, and likewise for square-integrability, by the Casselman criterion.
+Let $M$ be any finite-dimensional $\mathbf{H}_{\bar{\mathcal{K}}}$-module.
+We say that $M$ is $V$-tempered if all eigenvalues $\nu$ of $\theta_{\lambda}$ for all dominant $\lambda\in X_{*}(\mathbf{A})$ satisfy $V(\nu)\leq 0$.
 The representation theory of $J$ is very well understood.
@@
 Let $E$ be a $J$-module.
-Then $E\otimes_{\mathbb{C}}\kappa$ is a $J\otimes_{\mathbb{C}}\kappa$-module.
-Hence $\mathbf{H}_{\kappa}$ acts on $E$ via $\phi$.
-Denote the resulting $\mathbf{H}_{\kappa}$ module by ${}^{\phi}E$.
+Then $E\otimes_{\mathbb{C}}\mathcal{K}$ is a $J\otimes_{\mathbb{C}}\mathcal{K}$-module.
+Hence $\mathbf{H}_{\mathcal{K}}$ acts on $E$ via $\phi$.
+Denote the resulting $\mathbf{H}_{\mathcal{K}}$ module by $^{\phi}E$.
 1.5.1 Involutions on $\mathbf{H}$
+For $x\in W_{\mathrm{aff}}$, let $\omega(x)\in\pi_{1}(G)$ label the $W\ltimes\mathbb{Z}\Phi^{\vee}$-coset of $W_{\mathrm{aff}}$ containing $x$, and write $\omega(x)=\omega(x)_{f}\omega(x)_{t}\in W\ltimes X_{*}$.
+Then every $y\leq x$ is also in the coset of $\omega(x)$.
 Definition 5.
@@
 The $j$-involution exchanges the $\{C_{w}\}$ and $\{C_{w}^{\prime}\}$-bases [KL79].
-The reason for our choice of conventions, which match those of [affineIII] and [affineIV], is the presence of the Goldman involution and its exchange of temperedness and anti-temperedness in the relationship between $H$-modules and $J$-modules; see Theorem 4 and Lemma 1 below.
+The reason for our choice of conventions, which differ slightly from those of [affineIII] and [affineIV], is the presence of the involution $\dagger$ and its exchange of temperedness and anti-temperedness in the relationship between $H$-modules and $J$-modules; see Theorem 4 and Lemma 2 below.
 Definition 6.
-The Goldman involution $h\mapsto h^{\dagger}$ is the $\mathcal{A}$-algebra involution of $\mathbf{H}$ defined by setting
-$T_{w}^{\dagger}=(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1}.$
-Given an $\mathbf{H}$ (or $\mathbf{H}_{\kappa}$)-module $M$, define ${}^{\dagger}M$ to be the same vector space with the $\mathbf{H}$-action twisted by this involution.
-Let $\Delta$ be the usual positive system determined by $\mathbf{B}$ and let $\Delta^{-}$ be the geometric system. Then we have
+  1. (a)
+    Define the $\mathcal{A}$-algebra involution $^{\dagger}(-)$ of $\mathbf{H}$ by setting
+    $^{\dagger}T_{w}=(-1)^{\ell(w_{f})}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1},~~~w=w_{f}\lambda\in W_{\mathrm{aff}}.$
+    Note that the sign factor depends only on $w_{f}$.
+  2. (b)
+    Let $h\mapsto{}^{*}h$ be the $\mathcal{A}$-involution defined in terms of the Bernstein presentation of $\mathbf{H}$: by
+    $^{*}T_{w}=(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1},~~~w\in W$
+    and
+    $^{*}\theta_{\lambda}=\theta_{\lambda}^{-1},~~~\lambda\in X_{*}.$
+  3. (c)
+    ([PrasadDuke, Section 5.a].) Let $\bar{\kappa}$ be the $\mathcal{A}$-linear involution defined by
+    $\bar{\kappa}(T_{w})=T_{\bar{\kappa}(w)}$
+    induced by the Coxeter group automorphism given by $\bar{\kappa}(s)=w_{0}sw_{0}$ for $s\in S\setminus\left\{s_{0}\right\}$ and $\bar{\kappa}(s_{0})=s_{0}$, so that $\bar{\kappa}(\lambda)=-w_{0}(\lambda)$ for $\lambda\in X_{*}$; equivalently
+    $\bar{\kappa}(T_{w})=T_{w_{0}ww_{0}^{-1}},~~~w\in W,~~~\bar{\kappa}(\theta_{\lambda})=\theta_{-w_{0}(\lambda)},~~~\lambda\in X_{*}.$
+  4. (d)
+    ([Barbasch-Moy, Section 5].) Let $h\mapsto{}^{\bullet}h$ be the $\mathcal{A}$-linear anti-involution defined by
+    $^{\bullet}T_{w}=T_{w^{-1}},~~~w\in W,~~~{}^{\bullet}\theta_{\lambda}=\theta_{\lambda},~~\lambda\in X_{*}.$
+  5. (e)
+    ([Barbasch-Moy, Section 5].) Let $h\mapsto{}^{\star}h$ be the $\mathcal{A}$-linear anti-involution defined by
+    $^{\star}T_{w}=T_{w^{-1}},~~~w\in W,~~~{}^{\star}\theta_{\lambda}=T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}T_{w_{0}}^{-1},~~\lambda\in X_{*}.$
+By [OpdamSpectral, Prop.
+2.9], $^{\star}T_{w}=T_{w^{-1}}$ for $w\in W_{\mathrm{aff}}$.
+When $G$ is simply-connected, the involution $\dagger$ agrees with the Goldman involution of $\mathbf{H}$.
+Now we have
 Lemma 2.
 We have
-  1. 1.
-    $\overline{h^{\dagger}}=(\bar{h})^{\dagger}=j(h)$
-    as $\mathcal{A}$-antilinear involutions. In particular, $C_{w}^{\dagger}=j(C_{w})=(-1)^{\ell(w)}C^{\prime}_{w}$.
-  2. 2.
-    Let $\lambda$ be dominant for $\Delta$ and let $\theta^{\Delta}_{\lambda}$ and $\theta^{\Delta^{-}}_{\lambda}$ denote Bernstein elements with respect to $\Delta$ and $\Delta^{-}$, respectively.
-Then
-    ${\theta^{\Delta}_{\lambda}}^{\dagger}=(-1)^{\ell(\lambda)}\theta^{\Delta^{-}}_{\lambda}.$
+  1. (a)
+    We have $^{\dagger}C_{x}=(-1)^{\ell(\omega(x)_{f})}j(C_{w})=(-1)^{\ell(x)+\ell(\omega(x)_{f})}C^{\prime}_{x}$ for all $x\in W_{\mathrm{aff}}$.
+  2. (b)
+    We have $^{\dagger}\theta_{\lambda}=T_{w_{0}}\theta_{-\bar{\kappa}(\lambda)}T_{w_{0}}^{-1}=T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}^{-1}T_{w_{0}}^{-1}$.
+  3. (c)
+    We have ${}^{\bullet}{}^{\star}{}^{\dagger}(-)={}^{*}(-)$ as involutions on $\mathbf{H}$.
+  4. (d)
+    We have $^{\bullet}(-)\circ{}^{\star}(-)=T_{w_{0}}^{-1}\bar{\kappa}(-)T_{w_{0}}$ as automorphisms of $\mathbf{H}$.
+  5. (e)
+    We have the commutative diagram
+[…]
 Proof.
-We compute
-$\overline{\sum_{x}b_{x}T_{x}^{\dagger}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\overline{\mathbf{q}^{\ell(x)}\bar{T}_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}=j\left(\sum_{x}b_{x}T_{x}\right)$
+By definition, if $x=\omega x^{\prime}$ for $x^{\prime}\in W\ltimes\mathbb{Z}\Phi^{\vee}$ and $\omega\in\pi_{1}(G)$, then $C^{\prime}_{x}=T_{\omega}C^{\prime}_{x^{\prime}}$ and $C_{x}=T_{\omega}C_{x^{\prime}}$.
+Further, if $\omega=\omega_{f}\omega_{t}\in W_{\mathrm{aff}}$ has $\ell(\omega)=0$, then $^{\dagger}T_{\omega}=(-1)^{\ell(\omega_{f})}T_{\omega}$.
+Therefore it suffices to show that
+$\overline{{}^{\dagger}h}={}^{\dagger}\left(\bar{h}\right)=j(h)$
+as $\mathcal{A}$-antilinear automorphisms of $\mathbf{H}$ for $G$ simply connected.
+In this case, $\ell(\lambda)\in 2\mathbb{Z}$ for any dominant $\lambda\in X_{*}$, whence $^{\dagger}(-)$ agrees with the involution $T_{w}\mapsto(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1}$, $w\in W\ltimes\mathbb{Z}\Phi^{\vee}$.
+Therefore we compute
+$\overline{\sum_{x}b_{x}{}^{\dagger}T_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\overline{\mathbf{q}^{\ell(x)}\bar{T}_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}=j\left(\sum_{x}b_{x}T_{x}\right)$
 whereas
-$\left(\overline{\sum_{x}b_{x}T_{x}}\right)^{\dagger}=\left(\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}^{\dagger}\right)^{\dagger}.$
-Thus we have $\bar{C_{w}^{\dagger}}=(-1)^{\ell(w)}C^{\prime}_{w}$, whence the first claim.
-Next,
-${\theta^{\Delta}_{\lambda}}^{\dagger}=(-1)^{\ell(\lambda)}\mathbf{q}^{-\frac{\ell(\lambda)}{2}}\mathbf{q}^{\ell(\lambda)}T_{-\lambda}^{-1}=(-1)^{\ell(\lambda)}\theta^{\Delta^{-}}_{\lambda},$
-noting that $-\lambda$ is dominant with respect to $\Delta^{-}$ and so can appear as a label with respect to $\Delta^{-}$. ∎
+$^{\dagger}\left(\overline{\sum_{x}b_{x}T_{x}}\right)={}^{\dagger}\left(\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}{}^{\dagger}T_{x}\right).$
+Thus we have $\overline{{}^{\dagger}C_{w}}=(-1)^{\ell(w)}C^{\prime}_{w}$ for $G$ simply-connected, whence (a).
+It suffices to prove (b) for $\lambda$ dominant, in which case
+$^{\dagger}\theta_{\lambda}=\mathbf{q}^{-\frac{\ell(\lambda)}{2}}{}^{\dagger}T_{\lambda}=\mathbf{q}^{\frac{\ell(\lambda)}{2}}T_{-\lambda}^{-1}=\mathbf{q}^{\frac{\ell(\lambda)}{2}}\left({}^{\star}T_{\lambda}\right)^{-1}=\left({}^{\star}\theta_{\lambda}\right)^{-1}=\left(T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}T_{w_{0}}^{-1}\right)^{-1}=T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}^{-1}T_{w_{0}}^{-1},$
+where we used the equivalence of the definitions in Definition 6 (e). This shows (b).
+For (c), we again compute, for $\lambda$ dominant, that
+$^{\bullet}{}^{\star}{}^{\dagger}\theta_{\lambda}={}^{\bullet}{}^{\star}\left(T_{w_{0}}\theta_{-\bar{\kappa}(\lambda)}T_{w_{0}}^{-1}\right)={}^{\bullet}\theta_{-\lambda}=\theta_{-\lambda}.$
+Agreement on $T_{w}$ for $w\in W$ follows from the fact that $^{\bullet\star}(-)$ is the identity on the finite Hecke algebra. This shows (c).
+On the Bernstein subalgebra, part (d) follows from the definitions.
+On the finite Hecke algebra, we must show that the right hand side is the identity automorphism.
+For $s$ a finite reflection, write $w_{0}sw_{0}=\bar{\kappa}(s)$, again a finite reflection, so that $\bar{\kappa}(s)w_{0}s=w_{0}$, and
+$T_{w_{0}}T_{s}T_{w_{0}}^{-1}=T_{w_{0}}T_{s}T_{s}^{-1}T_{w_{0}s}^{-1}=T_{w_{0}}T_{w_{0}s}^{-1}=T_{\bar{\kappa}(s)}T_{w_{0}s}T_{w_{0}s}^{-1}=T_{\bar{\kappa}(s)}.$
+Therefore on the finite Hecke algebra, $\bar{\kappa}$ is given by conjugation by $T_{w_{0}}$.
+Commutativity of the right square in (e) follows from Lemma 1, and commutativity of the left square follows from (c) and (d).
+Given an $\mathbf{H}$ (or $\mathbf{H}_{\mathcal{K}}$)-module $M$, define $^{\dagger}M$ to be the same vector space with the $\mathbf{H}$-action twisted by this involution, and likewise for other automorphisms.
+Corollary 1.
+Twisting by $^{\dagger}(-)$ exchanges tempered and anti-tempered $H$-modules.
+Proof.
+Immediate from Lemma 2 (b). ∎
+In fact, more is true: by [Kato, Thm. 2], $^{\dagger}M$ is the Aubert-Zelevinski dual of $M$.
+Remark 4.
+There is another natural auto-equivalence of the category of admissible representations of $G$ which exchanges tempered and anti-tempered representations, namely Bernstein’s cohomlogical duality.
+This functor differs from Aubert-Zelevinksi duality by the contragredient [SchneiderStuhler] (c.f.
+[PrasadNori]).
+On semisimple $H$-modules $\pi$, we have $^{\bar{\kappa}}\pi=\tilde{\pi}$, by [PrasadDuke, Prop.
+6.3].
+Therefore twist by the involution $^{*}(-)$, which manifestly exchanges tempered and anti-tempered representations, induces on semisimple modules the cohmological duality.
+However, it does not do so in general, because $\bar{\kappa}$ does not induce the contragredient in general.
+Indeed, for $G=\mathrm{SL}_{2}(F)$, $\bar{\kappa}=\mathrm{id}$ but not all non-unitary principal series of $G$ are self-dual.
+Therefore the operation $^{*}(-)$ is less natural than $^{\dagger}(-)$; this is perhaps reflected by the fact the formulas in Theorem 2 are nicer than those for $(\phi\circ{}^{*}(-))^{-1}(t_{w})$, which are related to those of the Theorem by Lemma 2 (e).
 We now summarize the relationship between representations of $\mathbf{H}$ and of $J$.
 Theorem 4 ([affineIV], Prop. 2.11, Thm. 4.2, Prop. 4.4).
-Fix the geometric notion of dominance for $X_{*}(\mathbf{A})$. Then there are bijections of sets
+There are bijections of sets
 […]
-Moreover, for a simple $J$-module $E$,
+where $K(u,s,\rho)$ is a standard module as in [KLDeligneLanglands]. Moreover, for a simple $J$-module $E$,
   1. 1.
@@
   2. 2.
-    There is a unique two-sided cell $\mathbf{c}=\mathbf{c}(E)$ of $\tilde{W}$ such that $\mathrm{trace}\left({E}\,,{t_{w}}\right)\neq 0$ implies $w\in\mathbf{c}$.
+    There is a unique two-sided cell $\mathbf{c}=\mathbf{c}(E)$ of $W_{\mathrm{aff}}$ such that $\mathrm{trace}\left({E}\,,{t_{w}}\right)\neq 0$ implies $w\in\mathbf{c}$.
   3. 3.
@@
     where $M\simeq{}^{\phi}E$.
-In particular, $\mathrm{trace}\left({E}\,,{t_{w}}\right)$ is independent of $\mathbf{q}$, and upon specializing $\mathbf{q}=q$ a prime power, will be a regular function in the twisting character in the setting of the Payley-Weiner theorem for the Iwahori-Hecke algebra of the $p$-adic group $G$, as we will explain in greater detail below.
-That is, in our conventions, simple tempered $\mathbf{H}|_{\mathbf{q}=q}$-modules in the sense of harmonic analysis will extend to simple $J$-modules via ϕ()italic-$\phi\circ()^{\dagger}$.
-We will comment even further on the necessity of twisting by the Goldman involution in Section 2.4.
-1.6 Acknowledgements
-The author thanks Alexander Braverman for introducing him to this material and the problem, for suggesting the use of the Plancherel formula, and for helpful conversations.
-The author also thanks Roman Bezrukavnikov, Itai Bar-Natan, Gurbir Dhillon, Malors Espinosa-Lara, Julia Gordon, George Lusztig, and Matthew Sunohara for helpful conversations.
-The author thanks Maarten Solleveld for explaining the generalization in [Solleveld] of the results of [fomFormalDegUnip], suggesting the reference [APM], and for careful reading of an early version of this manuscript.
-Some of this work was completed at the Max Planck Institute in Mathematics for Bonn, the directors and staff of which the author thanks for their hospitality. This research was supported by NSERC.
+In particular, $\mathrm{trace}\left({E}\,,{t_{w}}\right)$ is independent of $\mathbf{q}$, and upon specializing $\mathbf{q}=q$ a prime power, will be a regular function in the twisting character in the setting of the Payley-Wiener theorem for the Iwahori-Hecke algebra of the $p$-adic group $G$, as we will explain in greater detail below.
+We will comment even further in Section 2.4 on the necessity of twisting by some $\mathbf{H}$-involution exchanging tempered and anti-tempered modules, and how it is sufficient to twist $\phi$ by either $(-)^{*}$ or $(-)^{\dagger}$, but that the latter twist leads to nicer formulas.
 2 Harish-Chandra’s Plancherel formula
@@
 2.1 Tempered and discrete series representations
+We take our definitions of discrete series and tempered representations from [Waldspurger, III.1] and [Waldspurger, III.2] respectively.
+By parabolic induction we always understand normalized induction.
 Definition 7.
-A smooth representation $\pi$ of $G$ is tempered if $\pi$ is unitary, of finite length, and all matrix coefficients of $\pi$ belong to $L^{2+\epsilon}(G/Z(G))$ for all $\epsilon>0$.
+A smooth admissible representation $\omega$ of $G$ belongs to the discrete series if $\omega$ admits a unitary central character and all matrix coefficients of $\omega$ are square-integrable modulo $Z(G)$.
+We write $\mathcal{E}_{2}(G)$ for the space of irreducible discrete series, and $\mathcal{E}_{2}(G)^{I}$ for the space of irreducible discrete series with nontrivial Iwahori-fixed vectors.
+Let $v$ be the $K$-fixed vector in the self-contragredient representation $\operatorname{Ind}_{B}^{G}(\mathrm{triv})^{K}$ such that $v(1)=1$.
+Define $\Xi(g)=\left\langle\pi(g)v,v\right\rangle$ to be the corresponding matrix coefficient.
+Definition 8.
+A smooth function $f$ on $G$ is tempered if there is $C>0$ and $r\in\mathbb{R}$ such that
+$|f(g)|\leq C\Xi(g)\left(1+\log\left\|g\right\|\right)^{r},$
+where $\left\|g\right\|\geq 1$ is defined as in [Waldspurger, p.242].
+Definition 9.
+A smooth admissible representation $\pi$ of $G$ is tempered if all its matrix coefficients are tempered functions in the sense above.
 We write $\mathcal{M}_{t}(G)$ for the category of tempered representations of $G$.
-In particular, the central character of such a representation takes values in the circle group $\mathbb{T}\subset\mathbb{C}^{\times}$.
-Definition 8.
-A smooth finite-length representation $\omega$ of $G$ belongs to the discrete series if all matrix coefficients of $\omega$ are square-integrable modulo $Z(G)$.
-We write $\mathcal{E}_{2}(G)$ for the space of unitary discrete series, and $\mathcal{E}_{2}(G)^{I}$ for the space of unitary discrete series with nontrivial Iwahori-fixed vectors.
+If a tempered representation admits a central character, the central character takes values in the circle group $\mathbb{T}\subset\mathbb{C}^{\times}$.
 The tempered representations are built from the discrete series according to the following theorem of Harish-Chandra, as related in [Waldspurger, Prop. III.4.1].
@@
 Every simple tempered representation is a direct summand of a representation of this form.
-By induction we always understand normalized induction.
 2.1.1 Formal degrees of discrete series representations
@@
 This normalization gives in our setting $\mu_{\mathrm{HII}}(K)=q^{\dim\mathbf{G}}\#\mathbf{G}(\mathbb{F}_{q})$.
-Hence, noting that #$\#\mathbf{G}(\mathbb{F}_{q})=P_{\mathbf{G}/\mathbf{B}}(q)\cdot\#\mathbf{B}(\mathbb{F}_{q})$ and that, as $\mathbb{F}_{q}$ is perfect, #$\#B(\mathbb{F}_{q})$ is a polynomial in $q$, we have
+Hence, noting that $\mathbf{G}(\mathbb{F}_{q})=P_{\mathbf{G}/\mathbf{B}}(q)\cdot\#\mathbf{B}(\mathbb{F}_{q})$ and that, as $\mathbb{F}_{q}$ is perfect, $B(\mathbb{F}_{q})$ is a polynomial in $q$, we have
 $\mu_{I}=\frac{1}{q^{\dim\mathbf{G}}\#\mathbf{B}(\mathbb{F}_{q})}\mu_{\mathrm{HII}},$
@@
 2.2.1 Unramified characters
-Let $X(\mathbf{G})=\mathrm{Hom}(\mathbf{G},{\mathbb{G}_{\mathrm{m}}})$ denote the rational characters of $\mathbf{G}$ defined over $F$.
+Let $X^{*}(\mathbf{G})=\mathrm{Hom}(\mathbf{G},{\mathbb{G}_{\mathrm{m}}})$ denote the rational characters of $\mathbf{G}$ defined over $F$.
 Let $\mathfrak{a}_{G}:=(X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{R})^{*}=(X^{*}(\mathbf{G})\otimes_{\mathbb{Z}}\mathbb{R})^{*}$ be the real Lie algebra of the maximal split central torus $\mathbf{A}=\mathbf{A}_{G}$ of $\mathbf{G}$ and let $\mathfrak{a}_{G}$$_{\mathbb{C}}$ be its complexification.
 We have a map
-$X(\mathbf{G})\to\mathrm{Hom}(G/G^{1},\mathbb{C}^{\times})=:\mathcal{X}(G)$
-given by $\chi\mapsto|\chi|_{F}$, where $|\chi|_{F}(g)=|\chi(g)|_{F}$ and $G^{1}=\bigcap_{\chi\in X(\mathbf{G})}\ker|\chi|_{F}$.
+$X^{*}(\mathbf{G})\to\mathrm{Hom}(G/G^{1},\mathbb{C}^{\times})=:\mathcal{X}(G)$
+given by $\chi\mapsto|\chi|_{F}$, where $|\chi|_{F}(g)=|\chi(g)|_{F}$ and $G^{1}=\bigcap_{\chi\in X^{*}(\mathbf{G})}\ker|\chi|_{F}$.
 This gives the unramified characters $\mathcal{X}(G)$ a complex manifold structure under which $\mathcal{X}(G)\simeq(\mathbb{C}^{\times})^{\dim_{\mathbb{R}}\mathfrak{a}_{G}}$.
 For indeed, we have
-${{\mathfrak{a}_{G}}_{\mathbb{C}}^{*}}$${\mathcal{X}(G)}$11{1}1
+$\mathfrak{a}_{G}$$_{\mathbb{C}}^{*}$$\mathcal{X}(G)$}1{1}
 given by
@@
 Denote $\mathrm{Im}\,\mathcal{X}(G)$ the set of unitary unramified characters taking values in the unit circle $\mathbb{T}\subset\mathbb{C}^{\times}$.
-This notation is justified, as if $|\chi|_{F}=|\chi^{\prime}|_{F}$, then $\mathrm{Re}(\chi)=\mathrm{Re}(\chi^{\prime})\in\mathfrak{a}_{G}$.
-Hence we can define $\mathrm{Im}\,\mathcal{X}(G)$ to be the unramified characters coming from pure imaginary elements of ${\mathfrak{a}_{G}}_{\mathbb{C}}$.
+This notation is justified, as if $|\chi|_{F}=|\chi^{\prime}|_{F}$, then $\mathrm{Re}(\chi)=\mathrm{Re}(\chi^{\prime})\in\mathfrak{a}_{G}^{*}$.
+Hence we can define $\mathrm{Im}\,\mathcal{X}(G)$ to be the unramified characters coming from pure imaginary elements of $\mathfrak{a}_{G}$$_{\mathbb{C}}^{*}$.
 The surjection $\mathrm{Im}\,\mathcal{X}(G)\to\mathrm{Im}\,\mathcal{X}(A)$ has finite kernel, and $\mathrm{Im}\,\mathcal{X}(A)$ is compact. We chose the Haar measure on it with volume one.
@@
 The group $\mathcal{X}(G)$ acts on admissible representations of $G$ by twisting: $\omega\mapsto\omega\otimes\nu$ for $\nu\in\mathcal{X}(G)$.
-This restricts to an action the unitary unramified characters on $\mathcal{E}_{2}(G)$.
-Pulling this action back to $i\mathfrak{a}_{G}$, and given a representation $\omega$, let $L^{*}$ be its stabilizer in $i\mathfrak{a}_{G}$, so that the orbit $\mathfrak{o}$ of $\omega$ is identified as $i{\mathfrak{a}_{G}}/L^{*}$.
-This gives $\mathfrak{o}=\mathrm{Im}\,\mathcal{X}(G)\cdot\omega$ the structure of a real submanifold of the larger orbit $\mathfrak{o}_{\mathbb{C}}={\mathfrak{a}_{G}}_{\mathbb{C}}/L^{*}=\mathcal{X}(G)\cdot\omega$.
+This restricts to an action of $\mathrm{Im}\,\mathcal{X}(G)$ on $\mathcal{E}_{2}(G)$.
+Pulling this action back to $i\mathfrak{a}_{G}^{*}$, and given a representation $\omega$, let $L^{*}$ be its stabilizer in $i\mathfrak{a}_{G}^{*}$, so that the orbit $\mathfrak{o}$ of $\omega$ is identified as $i{\mathfrak{a}_{G}}^{*}/L^{*}$.
+This gives $\mathfrak{o}=\mathrm{Im}\,\mathcal{X}(G)\cdot\omega$ the structure of a real submanifold of the larger orbit $\mathfrak{o}_{\mathbb{C}}={\mathfrak{a}_{G}}^{*}_{\mathbb{C}}/L^{*}=\mathcal{X}(G)\cdot\omega$.
 Definition 10.
@@
 If $G=\mathrm{SL}_{2}(F)$ and $M=A$ is the diagonal torus, we have $\mathfrak{a}_{G}^{*}\simeq\mathbb{R}$ and $R=\mathbb{Z}$ so that a fundamental domain for $\mathfrak{a}_{G}^{*}/\frac{2\pi}{\log q}R$ is $\left[-\frac{\pi}{\log q},\frac{\pi}{\log q}\right)$ and the canonical measure $d\nu=\frac{\log q}{2\pi}\mathrm{d}x$, where $\mathrm{d}x$ is the Lebesgue measure.
-To obtain quasicharacters of $G$, we associate to $\nu\in{\mathfrak{a}_{G}}_{\mathbb{C}}$ the quasicharacter $\chi_{\nu}(g)=q^{i\left\langle\nu,H_{G}(g)\right\rangle}$.
+To obtain quasicharacters of $G$, we associate to $\nu\in{\mathfrak{a}_{G}}^{*}_{\mathbb{C}}$ the quasicharacter $\chi_{\nu}(g)=q^{\left\langle\nu,H_{G}(g)\right\rangle}=|\nu(g)|_{F}$, where the second equality defines $H_{G}\colon G\to\mathfrak{a}_{G}$.
 Therefore to compute the integral of a function $f$ on $\mathcal{E}_{2}(G)$ supported on the unramified unitary characters of $A$, we compute
@@
 In general, $\mathcal{E}_{2}(M_{P})$ is a disjoint union of compact tori, and the Plancherel density descends to the quotients of these compact tori by certain finite groups, namely, the Weyl groups of $(P,A_{P})$.
-The set $\mathcal{E}_{2}(M_{P})^{I}$ is finite up to twist by unramified characters, by a result of Harish-Chandra [Waldspurger].
+The set $\mathcal{E}_{2}(M_{P})^{I}$ is finite up to twists by unramified characters, by a result of Harish-Chandra [Waldspurger].
 2.3 The Harish-Chandra Schwartz algebra
@@
 We have the obvious inclusion $\iota\colon H\mathrel{\hookrightarrow}\mathcal{C}^{I\times I}$.
-Define the subring $\mathcal{E}(G)$ of the endomorphism ring of the forgetful functor $\mathcal{M}(G)\to\mathbf{Vect}_{\mathbb{C}}$ from the category of all smooth representations, by replacing, in (1) above, $\mathcal{E}_{2}(M)$ with $\mathcal{M}(M)$, unitary characters with all unramified characters, and “smooth” with ”algebraic.” The matrix Paley-Wiener theorem of Bernstein [Bernstein] says that $f\mapsto\pi(f)$ is an isomorphism from the full Hecke algebra of $G$ onto $\mathcal{E}(G)$.
+Define the subring $\mathcal{E}(G)$ of the endomorphism ring of the forgetful functor $\mathcal{M}(G)\to\mathbf{Vect}_{\mathbb{C}}$ from the category of all smooth representations, by replacing, in (1) above, unitary characters with all unramified characters, and “smooth” with ”algebraic”, and adding
+  1. 3.
+    The endomorphisms $\eta_{\pi}$ are compatible with supercuspidal support, i.e. if $(M_{1},\sigma_{1})$ is the supercuspidal support of $\pi$, so that
+    $\pi=i_{P}^{G}(\omega\otimes\nu)\mathrel{\hookrightarrow}\pi_{1}=i_{P_{1}}^{G}(\sigma_{1}\otimes\nu_{1}),$
+    then $\eta_{\pi_{1}}$ preserves $\pi$ and $\eta_{\pi_{1}}|_{\pi}=\eta_{\pi}$.
+The matrix Paley-Wiener theorem of Bernstein [Bernstein] says that $f\mapsto\pi(f)$ is an isomorphism from the full Hecke algebra of $G$ onto $\mathcal{E}(G)$.
 Denote $\mathcal{E}^{I}$ and $\mathcal{E}_{t}$ the subrings of $I$-invariant endomorphisms.
+Remark 6.
+Property 3 means that unlike Schwartz functions, the Fourier transform a compactly supported function may be freely specified only on inductions of supercuspidal representations.
+In particular, an algebraic family of endomorphisms defined for all unramified characters may fail to come from a compactly-supported function.
+An example of a non-compactly supported function with regular Fourier transform is the element of $J$ given in Corollary 4, because $\pi(t_{1})=0$ except if $\pi$ is the Steinberg representation of $G$.
 For computational purposes such as ours, we require that these isomorphisms be explicit.
@@
 A character $\chi\colon M_{P}\to\mathbb{C}^{\times}$ of $M_{P}$ is non-strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x))|_{\infty}\geq 1$ for $|x|_{F}\geq 1$.
-We say a non-strictly positive character $\chi$ is strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x)|_{\infty}>1$ for $|x|_{F}\geq 1$.
-Of course, it suffices to test this for $x=\varpi^{-1}$. The condition for $\chi^{-1}$ to be non-strictly positive is then that $|\chi(\alpha^{\vee}(\varpi))|_{\infty}\geq 1$.
+We say a non-strictly positive character $\chi$ is strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x))|_{\infty}>1$ for $|x|_{F}>1$.
+Of course, it suffices to test this for $x=\varpi^{-1}$.
 Example 3.
-For $\mathbf{G}=\mathrm{SL}_{2}$, in the conventions fixed in Example 2, an unramified character $\chi$ of $A$ is non-strictly positive if $\chi^{-1}$ corresponds to $z$ such that $|z|\geq 1$.
-If $\mathbf{G}=\mathrm{GL}_{n}$ and $\nu$ corresponds to the vector $(z_{1},\dots,z_{n})\in\mathbb{C}^{n}$, then the condition that $\nu^{-1}$ is non-strictly positive translate to $|z_{n}|\geq|z_{n-1}|\geq\cdots\geq|z_{1}|$.
+For $\mathbf{G}=\mathrm{SL}_{2}$, in the conventions fixed in Example 2, an unramified character $\chi$ of $A$ is non-strictly positive if it corresponds to $z$ such that $|z|\geq 1$.
+If $\mathbf{G}=\mathrm{GL}_{n}$ and $\nu$ corresponds to the vector $(z_{1},\dots,z_{n})\in\mathbb{C}^{n}$, then the condition that $\nu$ is non-strictly positive translates to $|z_{1}|\geq|z_{2}|\geq\cdots\geq|z_{n}|$.
 Such conditions divide $\mathbb{C}^{n}$ into chambers, on which the Weyl group $\mathfrak{S}_{n}$ clearly acts simply-transitively.
-Interior points correspond to $\nu$ such that $\nu^{-1}$ is strictly positive.
+Interior points correspond to strictly positive $\nu$.
 Following op. cit., let $\mathcal{E}^{I}_{J}(G)$ denote the subring of $\mathcal{E}_{t}(G)$ defined by the following conditions on the endomorphisms $\eta_{\pi}$:
   1. 1.
-    For all $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$, the endomorphism $\eta_{\pi}=\eta_{\nu,\omega}$ is a rational function of $\nu$, regular on the set of $\nu$ such that $\nu^{-1}$ is non-strictly positive.
+    For all $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$, the endomorphism $\eta_{\pi}=\eta_{\nu,\omega}$ is a rational function of $\nu$, regular on the set of non-strictly positive $\nu$.
   2. 2.
@@
   1. 1.
-    Let $\pi$ be a tempered representation of $G$. Then the action of $H$ on $\pi^{I}$ extends uniquely to $J$.
+    Let $\pi$ be a tempered representation of $G$. Then the action of $H$ on $\pi^{I}$ extends uniquely to $J$ via $\phi\circ{}^{\dagger}(-)$.
   2. 2.
     Let $P=MN$ be a parabolic subgroup of $G$ with Levi subgroup $M$ and let $\omega$ an irreducible tempered representation of $M$.
-Let $\nu$ be a non-strictly positive character of $M$ and let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu^{-1})$.
+Let $\nu$ be a non-strictly positive character of $M$ and let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$.
 Then the action of $H(G,I)$ on $\pi^{I}$ extends uniquely to an action of $J$.
   3. 3.
-    The map $t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}$ defines a homomorphism
+    The action of $J$ on the representations $\pi^{I}$ in (2) extends rationally in $\nu$ to define a homomorphism
     $\eta\colon J\to\mathcal{E}_{J}^{I}(G).$
-Composing this isomorphism with the inverse Fourier transform, Braverman and Kazhdan define an algebra map
+    denoted
+    $t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}.$
+The proof of Theorem 7 (1) and (2) in [BK] uses Theorem 4 and [XiJAMS], to show that, in the notation of Section 1.5.1,
+$E(u,s,\rho)|_{H}\simeq{}^{*}K(u,s,\rho,q)={}^{*}i_{P}^{G}(\sigma\otimes\nu)$
+where the restriction is via $\phi$, and hence that
+$E(u,s,\rho)_{H,\phi\circ{}^{*}(-)}\simeq K(u,s,\rho,q)=i_{P}^{G}(\sigma\otimes\nu),$
+where the restriction is now via $\phi\circ{}^{*}(-)$.
+By Lemma 1, the map $\phi$ is $\bar{\kappa}$-linear, where $\bar{\kappa}$ is as in Definition 6. Of course, $\phi$ also intertwines conjugation by $T_{w_{0}}$ and $\phi(T_{w_{0}})$.
+As noted in Section 1.5.1, if $\pi=K(u,s,\rho)$ is simple tempered (or more generally, is any semisimple module), then by [PrasadDuke, Prop. 6.3],
+$E(u,s,\rho)|_{\phi\circ(-)^{\dagger}}={}^{\dagger}{}^{*}K(u,s,\rho)\simeq\widetilde{K(u,s,\rho)}$ (4)
+is simple tempered. Moreover, by Lemma 2 (e) and (4), we have, if $\Psi_{w_{0}}(h)=T_{w_{0}}hT_{w_{0}}^{-1}$, that
+${}^{\bar{\kappa}\circ\phi(\Psi_{w_{0}}^{-1})}E(u,s,\rho)|_{\phi\circ{}^{\dagger}(-)}$ $={}^{\phi(\Psi_{w_{0}}^{-1})}E(u,s,\rho)|_{\bar{\kappa}\circ\phi\circ{}^{\dagger}(-)}$
+$=E(u,s,\rho)|_{\phi(\Psi_{w_{0}}^{-1})\circ\bar{\kappa}\circ\phi\circ{}^{\dagger}(-)}$
+$=E(u,s,\rho)|_{\phi\circ\Psi_{w_{0}}^{-1}\circ\bar{\kappa}\circ{}^{\dagger}(-)}$
+$=E(u,s,\rho)|_{\phi\circ{}^{*}(-)}$
+$=K(u,s,\rho)$
+for any standard $H$-module $K(u,s,\rho)$. Therefore any standard $H$-module extends to a simple $J$-module via $\phi\circ{}^{\dagger}(-)$.
+Composing the morphism $\eta$ with the inverse Fourier transform, Braverman and Kazhdan define an algebra map
 $\tilde{\phi}\colon J\to\mathcal{C}^{I\times I}$
 sending
-$t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}\mapsto f_{w}=\sum_{x\in\tilde{W}}A_{x,w}T_{x}\in\mathcal{C}^{I\times I},$
+$t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}\mapsto f_{w}=\sum_{x\in W_{\mathrm{aff}}}A_{x,w}T_{x}\in\mathcal{C}^{I\times I},$
 where $A_{x,w}=f_{w}(IxI)$. By definition, $\eta_{\pi}(w)=\pi(f_{w})$ as endomorphisms of $\pi$. We will show later that $\tilde{\phi}$ is essentially the map $\phi^{-1}$.
@@
 Proof.
-Let $\pi$ be a tempered representation of $G$.
-By [BK], the $H$-action on it extends to a $J$-action, and by Corollary 2.6 of loc.
-cit., we have $\eta_{\pi}(\phi_{q}(f^{\dagger}))=\pi(f)$ in $\mathcal{E}_{t}^{I}$ for any $f\in H$.
-Therefore $\tilde{\phi}\circ\phi_{q}(f^{\dagger})=f$ by the Plancherel theorem.
-∎
-Remark 6.
-The presence of the Goldman involution is necessary for tempered representations of $J$ to restrict to tempered representations of $H$.
-For example, in [BK], Braverman and Kazhdan showed for $G=\mathrm{SL}_{2}(F)$, that $\mathcal{E}^{I}_{J}(G)=\operatorname{End}(\mathrm{St}^{I})\oplus\operatorname{End}_{\Phi,X_{*}}(\mathcal{S}^{I})$, matching the decomposition $J=\mathbb{C}t_{1}\oplus J_{0}$, where $\Phi$ is the Fourier transform on the Schwartz space of the basic affine space $\mathcal{S}$ [BKBasicAffine], and $\mathrm{St}$ is the Steinberg representation.
-In this case $J$ has three one-dimensional representations, whereas $H$ has one one-dimensional discrete series representation $\mathrm{St}$, two tempered one-dimensional representations not belonging to the discrete series, and one non-tempered one-dimensional representation, the trivial representation.
-It is easy to compute using the formulas of [D] that $\phi_{q}(T_{s})t_{1}=qt_{1}$, whereas $\phi_{q}(T_{s}^{\dagger})t_{1}=-t_{1}$.
-Thus for the Steinberg representation to restrict to the Steinberg representation, the twist is required.
+Let $\pi$ be a tempered representation of $G$. By Theorem 7 and (4), the $H$-action on it extends to a $J$-action via $\phi\circ^{\dagger}(-)$, such that
+$\eta_{\pi}(\phi_{q}({}^{\dagger}f))=\pi(f)$
+in $\mathcal{E}_{t}^{I}$ for any $f\in H$. Therefore $\tilde{\phi}\circ\phi_{q}({}^{\dagger}f)=f$ by the Plancherel theorem. ∎
 2.5 The Plancherel formula for $\mathrm{GL}_{n}$
@@
 For any representation $\sigma$ of $M$, if $f\in\operatorname{Ind}_{P}^{G}(\sigma)$ is $I$-fixed, then for $i_{M}\in I_{M}$, we have
-$f(i_{M})=\sigma(i_{M})\delta^{\frac{1}{2}}_{P}(i_{M})f(1)=f(1),$
-and so if $\delta^{\frac{1}{2}}_{P}(i_{M})=1$, then it must be that $f(1)\in\sigma^{I_{M}}$. The Iwahori factorization
-$I_{M}=N_{P}^{-}(\pi\mathcal{O})A(\mathcal{O})N_{M}(\mathcal{O})$
-and the fact that $\mathrm{Ad}_{\mathfrak{n}_{P}}$ is a representation implies that
-$\delta^{\frac{1}{2}}_{P}(i_{M})=|\det\mathrm{Ad}_{\mathfrak{n}_{P}}(N_{P}^{-}(i_{M})A(i_{M})N_{M}(i_{M}))|_{F}=|\det\mathrm{Ad}_{\mathfrak{n}_{P}}(A(i_{M}))|_{F}=1$
-by unipotence of the groups $N_{P}^{-}$ and $N_{M}$ and the fact that the image of $i_{M}$ in $M(\mathbb{F}_{q})$ lies in the Borel. ∎
+$f(i_{M})=\sigma(i_{M})\delta^{\frac{1}{2}}_{P}(i_{M})f(1)=f(1).$
+As $\delta_{P}=1$ on every compact subgroup of $P$, we have $\delta^{\frac{1}{2}}_{P}(i_{M})=1$, and $f(1)\in\sigma^{I_{M}}$. ∎
 Let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$ be a tempered representation and let $(u,s)$ be the KL-parameter of its discrete support.
@@
 Then $\pi^{I}$ is a tempered irreducible $H$-module, and is of the form $K(u,s,\rho,q)$ for $u,s\in\mathbf{G}^{\vee}(\mathbb{C})$ and $\rho$ a representation of $\pi_{0}(Z_{G^{\vee}}(u,s))$ with $s$ compact.
-By [BK], Corollary 2.6, $K(u,s,\rho,q)$ extends to a $J$-module.
-By definition of Lusztig’s bijection, $\pi(f_{w})\neq 0$ only if $w$ is in the two-sided cell $\mathbf{c}=\mathbf{c}(u)$ of $\tilde{W}$ corresponding to $u$.
+By Theorem 7, $K(u,s,\rho,q)$ extends to a $J$-module.
+By definition of Lusztig’s bijection, $\pi(f_{w})\neq 0$ only if $w$ is in the two-sided cell $\mathbf{c}=\mathbf{c}(u)$ of $W_{\mathrm{aff}}$ corresponding to $u$.
 On the other hand, $K(u,s,\rho,q)$ is induced from a square-integrable standard module $K_{M}(u,s,\tilde{\rho},q)$ of $H(M,I_{M})$.
-But now [KLDeligneLanglands, Theorem 8.3] says that $(u,s)$ must be preicsely as in the statement of the proposition.
+But now [KLDeligneLanglands, Theorem 8.3] says that $(u,s)$ must be precisely as in the statement of the proposition.
 Thus only such summands $(f_{w})_{M}$ in the Plancherel decomposition of $f_{w}$ are nonzero.
@@
 Note that all the positive roots $\Sigma^{+}(\theta)$ of $(P_{\theta},N_{\theta})$ are in $N_{P}$.
-Denoting by $[\alpha]$ the coset representing a root $\alpha$ of $G$ restricted to $P$, the positive roots in $N_{P}$ are the $\alpha_{ij}$ such that $[\alpha_{ij}]=[\beta_{n_{1}+\cdots+n_{k}}]$.
+Denoting by $\alpha]$ the coset representing a root $\alpha$ of $G$ restricted to $P$, the positive roots in $N_{P}$ are the $\alpha_{ij}$ such that $\alpha_{ij}]=[\beta_{n_{1}+\cdots+n_{k}}]$.
 Example 4.
@@
 $\mathrm{trace}\left({\pi}\,,{f}\right)\in\mathbb{C}[z_{1},\dots,z_{k},z_{1}^{-1},\dots,z_{k}^{-1}].$
-That is, the trace is a regular function on (×$(\mathbb{C}^{\times})^{k}$.
+That is, the trace is a regular function on $(\mathbb{C}^{\times})^{k}$.
 Proof.
@@
 When $\nu$ is non-strictly positive with respect to $M$, by Theorem 7, $\mathrm{trace}\left({\pi}\,,{f}\right)$ has poles only of the form $z_{i}^{n_{i}}=0$.
-The claim now follows from the $W$-invariance, and thus $\mathrm{trace}\left({\pi}\,,{f}\right)$ is a regular function on (×$(\mathbb{C}^{\times})^{k}$.
+The claim now follows from the $W$-invariance, and thus $\mathrm{trace}\left({\pi}\,,{f}\right)$ is a regular function on $(\mathbb{C}^{\times})^{k}$.
 ∎
@@
 Lemma 6.
-Let $d\in\tilde{W}$ be a distinguished involution and $\pi$ be a tempered representation induced from one of the Levi subgroups attached to the two-sided cell containing $d$ in the sense of Proposition 1.
+Let $d\in W_{\mathrm{aff}}$ be a distinguished involution and $\pi$ be a tempered representation induced from one of the Levi subgroups attached to the two-sided cell containing $d$ in the sense of Proposition 1.
 Then $\mathrm{trace}\left({\pi}\,,{f_{d}}\right)$ is constant and a natural number.
@@
 ∎
-3 Proof of Theorem 13 for $\mathrm{GL}_{n}$, and of Theorem 2 for general $\mathbf{G}$
+3 Proof of Theorem 2 for general $\mathbf{G}$, and the case of $\mathrm{GL}_{n}$
 3.1 The functions $f_{w}$ for $\mathrm{GL}_{n}$
 In this section, $q>1$.
-To compute with the Plancherel formula, we will need to apply the residue theorem successively in each variables $z_{i}$, and in doing so will we will need to sum over a certain tree that will track, for each variable, at which residues we evaluated.
+To compute with the Plancherel formula, we will need to apply the residue theorem successively in each variable $z_{i}$, and in doing so will we will need to sum over a certain tree that will track, for each variable, at which residues we evaluated.
 Upon integrating with respect to each variable $z_{i}$, we will have poles of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$.
@@
 {forest}
-[$z_{1}=0$ [$z_{2}=0$ [$z_{3}=0$] [$z_{3}=(q^{34})^{-1}z_{4}$] ] [$z_{2}=(q^{23})^{-1}z_{3}$ [$z_{3}=0$] [$z_{3}=(q^{34})^{-1}z_{4}$] ] [$z_{2}=(q^{24})^{-1}z_{4}$ [$z_{4}=0$] [$z_{4}=(q^{43})^{-1}z_{3}$] ] ]
+$z_{1}=0$ $z_{2}=0$ $z_{3}=0$] $z_{3}=(q^{34})^{-1}z_{4}$] ] $z_{2}=(q^{23})^{-1}z_{3}$ $z_{3}=0$] $z_{3}=(q^{34})^{-1}z_{4}$] ] $z_{2}=(q^{24})^{-1}z_{4}$ $z_{4}=0$] $z_{4}=(q^{43})^{-1}z_{3}$] ] ]
 Of course, to evaluate the entire integral for $M$, we would also need to consider trees whose roots are decorated with $z_{1}=(q^{12})^{-1}z_{2}$, and so on, for a total of four trees.
 Definition 13.
-Given a Levi subgroup $M$ with $N+1$ blocks, a bookkeeping tree $T$ for $M$ is a rooted tree with $N$ levels such that the vertices on the $i$-th level below the root each have $N-i$ child vertices, and each vertex is decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$, where the index $j$ does not appear along the path from the vertex to the root, and the parent root is decorated with an equation $z_{k}=(q^{ki})^{-1}z_{i}$ for some $k$.
+Given a Levi subgroup $M$ with $N+1$ blocks, a bookkeeping tree $T$ for $M$ is a rooted tree with $N$ levels such that the vertices on the $i$-th level below the root each have NiN-i child vertices, and each vertex is decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$, where the index $j$ does not appear along the path from the vertex to the root, and the parent root is decorated with an equation $z_{k}=(q^{ki})^{-1}z_{i}$ for some $k$.
 Moreover, we require that the root be decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$ for $i$ minimal.
@@
 Note that as $l_{i_{i}}+l_{i_{k+1}}\leq n$, the leftmost denominators in (12) satisfy the conclusion of the corollary also.
-Finally, $Q_{rk}q^{i_{r}i_{k+1}}=q^{l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}}$, and $R_{rk}$ is likewise always a polynomial in $q$ (as opposed to $q^{-1}$) divisible by $1-q$.
+Finally, $Q_{rk}q^{i_{r}i_{k+1}}=q^{l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}}$, and $R_{rk}$ is likewise always a polynomial in $q$ (as opposed to $q^{-1}$) divisible by 1q1-q.
 Again using that $l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}\leq n$, we are done with the first statement.
@@
 $d\pi_{\mu_{I}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}d\pi_{\mu_{K}}$ (13)
-for tempered principal series. Now let $w\in\mathbf{c}_{0}$. By Lemma 4 and the Satake isomorphism, there is a function $h_{w}$ in the spherical Hecke algebra such that for all principal series $\pi$,
+for tempered principal series.
+Now let $w\in\mathbf{c}_{0}$ and consider $f_{w}=\tilde{\phi}(t_{w})$.
+By Lemma 5 and the Satake isomorphism, there is a function $h_{w}$ in the spherical Hecke algebra such that for all principal series $\pi$,
 $\mathrm{trace}\left(\pi_{I}(f_{w})\right)=\mathrm{trace}\left(\pi_{K}(h_{w})\right)$
 as regular Weyl-invariant functions of the Satake parameter. By the Plancherel formula and (13),
-$f_{w}(1)=\int{\mathrm{trace}\left(\pi_{I}(f_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{K}}}=\frac{h_{w}(1)}{P_{\mathbf{G}/\mathbf{B}}(q)}.$ (12)
+$f_{w}(1)=\int{\mathrm{trace}\left(\pi_{I}(f_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{K}}}=\frac{h_{w}(1)}{P_{\mathbf{G}/\mathbf{B}}(q)},$ (14)
 In particular, if $w=d$ is a distinguished involution in the lowest two-sided cell, then
@@
 cit.
-there is the assumption of simple-connectedness, but it is easy to see that the distinguished involutions for the extended affine Weyl group $\tilde{W}(\tilde{\mathbf{G}}(F))$ of the universal cover $\tilde{\mathbf{G}}$ are distinguished involutions for $\tilde{W}$ using the definition in [affineII] and uniqueness of the $\{C_{w}\}$-basis, and that the lowest cell is just the lowest cell of $\tilde{W}(\tilde{\mathbf{G}}(F))$ intersected with $\tilde{W}$.) ∎
+there is the assumption of simple-connectedness, but it is easy to see that the distinguished involutions for the extended affine Weyl group $\tilde{W}(\tilde{\mathbf{G}}(F))$ of the universal cover $\tilde{\mathbf{G}}$ are distinguished involutions for $W_{\mathrm{aff}}$ using the definition in [affineII] and uniqueness of the $\{C_{w}\}$-basis, and that the lowest cell is just the lowest cell of $\tilde{W}(\tilde{\mathbf{G}}(F))$ intersected with $W_{\mathrm{aff}}$.) ∎
 Once we have established injectivity of $\tilde{\phi}$, we will show by a counting argument that $\operatorname{rank}\left(\pi(t_{d})\right)=1$ for any distinguished involution $d$, in the case $G=\mathrm{GL}_{n}$, see Theorem 13.
@@
 Corollary 3.
-Let $w\in\tilde{W}$.
+Let $w\in W_{\mathrm{aff}}$.
 Then $f_{w}(1)$ is a rational function of $q$ with denominator dividing a power of the Poincaré polynomial of $\mathbf{G}$.
-The numerator is a Laurent polynomial $p_{1}(q^{1/2})+p_{2}(q^{-1/2})$ in $q^{1/2}$, where the degree of $p_{1}$ is bounded uniformly in terms of $\tilde{W}$.
+The numerator is a Laurent polynomial $p_{1}(q^{1/2})+p_{2}(q^{-1/2})$ in $q^{1/2}$, where the degree of $p_{1}$ is bounded uniformly in terms of $W_{\mathrm{aff}}$.
 The denominator of $f_{w}(1)$ depends only on the two-sided cell containing $w$.
@@
 where we view $\nu$ as a semisimple conjugacy class in $M_{\mathbf{c}}$, and $V(\lambda)$ is the irreducible representation of $M_{\mathbf{c}}$ of highest weight $\lambda$.
-Then we have that $f_{\lambda}(1)\neq 0$ only if $\lambda$ is of height 00 with respect to the basis $\varepsilon_{i}\colon\mathrm{diag}(a_{1},\dots,a_{n})\mapsto a_{i}$.
+Then we have that $f_{\lambda}(1)\neq 0$ only if $\lambda$ is of height 0 with respect to the basis $\varepsilon_{i}\colon\mathrm{diag}(a_{1},\dots,a_{n})\mapsto a_{i}$.
 The proofs of Lemma 7 and Corollary 3 will use the notation of the proof of Theorem 11, and we defer them until after the proof of the theorem.
@@
 which is one of the factors in the product $(1-Q_{r,k-1}q_{i_{r}i_{k}})(Q_{r,k-1}(q_{i_{r}i_{k}})^{-1})$.
-The surviving factor in the numerator at index $(k-1,r)$ is then equal to $(1-Q_{r,k-1}q^{g_{i_{r}}-g_{i_{k}}})$.
+The surviving factor in the numerator at index (k1,r)(k-1,r) is then equal to $(1-Q_{r,k-1}q^{g_{i_{r}}-g_{i_{k}}})$.
 In short, the above factors in the denominator cancel with a numerator occurring with the same $r$-index but $k$-index one lower.
-Such a factor occurs whenever $r<k-1$ (note that this inequality does not hold when $k=1$ and $r=0$).
-When $r=k-1$, we have
+Such a factor occurs whenever r<k1r<k-1 (note that this inequality does not hold when $k=1$ and $r=0$).
+When r=k1r=k-1, we have
 $1-Q_{k-1,k}(q^{i_{k}i_{i+1}})^{-1}=1-q^{i_{k}i_{k+1}}q^{i_{k-1}i_{k}}(q^{i_{k-1}i_{k+1}})^{-1}=q^{g_{i_{k}}+g_{i_{k+1}}+1}q^{g_{i_{k-1}}-g_{i_{k+1}}}=1-q^{2g_{i_{k}}+1},$
@@
 First, observe that evaluating a residue of the form $z_{i}=(q^{ij})^{-1}z_{j}$ increases the sum of all powers by $1$; a factor $z_{j}$ is contributed to the resulting integrand.
-The sum of all powers is likewise increased by $1$ when evaluating the residue at a simple pole of $z_{i}^{-1}$ at 00.
-To compute the residue at a pole of $z_{i}^{-e_{i}}$ at 00 for $e_{i}>1$, consider the Taylor expansion at 00 of $\Gamma^{ij}$.
+The sum of all powers is likewise increased by $1$ when evaluating the residue at a simple pole of $z_{i}^{-1}$ at 0.
+To compute the residue at a pole of $z_{i}^{-e_{i}}$ at 0 for $e_{i}>1$, consider the Taylor expansion at 0 of $\Gamma^{ij}$.
 We have
@@
 Multiplying these series and further multiplying by the denominator $z_{i}^{2}-(q_{ij}+q_{ij}^{-1})z_{i}z_{j}+z_{j}$, it follows that the Taylor expansion of $\Gamma^{ij}$ is
-$1+\frac{q^{ij}+(q^{ij})^{-1}-q_{ij}-q_{ij}^{-1}}{z_{j}}z_{i}+\frac{(q^{ij})^{2}+2+(q^{ij})^{-2}-(q_{ij}+q_{ij}^{-1})(q^{ij}+q^{-ij})}{z_{j}}z_{i}^{2}+\cdots\\ +\frac{-(q_{ij}+q_{ij}^{-1})((q^{ij})^{n-1}+\cdots+(q^{ij})^{-n+1})+(q^{ij})^{n-2}+\cdots+(q^{ij})^{-n+2}+(q^{ij})^{n}+\cdots+(q^{ij})^{-n}}{z_{j}^{n}}z_{i}^{n}\\ +\cdots.$ (17)
+$1+\frac{q^{ij}+(q^{ij})^{-1}-q_{ij}-q_{ij}^{-1}}{z_{j}}z_{i}+\frac{(q^{ij})^{2}+2+(q^{ij})^{-2}-(q_{ij}+q_{ij}^{-1})(q^{ij}+q^{-ij})}{z_{j}^{2}}z_{i}^{2}+\cdots\\ +\frac{-(q_{ij}+q_{ij}^{-1})((q^{ij})^{n-1}+\cdots+(q^{ij})^{-n+1})+(q^{ij})^{n-2}+\cdots+(q^{ij})^{-n+2}+(q^{ij})^{n}+\cdots+(q^{ij})^{-n}}{z_{j}^{n}}z_{i}^{n}\\ +\cdots.$ (19)
 It is clear that the salient point of (19), that $z_{i}^{n}$ appears with a coefficient proportional to $z_{j}^{-n}$, holds also for the all the products like those in (18).
-Thus computing a residue at the pole of $z_{i}^{-e_{i}}$ at 00 will result in a new integrand, the total degree of which has increased by $e_{i}-e_{i}+1=1$.
-It now follows that after integrating with respect to $N-1$ variables, the final integral to be computed will be a constant times $(2\pi i)^{-1}\int_{\mathbb{T}}z_{N}^{e_{0}+\cdots+e_{N}+N-1}\mathrm{d}z_{N}$.
+Thus computing a residue at the pole of $z_{i}^{-e_{i}}$ at 0 will result in a new integrand, the total degree of which has increased by $e_{i}-e_{i}+1=1$.
+It now follows that after integrating with respect to N1N-1 variables, the final integral to be computed will be a constant times $(2\pi i)^{-1}\int_{\mathbb{T}}z_{N}^{e_{0}+\cdots+e_{N}+N-1}\mathrm{d}z_{N}$.
 This is nonzero if and only if $e_{0}+\cdots+e_{N}=-N$.
@@
 Therefore either $e_{i_{0}}+e_{i_{1}}+1=e_{i_{2}}=\cdots=e_{-N}=-1$, or we may again assume without loss of generality that the exponent of some $z_{i_{j}}$ is nonnegative, and proceed with evaluating (20) by integrating with respect to $z_{i_{j}}$.
-We may continue in this way, never having to deal with more than a simple pole at 00.
+We may continue in this way, never having to deal with more than a simple pole at 0.
 (Of course, the resulting order of integration need not be the same as the order used in the proofs of Theorem 11 and Corollary 3.)
@@
 This and the calculations in the proof of Theorem 11 make clear that the only positive powers of $q^{1/2}$ that appear in the any denominator are those that appeared as in Theorem 11.
-These are controlled by the possible block sizes of $M$, and hence are bounded in terms of $\tilde{W}$.
+These are controlled by the possible block sizes of $M$, and hence are bounded in terms of $W_{\mathrm{aff}}$.
 Throughout this procedure, the denominator has been contributed to only by the Plancherel density itself, which depends only on $M$.
@@
 Theorem 12.
-Let $w\in\tilde{W}$.
-Then $f_{w}(1)$ is a rational function of $q$ with poles drawn from a finite set of roots of unity depending only on $\tilde{W}$.
+Let $w\in W_{\mathrm{aff}}$.
+Then $f_{w}(1)$ is a rational function of $q$ with poles drawn from a finite set of roots of unity depending only on $W_{\mathrm{aff}}$.
 The numerator is a Laurent polynomial in $q^{1/2}$.
@@
 First, we note that the reasoning for the lowest cell used in the proof of Corollary 2 holds for arbitrary $G$. Therefore (14) proves the last claim just as for $\mathrm{GL}_{n}$.
-More generally, let $w\in\tilde{W}$ and $M_{P}$ be a Levi subgroup corresponding to the two-sided cell containing $w$.
+More generally, let $w\in W_{\mathrm{aff}}$ and $M_{P}$ be a Levi subgroup corresponding to the two-sided cell containing $w$.
 Let $N=\operatorname{rk}A_{P}$.
@@
 3.3 Relating $t_{w}$ and $f_{w}$
-[…]
+We will now relate the Schwartz functions $f_{w}$ on $G$ to the elements $\phi^{-1}(t_{w})$ of a completion $\mathbfcal{H}^{-}$ of $\mathbf{H}$, whose definition we will now recall.
+In this section $\mathbf{G}$ is general.
 3.3.1 Completions of $\mathbf{H}$ and $J\otimes_{\mathbb{C}}\mathcal{A}$
-[…]
+Let $\hat{\mathcal{A}}=\mathbb{C}((\mathbf{q}^{-1/2}))$ and $\hat{\mathcal{A}}^{-}=\mathbb{C}[[\mathbf{q}^{-1/2}]]$. Write $\mathbfcal{H}^{-}$ for the $\hat{\mathcal{A}}-$algebra
 […]
@@
 of $J\otimes_{\mathbb{C}}\mathcal{A}$.
-[…]
+In [affineII], Lusztig shows that $\phi$ extends to an isomorphism of $\hat{\mathcal{A}}-$algebras $\mathbfcal{H}^{-}_{C}\to\mathbfcal{J}$.
+In this way the elements $t_{w}\in J\subset\mathbfcal{J}$ may be identified with elements of $\mathbfcal{H}^{-}_{C}$ via $\phi$.
 Lemma 8.
-[…]
+We have $\mathbfcal{H}^{-}_{C^{\prime}}\subset\mathbfcal{H}^{-}$. The inclusion is continuous.
 Proof.
@@
 3.3.2 The functions $f_{w}$ and the basis elements $t_{w}$
-[…]
+We shall now explain how the map $\tilde{\phi}\colon J\to\mathcal{C}(G)^{I}$ induces a map of $\mathcal{A}$-algebras $\hat{\phi}\colon J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$.
 Proposition 2.
-[…]
+There is a map of $\mathcal{A}$-algebras $\hat{\phi}\colon J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$ such that if $\hat{\phi}(t_{w})=\sum_{x}a_{x,w}T_{x}$, then $a_{x,w}(q)=f_{w}(x)$.
+Moreover, there is a constant $N$ depending only on $W_{\mathrm{aff}}$ such that $a_{x,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $x,w\in W_{\mathrm{aff}}$.
 The most difficult part of the proof of the proposition is showing that $a_{x,w}\to 0$ as $\ell(x)\to\infty$. To this end we have
 Lemma 9.
-Let $w\in\tilde{W}$.
-Then the degree in $q$ of the numerator of $f_{w}(1)$ is bounded uniformly in $w$ by some $N$ depending only on $\tilde{W}$, and hence in the notation of Proposition 2 above, we have $a_{1,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $w\in\tilde{W}$.
+Let $w\in W_{\mathrm{aff}}$.
+Then the degree in $q$ of the numerator of $f_{w}(1)$ is bounded uniformly in $w$ by some $N$ depending only on $W_{\mathrm{aff}}$, and hence in the notation of Proposition 2 above, we have $a_{1,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $w\in W_{\mathrm{aff}}$.
 Remark 12.
@@
 where $h_{1},k_{1}\in\mathbb{C}[q]$ and $h_{2},k_{2}\in q^{-1}\mathbb{C}[q^{-1}]$, and the sum is over the finitely many, up to unitary twist, pairs $(M,\omega)$, for $\omega\in\mathcal{E}_{2}(M)$, such that $\mathrm{trace}\left({\pi_{\omega}}\,,{f_{w}}\right)\neq 0$.
-Fix $w\in\tilde{W}$. Then
+Fix $w\in W_{\mathrm{aff}}$. Then
 $|f_{w}(1)|=\left|\sum_{(P,\omega)}\int_{\mathcal{O}(\omega)}\mathrm{trace}\left({\pi_{\omega}}\,,{f_{w}}\right)d(\omega)\,\mathrm{d}\mu_{M_{P}}(z)\right|\leq\frac{h_{1}(q)+h_{2}(q^{-1})}{k_{1}(q)+k_{2}(q^{-1})}.$
@@
 for $q$ sufficiently large. Indeed, let $\epsilon>0$ be given and choose $q\gg 1$ such that
-$\left||(f_{1}(q)+f_{2}(q^{-1}))||(k_{1}(q)+k_{2}(q^{-1}))|-|f_{1}(q)k_{1}(q)|\right|\leq\epsilon$
+[…]
 and
-$\left||(d_{1}(q)+d_{2}(q^{-1}))||(h_{1}(q)+h_{2}(q^{-1}))|-|d_{1}(q)h_{1}(q)|\right|\leq\epsilon,$
+[…]
 and (24) holds. Then we have
@@
 Now, the denominator of $f_{w}(1)$, and hence $\deg d_{1}$, depends only on the two-sided cell containing $w$, again by Corollary 3 and Theorem 12.
-We can also bound $\deg h_{1}$ uniformly in terms of $\tilde{W}$, as it depends only on the Plancherel measure and the finitely many possible formal degrees appearing in the parametrization of the $I$-spherical part of the tempered dual of $G$.
+We can also bound $\deg h_{1}$ uniformly in terms of $W_{\mathrm{aff}}$, as it depends only on the Plancherel measure and the finitely many possible formal degrees appearing in the parametrization of the $I$-spherical part of the tempered dual of $G$.
 This proves the lemma.
@@
 Proof of Proposition 2.
-[…]
-By Corollary 2 and Theorem 12 , there is a formal power series in $\hat{\mathcal{A}}^{-}$ with constant term equal to $1$ that specializes to the denominator of $A_{1,w}$ when $\mathbf{q}=q$.
+Let $w\in W_{\mathrm{aff}}$.
+Write $\tilde{\phi}(t_{w})=f_{w}=\sum_{x}A_{x,w}T_{x}$ as Schwartz functions on $G$ so that $A_{x,w}=f_{w}(x)$.
+We must show that there is a unique element $a_{x,w}\in\hat{\mathcal{A}}$ such that $a_{x,w}(q)=A_{x,w}$ as complex numbers.
+We will then check that $a_{x,w}\to 0$ rapidly enough as $\ell(x)\to\infty$ for $\sum_{x}a_{x,w}T_{x}$ to define an element of $\mathbfcal{H}^{-}$.
+By Corollary 3 and Theorem 12 , there is a formal power series in $\hat{\mathcal{A}}^{-}$ with constant term equal to 11 that specializes to the denominator of $A_{1,w}$ when $\mathbf{q}=q$.
 Moreover, there is a unique formal Laurent series $a_{1,w}\in\hat{\mathcal{A}}$ such that $a_{1,w}(q)=A_{1,w}$ for all prime powers.
@@
 $q^{-\ell(x)}(f\star T_{x^{-1}})(1)=q^{-\ell(x)}\int_{G}{f(g)T_{x^{-1}}(g^{-1})}\,\mathrm{d}{\mu_{I}(g)}\\ =q^{-\ell(x)}\int_{IxI}{f(g)}\,\mathrm{d}{\mu_{I}(g)}=q^{-\ell(x)}\mu_{I}(IxI)f(x)=f(x).$
-By definition, $f_{w}(x)=A_{x,w}$. On the other hand, according to Lemma 2, we have
-$q^{-\ell(x)}(f_{w}\star T_{x^{-1}})(1)$ $=q^{-\ell(x)}\left(\tilde{\phi}(t_{w})\star\tilde{\phi}\left(\phi_{q}(T_{x^{-1}}^{\dagger})\right)\right)(1)$
-$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}(T_{x^{-1}}^{\dagger})\right)(1)$
-$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})-\ell(y)}Q_{y,x^{-1}}(q)(C^{\prime}_{y})^{\dagger}\right)\right)(1)$ (24)
-$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)(C_{y})\right)\right)(1)$ (25)
-$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\\ d\in\mathcal{D}\\ a(d)=a(r)\end{subarray}}h_{y,d,r}t_{r}\right)(1)$ (26)
-$=q^{-\ell(x)}\tilde{\phi}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ d\in\mathcal{D}\\ a(d)=a(r)=a(w)\end{subarray}}h_{y,d,r}t_{w}t_{r}\right)(1)$
-$=q^{-\ell(x)}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ a(r)=a(w)\end{subarray}}h_{y,d_{w},r}(f_{w}\star f_{r})(1),$ (27)
+By definition, $f_{w}(x)=A_{x,w}$.
+On the other hand, according to Lemma 3, we have, for $\omega(x^{-1})_{f}$ as defined above Lemma 2,
+$q^{-\ell(x)}(f_{w}\star T_{x^{-1}})(1)$ $=q^{-\ell(x)}\left(\tilde{\phi}(t_{w})\star\tilde{\phi}\left(\phi_{q}({}^{\dagger}T_{x^{-1}})\right)\right)(1)$
+$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}({}^{\dagger}T_{x^{-1}})\right)(1)$
+$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})-\ell(y)}Q_{y,x^{-1}}(q){}^{\dagger}C^{\prime}_{y}\right)\right)(1)$ (26)
+$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}(-1)^{\ell(\omega(x^{-1})_{f})}Q_{y,x^{-1}}(q)(C_{y})\right)\right)(1)$ (27)
+$=(-1)^{\ell(x^{-1})}(-1)^{\ell(\omega(x^{-1})_{f})}q^{-\ell(x)}\tilde{\phi}\left(t_{w}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\\ d\in\mathcal{D}\\ a(d)=a(r)\end{subarray}}h_{y,d,r}t_{r}\right)(1)$ (28)
+$=(-1)^{\ell(x^{-1})+\ell(\omega(x^{-1})_{f})}q^{-\ell(x)}\tilde{\phi}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ d\in\mathcal{D}\\ a(d)=a(r)=a(w)\end{subarray}}h_{y,d,r}t_{w}t_{r}\right)(1)$
+$=(-1)^{\ell(x^{-1})+\ell(\omega(x^{-1})_{f})}q^{-\ell(x)}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d_{w}\\ a(r)=a(w)\end{subarray}}h_{y,d_{w},r}(f_{w}\star f_{r})(1),$ (29)
 where $d_{w}$ is the unique distinguished involution in the right cell containing $w$.
 In line (26), we rewrote $T_{x^{-1}}$ in terms of the $C^{\prime}$-basis of $H$, using the inverse Kazhdan-Lusztig polynomials $Q_{y,x^{-1}}$.
-In line (25), we applied the Goldman involution (see Lemma 1).
+In line (27), we applied the involution $^{\dagger}(-)$ (see Lemma 2).
 In line (28) we applied Lusztig’s map $\phi_{q}$, and then in line (29), we applied the map $\tilde{\phi}$.
@@
 $q^{-\ell(x)}q^{\frac{\ell(y)}{2}}q^{\frac{\ell(x^{-1})-\ell(y)-1}{2}}=q^{-\ell(x)}q^{\frac{\ell(y)}{2}}q^{\frac{\ell(x)-\ell(y)-1}{2}}=q^{\frac{-\ell(x)-1}{2}}\to 0$ (30)
-as $\ell(x)\to\infty$. This completes the definition of $\phi_{1}$ as a map of $\mathcal{A}$-modules.
-It is easy to see that $\phi_{1}$ is a morphism of rings, essentially because $\tilde{\phi}$ is. Indeed, we have
-$\phi_{1}(t_{w}t_{w^{\prime}})=\sum_{z}\gamma_{w,w^{\prime},z^{-1}}\sum_{x}a_{x,z}T_{x}$ (29)
+as $\ell(x)\to\infty$. This completes the definition of $\hat{\phi}$ as a map of $\mathcal{A}$-modules.
+It is easy to see that $\hat{\phi}$ is a morphism of rings, essentially because $\tilde{\phi}$ is. Indeed, we have
+$\hat{\phi}(t_{w}t_{w^{\prime}})=\sum_{z}\gamma_{w,w^{\prime},z^{-1}}\sum_{x}a_{x,z}T_{x}$ (31)
 while on the other hand
-$\phi_{1}(t_{w})\phi_{1}(t_{w^{\prime}})=\sum_{x}a_{x,w}T_{x}\cdot\sum_{y}a_{y,w^{\prime}}T_{y}$ (30)
+$\hat{\phi}(t_{w})\hat{\phi}(t_{w^{\prime}})=\sum_{x}a_{x,w}T_{x}\cdot\sum_{y}a_{y,w^{\prime}}T_{y}$ (32)
 and when $\mathbf{q}=q$, we have that (32) becomes by definition
 $\tilde{\phi}(t_{w})\star\tilde{\phi}(t_{w^{\prime}})=\tilde{\phi}\left(\sum_{z}\gamma_{w,w^{\prime},z^{-1}}t_{z}\right)=\sum_{z}\sum_{x}\gamma_{w,w^{\prime},z^{-1}}A_{x,z}T_{x}.$
-[…]
+Hence for infinitely many prime powers we have that the specializations of (31) agrees with those of (32), and hence (31) is equal to (32) in $\mathbfcal{H}^{-}$.
+A similar argument shows that $\hat{\phi}$ preserves units.
+∎
 Remark 13.
-[…]
+The proof, specifically (30), gives a necessary condition for an element of $\mathbfcal{H}^{-}$ to belong to the image of $\hat{\phi}$: the coefficients must decay asymptotically at least as fast as $\mathbf{q}^{-\frac{\ell(x)}{2}}$.
 Proposition 3.
 There is a commutative diagram
-[…]
-[…]
+$\mathbf{H}$$J\otimes_{\mathbb{C}}\mathcal{A}$$\mathbfcal{H}^{-}$$\mathbfcal{H}^{-}_{C}$$\mathbfcal{H}^{-}_{C^{\prime}},$$\scriptstyle{\phi}$$\scriptstyle{\hat{\phi}}$$\scriptstyle{\phi^{-1}}$$\scriptstyle{{}^{\dagger}(-)}$
+and we have $\hat{\phi}={}^{\dagger}(-)\circ\phi^{-1}$ as morphisms of $\mathcal{A}$-algebras $J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$.
+In particular, $a_{x,w}$ has integer coefficients for all $x,w\in W_{\mathrm{aff}}$.
 Proof.
-[…]
-Note that as $\tilde{\phi}\circ\phi_{q}=()^{\dagger}$ on $H$ for all $q$, we have that $\phi_{1}=()^{\dagger}\circ\phi^{-1}$ on $\phi(\mathbf{H})$. This says that the diagram commutes.
-[…]
-$\phi_{1}\left(\sum_{w}b_{w}t_{w}\right)=\sum_{y}b^{\prime}_{y}T_{y},$
+The second claim follows from the first if we show that $\hat{\phi}$ extends to a continuous morphism $\mathbfcal{J}\to\mathbfcal{H}^{-}$, by density of $\phi(\mathbf{H})$ in $\mathbfcal{J}\simeq\mathbfcal{H}^{-}_{C}$, and the third claim follows from the second and the fact that the completions we consider are actually defined over $\mathbb{Z}$ [affineII, Thm.
+2.8].
+Note that as $\tilde{\phi}\circ\phi_{q}={}^{\dagger}(-)$ on $H$ for all $q$, we have that $\hat{\phi}={}^{\dagger}(-)\circ\phi^{-1}$ on $\phi(\mathbf{H})$. This says that the diagram commutes.
+We now show that $\hat{\phi}$ extends to a continuous map $\mathbfcal{J}\to\mathbfcal{H}^{-}_{C}$. Let $\sum_{w}b_{w}t_{w}$ define an element of $\mathbfcal{J}$ and define
+$\hat{\phi}\left(\sum_{w}b_{w}t_{w}\right)=\sum_{y}b^{\prime}_{y}T_{y},$
 where
@@
 Therefore $b^{\prime}_{y}$ is well-defined, and as $a_{y,w}\to 0$ as $\ell(y)\to\infty$, we have $b^{\prime}_{y}\to 0$ as $\ell(y)\to\infty$.
-Therefore $\phi_{1}$ extends to $\mathbfcal{J}$.
-To show continuity, it suffices to show that if $\{\sum_{w}b_{w,n}t_{w}\}_{n}$ is a sequence of elements of $\mathbfcal{J}$ tending to 00 as $n\to\infty$, then
-$\sum_{w}b_{w,n}\phi_{1}(t_{w})=\sum_{y}b^{\prime}_{y,n}T_{y}\to 0$
-[…]
-[…]
+Therefore $\hat{\phi}$ extends to $\mathbfcal{J}$.
+To show continuity, it suffices to show that if $\{\sum_{w}b_{w,n}t_{w}\}_{n}$ is a sequence of elements of $\mathbfcal{J}$ tending to 0 as $n\to\infty$, then
+$\sum_{w}b_{w,n}\hat{\phi}(t_{w})=\sum_{y}b^{\prime}_{y,n}T_{y}\to 0$
+as $n\to\infty$ in $\mathbfcal{H}^{-}$, where $b^{\prime}_{y,n}=\sum_{w}b_{w,n}a_{y,w}$.
+For all $R>0$, there is $N>0$ such that $n>N$ implies $b_{x,n}\in(\mathbf{q}^{-1/2})^{R}\hat{\mathcal{A}}^{-}$ for all $x$.
+We have seen that there is $M$ depending only on $W_{\mathrm{aff}}$ such that $a_{y,w}\in(\mathbf{q}^{1/2})^{M}\hat{\mathcal{A}}^{-}$ for all $w,y$.
+Therefore $b^{\prime}_{y,n}\to 0$ as $n\to\infty$, because $b_{w,n}\to 0$ as $n\to\infty$.
+∎
+Note that Proposition 3 means in particular that $\hat{\phi}(J)\subset\mathbfcal{H}^{-}_{C^{\prime}}$.
 Corollary 4.
 We have
-$(\phi\circ()^{\dagger})^{-1}(t_{1})=f_{1}(1)\sum_{w\in\tilde{W}}(-1)^{\ell(w)}q^{-\ell(w)}T_{w}.$
+$(\phi\circ^{\dagger}(-))^{-1}(t_{1})=f_{1}(1)\sum_{w\in W_{\mathrm{aff}}}(-1)^{\ell(w)}q^{-\ell(w)}T_{w}.$
 Proof.
-If $w=1$, everything in (27) reduces to $r=d_{w}=y=1$, and we need only recall that $Q_{1,x}(\mathbf{q})=1$ for any $x\in\tilde{W}$.
+If $w=1$, everything in (29) reduces to $r=d_{w}=y=1$, and we need only recall that $Q_{1,x}(\mathbf{q})=1$ for any $x\in W_{\mathrm{aff}}$.
 This follows from unicity of the inverse Kazhdan-Lusztig polynomials and the identity $\sum_{x\leq w}(-1)^{\ell(x)}P_{x,w}(\mathbf{q})=0$ for $w\neq 1$ [BjornerBrenti, Exercise 5.17].
@@
 We must show that $\tilde{\phi}(j)\neq 0$.
-By injectivity of $\phi_{1}$, we have $\phi_{1}(j)\neq 0$.
-By definition of the map $\phi_{1}$, this means that there exists $q_{0}>1$, $u,s\in G^{\vee}$ such that $us=su$ with $s$ compact, and a representation $\rho$ of $\pi_{0}(Z_{G^{\vee}}(u,s))$ such that $jK(u,s,\rho,q_{0})\neq 0$.
+By injectivity of $\hat{\phi}$, we have $\hat{\phi}(j)\neq 0$.
+By definition of the map $\hat{\phi}$, this means that there exists $q_{0}>1$, $u,s\in G^{\vee}$ such that $us=su$ with $s$ compact, and a representation $\rho$ of $\pi_{0}(Z_{G^{\vee}}(u,s))$ such that $jK(u,s,\rho,q_{0})\neq 0$.
 But then $jK(u,s,\rho,q)\neq 0$, $K(u,s,\rho,q)$ being a different specialization of the restriction of the same $J$-module $E(u,s,\rho)$ as for $K(u,s,\rho,q_{0})$, and $K(u,s,\rho,q)$ is also tempered.
@@
 ∎
-3.4 Proof of Theorem 13 for $\mathrm{GL}_{n}$
+3.4 The case of $\mathrm{GL}_{n}$
 Theorem 13.
-Let $\tilde{W}$ be of type $\tilde{A}_{n}$.
+Let $W_{\mathrm{aff}}$ be of type $\tilde{A}_{n}$.
 Then statements 1 and 2 in the statement of Theorem 2 are true, together with a stronger version of statement 3: Let $u=(r_{1},\dots,r_{k})$ be a unipotent conjugacy class in $\mathrm{GL}_{n}(\mathbb{C})$.
-Let $d$ be a distinguished involution in the two-sided cell $\mathbf{c}(u)\subset\tilde{W}(\mathrm{GL}_{n})$ corresponding to $u$.
+Let $d$ be a distinguished involution in the two-sided cell $\mathbf{c}(u)\subset W_{\mathrm{aff}}(\mathrm{GL}_{n})$ corresponding to $u$.
 Let $\pi$ be the unique family of parabolic inductions that $t_{d}$ does not annihilate.
@@
 This proves the first claim.
-The second claim now follows from the first claim and the first statement of Corollary 2 and the fact that $\tilde{W}$ has finitely-many two-sided cells.
+The second claim now follows from the first claim and the first statement of Corollary 3 and the fact that $W_{\mathrm{aff}}$ has finitely-many two-sided cells.
 Finally, let $u=(r_{1},\dots,r_{k})$.
@@
 Considering traces and invoking Proposition 3, we see that the denominator of every $a_{x,w}$ divides a power of $P_{W}(\mathbf{q})$.
-Therefore there is $N=N_{\tilde{W}}$ depending only $\tilde{W}$ such that we can take $P^{1}_{\mathbf{G}}(\mathbf{q})=P_{W}(\mathbf{q})^{N}$.
+Therefore there is $N=N_{W_{\mathrm{aff}}}$ depending only $W_{\mathrm{aff}}$ such that we can take $P^{1}_{\mathbf{G}}(\mathbf{q})=P_{W}(\mathbf{q})^{N}$.
 This proves the second statement of the Theorem.
@@
 Let $\pi$ be a simple tempered representation of $G$ with $I$-fixed vectors with Kazhdan-Lusztig parameter $(u,s,\rho)$.
-Let $w_{\mathcal{P}}$ be the longest element in the parabolic subgroup of $\tilde{W}$ defined by $\mathcal{P}$ and $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
+Let $w_{\mathcal{P}}$ be the longest element in the parabolic subgroup of $W_{\mathrm{aff}}$ defined by $\mathcal{P}$ and $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
   1. 1.
@@
 We have
-$\left(\phi\circ()^{\dagger}\right)(C^{\prime}_{w_{\mathcal{P}}})t_{w_{P}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{\overset{d\in\mathcal{D}}{z\sim_{L}d}}h_{w_{\mathcal{P}},d,z}t_{z}t_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{z}h_{w_{\mathcal{P}},w_{\mathcal{P}},z}t_{z}=\operatorname{vol}(\mathcal{P})t_{w_{P}},$
+$\left(\phi\circ^{\dagger}(-)\right)(C^{\prime}_{w_{\mathcal{P}}})t_{w_{P}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{\overset{d\in\mathcal{D}}{z\sim_{L}d}}h_{w_{\mathcal{P}},d,z}t_{z}t_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{z}h_{w_{\mathcal{P}},w_{\mathcal{P}},z}t_{z}=\operatorname{vol}(\mathcal{P})t_{w_{P}},$
 as $t_{z}t_{w_{\mathcal{P}}}\neq 0$ only if $z\sim_{L}w_{\mathcal{P}}$, and $C_{w_{\mathcal{P}}}C_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\operatorname{vol}(\mathcal{P})C_{w_{\mathcal{P}}}$.
@@
 [labels=0,1,2] B[1]2.
-There are five conjugacy classes of parahoric subgroup, each obtained by projection from $\mathrm{GSp}_{4}(F)$: the maximal parahoric $\mathrm{PGSp}_{4}(\mathcal{O})$, the image of the paramodular group $\mathsf{K}$ corresponding to $\{0,2\}$, the image of the Siegel parahoric subgroup $\mathsf{Si}$ corresponding to $\{1\}$, the image of the Klingen parahoric $\mathsf{Kl}$ corresponding to $\{0\}$, and the Iwahori subgroup.
+There are five conjugacy classes of parahoric subgroups, each obtained by projection from $\mathrm{GSp}_{4}(F)$: the maximal parahoric $\mathrm{PGSp}_{4}(\mathcal{O})$, the image of the paramodular group $\mathsf{K}$ corresponding to $\{0,2\}$, the image of the Siegel parahoric subgroup $\mathsf{Si}$ corresponding to $\{1\}$, the image of the Klingen parahoric $\mathsf{Kl}$ corresponding to $\{0\}$, and the Iwahori subgroup.
 The columns of the below table give all the $I$-spherical tempered representations of $G$ by denoting the representation $\operatorname{Ind}_{P}^{G}(\omega)$ by $\omega$.
@@
 Representations attached to the same cell but belonging to different packets are separated with a dotted line.
-$t_{w_{\mathcal{P}}}$\  $\pi=\operatorname{Ind}_{P}^{G}(-)$ $a$ $\nu$ $\nu\otimes\pm\mathrm{St}_{\mathrm{SO}_{3}}$ $\tau_{\mathrm{triv}}$ $\tau_{\mathrm{sgn}}$$\pm\tau_{2}$$\pm\mathrm{St}$ $\mathcal{P}$
+$t_{w_{\mathcal{P}}}\$ $\pi=\operatorname{Ind}_{P}^{G}(-)$ $a$ $\nu$ $\nu\otimes\pm\mathrm{St}_{\mathrm{SO}_{3}}$ $\tau_{\mathrm{triv}}$ $\tau_{\mathrm{sgn}}$    $\pm\tau_{2}$ $\pm\mathrm{St}$ $\mathcal{P}$
 $t_{w_{0}}$ 4 1 $\mathrm{PGSp}_{4}(\mathcal{O})$
@@
 $t_{s_{1}}$ 1 4 1 1 1 $\mathsf{Si}$
-$t_{1}$ 0 8 4 3 1 4 1 $I$
+$t_{1}$ 0 8 4 3 1 4 1 II
+Acknowledgements.
+The author thanks Alexander Braverman for introducing him to this material and the problem, for suggesting the use of the Plancherel formula, and for helpful conversations.
+The author also thanks Roman Bezrukavnikov, Itai Bar-Natan, Gurbir Dhillon, Malors Espinosa-Lara, Julia Gordon, George Lusztig, and Matthew Sunohara for helpful conversations.
+The author thanks Maarten Solleveld for explaining the generalization in [Solleveld] of the results of [fomFormalDegUnip], suggesting the reference [APM], and for careful reading of an early version of this manuscript.
+The author owes a great debt to the anonymous referee for careful reading resulting in many corrections and improvements both mathematical and expositional.
+Some of this work was completed at the Max Planck Institute in Mathematics for Bonn, the directors and staff of which the author thanks for their hospitality. This research was supported by NSERC.
+Competing interests:
+The author declares none.